### v1

# Mixing Time Bounds for the Gibbs Sampler under Isoperimetry

Abstract
We establish bounds on the conductance for the systematic-scan and random-scan Gibbs samplers when the target distribution satisfies a Poincaré or log-Sobolev inequality and possesses sufficiently regular conditional distributions. These bounds lead to mixing time guarantees that extend beyond the log-concave setting, offering new insights into the convergence behavior of Gibbs sampling in broader regimes. Moreover, we demonstrate that our results remain valid for log-Lipschitz and log-smooth target distributions. Our approach relies on novel three-set isoperimetric inequalities and a sequential coupling argument for the Gibbs sampler.
1 Introduction
Sampling from a target distribution has long been a central problem in many fields related to Statistics and Applied Probability. Markov chain Monte Carlo (MCMC) methods are among the most popular algorithms for generating samples from a potentially high-dimensional target distribution $\pi$. The idea is to simulate an ergodic Markov chain that admits $\pi$ as its stationary distribution, and use the states of the chain as approximate samples from $\pi$. The Gibbs sampler (Geman and Geman, 1984) is one of the most widely used MCMC methods in practical applications. This algorithm updates a single coordinate (or a block of coordinates) at each iteration using its conditional distribution under $\pi$, while keeping the other coordinates fixed. There are two common variants: systematic-scan Gibbs and random-scan Gibbs. In the systematic-scan case, coordinates are updated one by one in a fixed order. In random-scan Gibbs, one coordinate is chosen randomly at each step to be updated. The advantage of Gibbs sampling is that it simplifies high-dimensional sampling by reducing it to a series of lower-dimensional updates. Moreover, in many popular settings, such as hierarchical or exponential family models, the conditional distributions are tractable, eliminating the need for step-size selection.
Given the popularity of Gibbs sampling in practical applications, understanding the convergence properties of this algorithm has been of particular importance. Under mild conditions on $\pi$, qualitative convergence for both variants of the Gibbs sampler was established as early as in Roberts and Smith (1994). However, early literature on the quantitative convergence properties of the Gibbs sampler relied on strong assumptions regarding $\pi$. Exact convergence rates were established for multivariate Gaussian distributions (Amit and Grenander, 1991; Amit, 1996; Roberts and Sahu, 1997), as well as for models involving Dobrushin coefficients (Wu, 2006; Dyer et al., 2008; Wang and Wu, 2014; Wang, 2017, 2019). More recently, Yang and Rosenthal (2023) generalized standard convergence tools based on minorization and Lyapunov drift conditions for Gibbs samplers, yielding new results for Bayesian hierarchical models (Qin and Hobert, 2019, 2022); see also Biswas et al. (2022) for similar results using coupling techniques. For such models, Ascolani and Zanella (2024) established further bounds on the convergence rate of the Gibbs sampler, leveraging Bayesian asymptotics under specific random data-generating assumptions.
In this paper, we characterize the convergence of Gibbs sampling using the so-called mixing time, which quantifies the number of iterations needed for the Markov chain to be within a specified distance of its stationary distribution. A common technique for bounding the mixing time is to analyze a quantity known as the conductance of the Markov chain. One of the key components of this technique relies on the notion of isoperimetry for the target distribution. This is often formalized via Cheeger’s isoperimetric inequality (Cheeger, 1970), which holds for log-concave target distributions (Bobkov, 1999) and remains valid under bounded perturbations, as characterized by the criterion of Holley and Stroock (1987). The conductance approach has proven successful for the Hit-and-Run algorithm (Lovász and Simonovits, 1993; Lovász, 1999; Lovász and Vempala, 2004; Vempala, 2005), as well as for random-walk Metropolis-type algorithms (Dwivedi et al., 2019; Mangoubi and Vishnoi, 2019; Chen et al., 2020; Chewi et al., 2021; Zou et al., 2021; Wu et al., 2022; Mou et al., 2022; Chen and Gatmiry, 2023), among others. However, conductance-based proof techniques have seen limited use in the context of Gibbs sampling, with notable exceptions including sampling from uniform distributions over convex bodies (Narayanan and Srivastava, 2022; Laddha and Vempala, 2023), posterior distributions in Bayesian regression models (Lee and Zhang, 2024), and strongly log-concave targets (Wadia, 2024).
In the recent seminal work by Andrieu et al. (2024), the authors unify many ideas from the references mentioned above and provide clear guidelines for controlling the mixing time of general MCMC methods under the assumption of isoperimetry for the target distribution. Specifically, they consider log-concave target distributions which also satisfy either a Poincaré or a log-Sobolev inequality. However, log-concavity is not necessary for these functional inequalities to hold: a wide range of non-log-concave distributions still satisfy a Poincaré or log-Sobolev inequality, making the class of distributions satisfying such functional inequalities substantially broader. Nevertheless, when deriving explicit bounds on the mixing time of the random-walk Metropolis algorithm, even these authors exclusively focus on the strongly log-concave case. To the best of our knowledge, the work of Vempala and Wibisono (2019) on the unadjusted Langevin algorithm is the only existing result that establishes convergence bounds under a Poincaré or log-Sobolev inequality rather than relying on Cheeger’s isoperimetric inequality. It is therefore interesting to investigate whether the framework developed by Andrieu et al. (2024) can be applied to the Gibbs sampler targeting non-log-concave distributions, specifically those satisfying either a Poincaré or log-Sobolev inequality.
1.1 Our contributions
In this paper, we establish conductance bounds for the systematic-scan and random-scan Gibbs samplers over a broad class of target distributions supported on $\mathbb{R}^{d}$. In particular, consider a target distribution $\pi$ satisfying a $L^{q}$-Poincaré inequality with constant $\mathtt{p}_{q}$ and whose one-dimensional conditional distributions are Hölder continuous in total variation distance with constant $M$ and exponent $\beta$ (see Subsection 2.2 for formal definitions). Denote by $P_{\mathrm{SS}}$ and $P_{\mathrm{RS}}$ the Markov kernels corresponding to the systematic-scan and random-scan Gibbs samplers, respectively. Then, for sufficiently large $d$, our main results (Theorem 4.7 and Corollary 4.9) provide the following bounds on the conductance of the systematic-scan Gibbs kernel, $\Phi(P_{\mathrm{SS}})$, and the random-scan Gibbs kernel, $\Phi(P_{\mathrm{RS}})$:
$\Phi(P_{\mathrm{SS}})\gtrsim\frac{\mathtt{p}_{q}}{\left(Md\right)^{q/\beta}},$
$\Phi(P_{\mathrm{RS}})\gtrsim\frac{\mathtt{p}_{q}}{M^{q/\beta}(d\cdot\log d)^{q/\beta+1}},$
where $\gtrsim$ denotes inequalities that hold up to a universal constant. Bounds on the mixing time in total variation distance then follow directly from Lovász and Simonovits (1993, Corollary 1.5).
To prove the conductance bounds, we establish novel intermediate results that may also be of independent interest.
  • We prove new three-set isoperimetric inequalities for distributions satisfying a $L^{q}$-Poincaré (or log-Sobolev) inequality, without assuming log-concavity (Propositions 4.1 and 4.3). This is achieved by applying these functional inequalities to a carefully chosen test function.
  • Assuming that the one-dimensional conditional distributions under $\pi$ are Hölder continuous in total variation distance, we derive close coupling conditions for the kernels $P_{\mathrm{SS}}$ and $P^{N_{d}}_{\mathrm{RS}}$, where $N_{d}\sim d\cdot\log d$ (Propositions 4.4 and 4.6). These conditions are established through sequential coupling arguments. Here, $P^{N_{d}}_{\mathrm{RS}}$ denotes the random-scan Gibbs kernel iterated $N_{d}$ times, yielding a kernel that is supported on the entire space $\mathbb{R}^{d}$.
  • We show that if $\pi$ is log-Lipschitz continuous or log-smooth, then the one-dimensional conditional distributions under $\pi$ are Hölder continuous in total variation distance (Lemmas 5.4 and 5.5). The proof relies on the fact that log-Lipschitz continuity and log-smoothness are inherited by the marginal distributions of $\pi$, as shown in Lemma 5.3.
After establishing the three-set isoperimetric inequalities and close coupling conditions, bounds on the conductance follow from Proposition 3.10 below, which is a slight modification of Andrieu et al. (2024, Lemma 15).
1.2 Connections with recent work
As a special case, the conductance approach has been successfully applied to the random-scan Gibbs sampler targeting uniform distributions over convex bodies, where quantitative convergence guarantees have been established (Narayanan and Srivastava, 2022; Laddha and Vempala, 2023). Narayanan and Srivastava (2022) establish a close coupling condition for a multi-step Metropolis-within-Gibbs scheme, where the corresponding Markov kernel updates a randomly chosen coordinate using a single step of random-walk Metropolis while keeping the others fixed. They then compare the conductance of this auxiliary Markov chain with that of the original random-scan Gibbs sampler. Laddha and Vempala (2023) take a different approach, proving a three-set isoperimetric inequality for partitions of space into axis-disjoint regions. This allows them to control the mixing behavior without relying on auxiliary chains. In our work, we extend these ideas to general target distributions and establish three-set isoperimetric inequalities that hold for arbitrary partitions.
Additionally, the conductance approach has been used to establish rapid mixing of the Gibbs sampler targeting the posterior distributions in Bayesian probit, logit, and lasso regression models (Lee and Zhang, 2024). In each case, the Gibbs sampler is made tractable by augmenting the state space with latent variables. For such data-augmented chains, Lee and Zhang (2024) show that close coupling conditions follow directly from the conditional independence structure of the posterior distribution. Isoperimetric inequalities are obtained via transference inequalities to a transformed Markov chain with a log-concave target. Under suitable assumptions on the data-generating process, these authors then establish bounds on the mixing time that hold with high probability.
More generally, mixing time bounds for the random-scan Gibbs sampler targeting log-smooth and strongly log-concave distributions have recently been established by Ascolani et al. (2024a); Wadia (2024). In particular, Ascolani et al. (2024a) prove that the random-scan Gibbs sampler contracts in relative entropy and provide a sharp characterization of the contraction rate. Their approach relies on triangular transport maps to decompose relative entropy into coordinate-wise terms. Wadia (2024) follow a similar approach to Laddha and Vempala (2023), deriving a three-set isoperimetric inequality for partitions of space into axis-disjoint regions in the setting of general log-smooth and strongly log-concave distributions. The key insight behind their approach is that such distributions are approximately uniform on small enough domains. Both proof strategies rely heavily on the log-smoothness and log-concavity of $\pi$, so it is unclear whether they can be extended to more general settings, such as when the target distribution satisfies a Poincaré or log-Sobolev inequality.
1.3 Organization
The rest of this paper is structured as follows. Section 2 introduces the Gibbs sampler and outlines the relevant assumptions on the target distribution. Section 3 reviews standard definitions and key results on Markov chain mixing times. In particular, we explain how a three-set isoperimetric inequality and a close coupling condition can be used to derive lower bounds on conductance. Next, we establish three-set isoperimetric inequalities for target distributions satisfying a Poincaré or log-Sobolev inequality (Subsection 4.1) and derive close coupling conditions for the systematic-scan and random-scan Gibbs kernels (Subsection 4.2). We then use these results to prove conductance and mixing time bounds for both variants of the Gibbs sampler (Subsection 4.3). Finally, Section 5 extends these bounds to log-Lipschitz continuous and log-smooth target distributions. We also provide examples of non-log-concave target distributions for which our results can be applied to establish mixing time bounds for both variants of the Gibbs sampler, filling a gap not addressed by any other existing literature.
2 Preliminaries
2.1 The Gibbs sampler
We consider a probability distribution $\pi(\mathrm{d}x)$ supported on $\mathbb{R}^{d}$, equipped with its standard Borel $\sigma$-algebra. Let $P(x,\mathrm{d}y)$ be the Markov transition kernel of a discrete-time Markov chain on $\mathbb{R}^{d}$ with stationary (or invariant) distribution $\pi$. Here, the Markov chain is interpreted as the MCMC method for generating approximate samples from $\pi$. The kernel $P$ is said to be reversible with respect to $\pi$ if
$\pi(\mathrm{d}x)P(x,\mathrm{d}y)=\pi(\mathrm{d}y)P(y,\mathrm{d}x)\qquad\forall x,y\in\mathbb{R}^{d}.$
Reversibility with respect to $\pi$ implies stationarity of $\pi$, however the reverse implication is in general false. Among MCMC methods, we focus on the Gibbs sampler. This method has two commonly used variants: the random-scan Gibbs sampler and the systematic-scan Gibbs sampler, which we now introduce. Let $x=\left(x_{1},\dots,x_{d}\right)\in\mathbb{R}^{d}$. For $1\leq k\leq l\leq d$, we write
$x_{k:l}=(x_{k},x_{k+1},\dots,x_{l-1},x_{l})\in\mathbb{R}^{l-k+1}$
and
$x_{-k}=(x_{1},\dots,x_{k-1},x_{k+1},\dots,x_{d})\in\mathbb{R}^{d-1}.$
Given another point $y=\left(y_{1},\dots,y_{d}\right)\in\mathbb{R}^{d}$, we also write
$(x_{-k},y_{k})=(x_{1},\dots,x_{k-1},y_{k},x_{k+1},\dots,x_{d}).$ (1)
We additionally use the notation $\left[d\right]=\left\{1,\dots,d\right\}$. Let $X\sim\pi$. Denote the conditional distribution of $X_{k}$ given $X_{-k}=x_{-k}$ as $\pi(\cdot\mid x_{-k})$. For all $k\in[d]$, define the Markov kernel
$P_{k}(x,\mathrm{d}y)=\pi(\mathrm{d}y_{k}\mid x_{-k})\delta_{x_{-k}}(\mathrm{d}y_{-k}),$
where $\delta_{x_{-k}}$ denotes the $(d-1)$-dimensional Dirac delta distribution: for all measurable $S\subseteq\mathbb{R}^{d-1}$,
$\delta_{x_{-k}}(S)=\begin{cases}1&\text{if }x_{-k}\in S,\\ 0&\text{if }x_{-k}\notin S.\end{cases}$
Using the kernels $P_{k}$, we now define the Markov kernels corresponding to the random-scan and systematic-scan Gibbs samplers. For two Markov kernels $P$ and $Q$, we denote their composition as $PQ(x,\mathrm{d}y)=\int P(u,\mathrm{d}y)Q(x,\mathrm{d}u)$.
Definition 2.1.
The Markov kernel associated with the random-scan Gibbs sampler is defined as the mixture of the Markov kernels $P_{k}$:
$P_{\mathrm{RS}}=\frac{1}{d}\sum_{k=1}^{d}P_{k}.$
The Markov kernel associated with the systematic-scan Gibbs sampler is defined as the composition of the Markov kernels $P_{k}$:
$P_{\mathrm{SS}}=P_{d}P_{d-1}\dots P_{2}P_{1}.$
In words, the random-scan Gibbs sampler selects a single coordinate uniformly at random and resamples it from its conditional distribution, while leaving the other coordinates unchanged. In contrast, the systematic-scan Gibbs sampler sequentially updates each coordinate in a predetermined order, from 1 to $d$, resampling each from its respective conditional distribution.
It is worth noting that the Markov kernel associated with the random-scan Gibbs sampler is $\pi$-reversible, whereas the Markov kernel associated with the systematic-scan Gibbs sampler is generally not. To address this, one approach is to use the multiplicative reversibilization of the systematic-scan Gibbs sampler. The kernel for this version is defined as
$P_{\mathrm{SS}}^{*}P_{\mathrm{SS}}=P_{1}P_{2}\dots P_{d-1}P_{d}P_{d-1}\dots P_{2}P_{1},$
where $P^{*}$ represents the adjoint operator of a Markov kernel $P$. In this version, the coordinates are updated first in increasing order, from $1$ to $d$, and then in decreasing order, from $d$ to $1$.
2.2 Assumptions on the target distribution
[…]
Definition 2.2.
Let $q\geq 1$. The distribution $\pi$ is said to satisfy a $L^{q}$-Poincaré inequality ($\mathrm{PI}_{q}$) with constant $\mathtt{p}_{q}>0$ if for any locally Lipschitz function $f:\mathbb{R}^{d}\to\mathbb{R}$,
$\mathbb{E}\left[\left\lvert f-\mathbb{E}[f]\right\rvert^{q}\right]\leq\frac{1}{\mathtt{p}_{q}}\cdot\mathbb{E}\left[\left\lvert\nabla f\right\rvert^{q}\right].$ (2)
We take $\mathtt{p}_{q}$ to be the largest constant for which this inequality holds.
Definition 2.3.
Let $q\geq 1$. The distribution $\pi$ is said to satisfy a $L^{q}$-log-Sobolev inequality ($\mathrm{LSI}_{q}$) with constant $\mathtt{s}_{q}>0$ if for any locally Lipschitz function $f:\mathbb{R}^{d}\to\mathbb{R}$,
$\mathbb{E}\left[\left\lvert f\right\rvert^{q}\cdot\log\left\lvert f\right\rvert^{q}\right]-\mathbb{E}\left[\left\lvert f\right\rvert^{q}\right]\cdot\log\mathbb{E}\left[\left\lvert f\right\rvert^{q}\right]\leq\frac{1}{\mathtt{s}_{q}}\cdot\mathbb{E}\left[\left\lvert\nabla f\right\rvert^{q}\right].$ (3)
As before, we take $\mathtt{s}_{q}$ to be the largest constant for which this inequality holds.
These functional inequalities extend the classical assumption of (strong) log-concavity for $\pi$. Specifically, Bobkov (1999) proved that $\mathrm{PI}_{1}$ is satisfied by all log-concave distributions, and Bakry and Émery (1985) showed that $\mathrm{LSI}_{2}$ holds for all strongly log-concave distributions. As shown in Bobkov and Zegarlinski (2005, Proposition 6.2), $\mathrm{LSI}_{q}$ is also preserved by bounded perturbation à la Holley and Stroock (1987). More generally, Ledoux (1999) proved that $\mathrm{PI}_{2}$ and $\mathrm{LSI}_{2}$ respectively imply sub-exponential and sub-Gaussian tail bounds for the distribution $\pi$. Log-Sobolev inequalites have also been proven for mixture distributions (Chen et al., 2021).
Furthermore, if $\pi$ satisfies $\mathrm{LSI}_{q}$ with constant $\mathtt{s}_{q}$, then Bobkov and Zegarlinski (2005, Theorem 2.1) showed that $\pi$ satisfies $\mathrm{PI}_{q}$ with constant $\mathtt{p}_{q}\geq 4\mathtt{s}_{q}/\log 2$. In addition, it can be shown that these functional inequalities form a hierarchy: if $\pi$ satisfies $\mathrm{PI}_{q}$, then $\pi$ satisfies $\mathrm{PI}_{r}$ for any $1\leq q\leq r$ (Milman, 2009a, Proposition 2.5). An analogous result holds for log-Sobolev inequalities, as formalized in the following proposition.
Proposition 2.4.
Let $1\leq q\leq r$. Suppose that the distribution $\pi$ satisfies $\mathrm{LSI}_{q}$ with constant $\mathtt{s}_{q}$. Then, $\pi$ also satisfies $\mathrm{LSI}_{r}$ with constant
$\frac{5}{128}\left(\frac{4}{105}\right)^{r/q}\left(\frac{q}{r}\right)^{r}\mathtt{s}_{q}^{r/q}\leq\mathtt{s}_{r}.$
To the best of our knowledge, Proposition 2.4 does not appear explicitly in the literature. For completeness, we include a proof in Appendix A, which may also be of independent interest.
We now introduce one final regularity assumption on the conditional distributions of $\pi$. To formalize this, recall that the total variation (TV) distance between two probability distributions $\mu$ and $\nu$ supported on $\mathbb{R}^{d}$ is given by
$\left\|\mu-\nu\right\|_{\textsc{TV}}=\sup\left\{\left\lvert\mu(S)-\nu(S)\right\rvert:S\subseteq\mathbb{R}^{d}\text{ measurable}\right\}.$
Definition 2.5.
Let $M>0$ and $\beta\in(0,1]$. The conditional distributions of $\pi$ are said to be $(M,\beta)$-TV continuous if
$\left\|\pi(\cdot\mid x_{-k})-\pi(\cdot\mid y_{-k})\right\|_{\textsc{TV}}\leq M\left\lvert x_{-k}-y_{-k}\right\rvert^{\beta}\qquad\forall k\in[d],~{}\forall x,y\in\mathbb{R}^{d}.$
This type of continuity has been shown to hold for posterior distributions in Bayesian models, under assumptions such as Lipschitz continuity of the likelihood and sufficient integrability of the prior (Dolera and Mainini, 2020, 2023). Moreover, in Subsection 5.1, we will show how to verify that the conditional distributions of $\pi$ are TV-continuous under standard conditions on the potential function $U=-\log\pi$.
3 Convergence analysis of Markov chains
3.1 Mixing times of Markov chains
We begin by recalling key definitions and concepts related to the mixing of Markov chains. Let $\mu$ denote the initial distribution of the Markov chain associated to a kernel $P$. After $k$ steps, the distribution of the chain is given by $\mu P^{k}$, where
$\mu P^{k}(\mathrm{d}x)=\int P^{k}(y,\mathrm{d}x)\mu(\mathrm{d}y).$
It is often assumed that the initial distribution is “not too far” from the stationary distribution, a condition referred to as a warm start.
Definition 3.1.
Let $\Omega>0$. A $\pi$-invariant Markov chain with initial distribution $\mu$ is said to have a $\Omega$-warm start if
$\sup\left\{\frac{\mu(S)}{\pi(S)}:S\subseteq\mathbb{R}^{d}\text{ measurable}\right\}\leq\Omega.$
Note that the constant $\Omega$ can grow exponentially with the dimension $d$ even in fairly innocent examples. For instance, this occurs when the initial distribution is a Gaussian centered at the mode of a strongly log-concave target (Dwivedi et al., 2019). Nevertheless, when $\pi$ is strongly log-concave, Altschuler and Chewi (2024) showed that a distribution with a $\mathcal{O}(1)$-warm start can be efficiently obtained using the underdamped Langevin dynamics in $\mathcal{O}(\sqrt{d})$ iterations. We now define the mixing time of a Markov chain as the minimum number of iterations required for the chain to be “close” to its stationary distribution in total variation distance.
Definition 3.2.
Let $\zeta>0$. The $\zeta$-mixing time of a $\pi$-invariant Markov chain with transition kernel $P$ is defined as
$\tau(\zeta,P)=\inf\left\{k\geq 0:\left\|\mu P^{k}-\pi\right\|_{\textsc{TV}}\leq\zeta\right\}.$
A common approach to bound the mixing time is to study a quantity known as the conductance of the Markov chain, which is defined as follows.
Definition 3.3.
Consider a $\pi$-invariant Markov chain with transition kernel $P$. The conductance of $P$ is defined as
$\Phi(P)=\inf\left\{\frac{\int_{S}P(x,S^{c})\pi(\mathrm{d}x)}{\pi(S)}:S\subseteq\mathbb{R}^{d}\text{ measurable},\text{ }\pi(S)\leq\frac{1}{2}\right\}.$
The connection between the mixing time of a Markov chain and its conductance was established by Lovász and Simonovits (1993), who extended the earlier work of Jerrum and Sinclair (1988) to general state spaces. Importantly, this result does not require the transition kernel $P$ to be reversible.
Theorem 3.4 (Lovász and Simonovits, 1993, Corollary 1.5).
Consider a $\pi$-invariant Markov chain with transition kernel $P$ initialized at a $\Omega$-warm start. Then, for all $k\geq 0$,
$\left\|\mu P^{k}-\pi\right\|_{\textsc{TV}}\leq\sqrt{\Omega}\cdot\exp\left(-\frac{k}{2}\cdot\Phi(P)^{2}\right).$
Corollary 3.5.
Consider a $\pi$-invariant Markov chain with transition kernel $P$ initialized at a $\Omega$-warm start. Then, its $\zeta$-mixing time can be bounded by
$\tau(\zeta,P)\leq\frac{2}{\Phi(P)^{2}}\cdot\log\left(\frac{\sqrt{\Omega}}{\zeta}\right).$
The problem of upper bounding the mixing time is thus reduced to proving a lower bound on the conductance.
3.2 The spectral gap
Let
$L^{2}_{\pi}=\left\{f:\mathbb{R}^{d}\to\mathbb{R}:\int f(x)^{2}\pi(\mathrm{d}x)<\infty\right\}$
denote the Hilbert space of square $\pi$-integrable functions. For all $f,g\in L^{2}_{\pi}$, the scalar product associated to this space is
$\langle f,g\rangle=\int f(x)g(x)\pi(\mathrm{d}x),$
with corresponding $L^{2}_{\pi}$-norm $\left\|f\right\|_{2}=\sqrt{\langle f,f\rangle}$. Note that a Markov kernel $P$ can be seen as an operator that acts on functions $f\in L^{2}_{\pi}$ via
$Pf(x)=\int f(y)P(x,\mathrm{d}y).$
Let also $I$ denote the identity operator: $If=f$.
Definition 3.6.
Consider a $\pi$-invariant Markov chain with transition kernel $P$. Suppose that $P$ is $\pi$-reversible. The spectral gap of $P$ is defined as
$\lambda_{2}(P)=\inf\left\{\frac{\langle f,(I-P)f\rangle}{\left\|f\right\|_{2}}:f\in L^{2}_{\pi},\text{ }\mathbb{E}[f]=0\right\}.$
When $P$ is reversible, the spectral gap provides a bound on the $L^{2}_{\pi}$-convergence rate of the Markov chain: for any $f\in L^{2}_{\pi}$ with $\mathbb{E}\left[f\right]=0$, we have
$\left\|Pf\right\|_{2}\leq\left(1-\lambda_{2}\right)\left\|f\right\|_{2}.$
See (Levin and Peres, 2017, Exercise 12.4) for a proof. The spectral gap of a reversible Markov kernel $P$ is related to its conductance via the following bounds, also known as Cheeger’s inequalities (Cheeger, 1970).
Theorem 3.7 (Lawler and Sokal, 1988, Theorem 2.1).
Let $\Phi(P)$ and $\lambda_{2}(P)$ denote respectively the conductance and spectral gap of a reversible Markov kernel $P$. Then,
$\frac{1}{2}\cdot\Phi^{2}(P)\leq\lambda_{2}(P)\leq\Phi(P).$
A lower bound on the conductance thus translates to a lower bound on the spectral gap. Theorem 3.7 applies specifically to the reversible setting and covers the random-scan Gibbs sampler as well as reversibilized forms of the systematic-scan Gibbs sampler, such as the variant with kernel $P_{\mathrm{SS}}^{*}P_{\mathrm{SS}}$ discussed in Subsection 2.1. In this paper, however, we do not directly assume reversibility and instead focus on mixing time bounds as in Corollary 3.5. For non-reversible chains, various notions of spectral gaps have been proposed; see, for instance, Kontoyiannis and Meyn (2012) for definitions based on suitably weighted Banach spaces, and Chatterjee (2023) for finite state spaces. Other approaches (Fill, 1991; Paulin, 2015; Choi, 2020) use reversibilized Markov kernels to derive similar bounds on the convergence rate. However, the sharpness of such bounds in the non-reversible setting remains unclear, and a detailed investigation lies beyond the scope of this work.
3.3 Bounding the conductance
In order to bound the conductance of a $\pi$-invariant Markov kernel $P$, two main components are required: a three-set isoperimetric inequality for the distribution $\pi$, and a close coupling condition for the kernel $P$. In what follows, we write $\mathrm{Id}:\mathbb{R}\to\mathbb{R}$ for the identity function and use $\sqcup$ to denote the union of disjoint sets.
Definition 3.8.
Let $\Upsilon:(0,\infty)\to(0,\infty)$ and $\Psi:(0,1/2)\to(0,\infty)$ be two non-decreasing functions. Additionaly, assume that $\Psi/\mathrm{Id}$ is non-increasing. The distribution $\pi$ is said to satisfy a three-set isoperimetric inequality with functions $\Upsilon,\Psi$, if for all measurable partitions $S_{1}\sqcup S_{2}\sqcup S_{3}=\mathbb{R}^{d}$,
$\pi(S_{3})\geq\Upsilon(\mathcal{D}(S_{1},S_{2}))\cdot\Psi(\min\{\pi(S_{1}),\pi(S_{2})\}),$
where $\mathcal{D}(S_{1},S_{2})=\inf\{\left\lvert x-y\right\rvert:x\in S_{1},y\in S_{2}\}$ denotes the distance between $S_{1}$ and $S_{2}$.
When $\pi$ is log-concave (or a bounded perturbation thereof), a three-set inequality can be derived by controlling a quantity known as the isoperimetric profile of $\pi$ Milman, 2009b, Theorem 1.2; Andrieu et al., 2024, Section 3. Bounds on the isoperimetric profile have also been established when $\pi$ satisfies $\mathrm{PI}_{1}$ and $\mathrm{LSI}_{1}$ (Rothaus, 1985; Bobkov and Houdré, 1997). In all these cases, $\Upsilon$ is taken to be a multiple of the identity function. However, this approach generally breaks down for non-log-concave $\pi$, as shown in Milman (2009a, Subsection 1.2).
Definition 3.9.
Let $\delta,\varepsilon>0$. A Markov kernel $P$ is said to satisfy a $(\delta,\varepsilon)$-close coupling condition if
$\left\lvert x-y\right\rvert\leq\delta\implies\left\|P(x,\cdot)-P(y,\cdot)\right\|_{\textsc{TV}}\leq 1-\varepsilon\qquad\forall x,y\in\mathbb{R}^{d}.$
The close coupling condition ensures that when $x$ and $y$ are “close” in Euclidean distance, the probability distributions $P(x,\cdot)$ and $P(y,\cdot)$ are “close” in total variation distance. Assuming that $\pi$ satisfies a three-set isoperimetric inequality and $P$ satisfies a close coupling condition, we are able to establish a lower bound on the conductance of the Markov chain.
Proposition 3.10.
Consider a $\pi$-invariant Markov kernel $P$. Suppose that $\pi$ satisfies a three-set isoperimetric inequality with functions $\Upsilon,\Psi$, and $P$ satisfies a $(\delta,\varepsilon)$-close coupling condition. Then, for any measurable $S\subseteq\mathbb{R}^{d}$ with $\pi(S)\leq 1/2$,
$\int_{S}P(x,S^{c})\pi(\mathrm{d}x)\geq\frac{\varepsilon}{4}\cdot\min\left\{\pi(S),\Upsilon(\delta)\cdot\Psi\left(\frac{\pi(S)}{2}\right)\right\}.$
Corollary 3.11.
Under the conditions of Proposition 3.10, the conductance of the Markov kernel $P$ can be bounded by
$\Phi(P)\geq\frac{\varepsilon\cdot\Psi(1/4)}{2}\cdot\min\left\{\frac{1}{2\cdot\Psi(1/4)},\Upsilon(\delta)\right\}.$
The proof of Proposition 3.10 can be found in Appendix A and builds on several previous works; see Andrieu et al. (2024, Lemma 15) and the survey by Vempala (2005) for further details. Notably, Proposition 3.10 and Corollary 3.11 do not require the Markov kernel $P$ to be reversible. Conventionally, $\delta$ is chosen to scale inversely with the dimension $d$ to ensure that $\varepsilon$ remains of constant order. Under this regime, Corollary 3.11 yields
$\Phi(P)\gtrsim\min\{1,\Upsilon(\delta)\}=\Upsilon(\delta)$
for sufficiently large $d$.
4 Main results
4.1 Three-set isoperimetric inequalities via functional inequalities
We now establish three-set isoperimetric inequalities for distributions $\pi$ satisfying an $L^{q}$-Poincaré or $L^{q}$-log-Sobolev inequality for any $q\geq 1$. As noted in Subsection 3.3, traditional methods based on the isoperimetric profile cease to be effective when $\pi$ is not log-concave. To overcome these challenges, we take a more direct approach to derive three-set isoperimetric inequalities for distributions satisfying $\mathrm{PI}_{q}$ or $\mathrm{LSI}_{q}$ with $q\geq 1$. In this setting, $\Upsilon$ is no longer restricted to be proportional to the identity function. In what follows, denote the indicator function of a measurable set $S\subseteq\mathbb{R}^{d}$ as $\mathbbm{1}_{S}$. Additionally, define the $L^{q}_{\pi}$-norm of a function $f$ as
$\left\|f\right\|_{q}=\left(\int\left\lvert f(x)\right\rvert^{q}\pi(\mathrm{d}x)\right)^{1/q}.$
Proposition 4.1.
Let $q\geq 1$. If the distribution $\pi$ satisfies $\mathrm{PI}_{q}$ with constant $\mathtt{p}_{q}$, then $\pi$ satisfies a three-set isoperimetric inequality with functions
$\Upsilon(t)=\frac{\mathtt{p}_{q}t^{q}}{2^{2q}\left(1+2^{q}\mathtt{p}_{q}t^{q}\right)},\qquad\Psi(t)=t.$
Proof.
Let $S_{1}\sqcup S_{2}\sqcup S_{3}=\mathbb{R}^{d}$ be a partition of space with $\mathcal{D}(S_{1},S_{2})>0$. Inspired by Ledoux (2001, Proposition 1.7), define the function
$f(x)=\max\left\{0,1-\frac{\mathcal{D}(x,S_{1})}{\mathcal{D}(S_{1},S_{2})}\right\}.$
Observe that the map $x\mapsto\mathcal{D}(x,S_{1})$ is Lipschitz continuous with constant 1. Indeed, for any $\xi>0$, let $s_{\xi}\in S_{1}$ be such that $\left\lvert y-s_{\xi}\right\rvert\leq\mathcal{D}(y,S_{1})+\xi$. Then,
$\mathcal{D}(x,S_{1})$ $\leq\left\lvert x-s_{\xi}\right\rvert$
$\leq\left\lvert x-y\right\rvert+\left\lvert y-s_{\xi}\right\rvert$
$\leq\left\lvert x-y\right\rvert+\mathcal{D}(y,S_{1})+\xi,$
so that $\mathcal{D}(x,S_{1})-\mathcal{D}(y,S_{1})\leq\left\lvert x-y\right\rvert+\xi$. Swapping the roles of $x$ and $y$ and letting $\xi\to 0$, we obtain $\left\lvert\mathcal{D}(x,S_{1})-\mathcal{D}(y,S_{1})\right\rvert\leq\left\lvert x-y\right\rvert$. It follows that the function $f$ is equal to 1 on $S_{1}$, vanishes on $S_{2}$, and is Lipschitz continuous on a subset of $S_{3}$, with the modulus of its gradient equal to $\mathcal{D}(S_{1},S_{2})^{-1}$; elsewhere on $S_{3}$, $f$ vanishes. In particular, $f(x)\in[0,1]$ for $x\in S_{3}$. Applying (2) to $f$, we obtain
$\left\|f-\mathbb{E}[f]\right\|_{q}^{q}=\mathbb{E}\left[\left\lvert f-\mathbb{E}[f]\right\rvert^{q}\right]\leq\frac{1}{\mathtt{p}_{q}}\cdot\mathbb{E}\left[\left\lvert\nabla f\right\rvert^{q}\right]\leq\frac{\pi(S_{3})}{\mathtt{p}_{q}\cdot\mathcal{D}(S_{1},S_{2})^{q}}.$ (4)
We now compare $\mathbb{E}[\left\lvert f-\mathbb{E}[f]\right\rvert^{q}]$ to the corresponding quantity for the indicator function $\mathbbm{1}_{S_{1}}$, namely $\mathbb{E}\left[\left\lvert\mathbbm{1}_{S_{1}}-\mathbb{E}[\mathbbm{1}_{S_{1}}]\right\rvert^{q}\right]$. By the triangle inequality and Jensen’s inequality, we get
$\mathbb{E}\left[\left\lvert\mathbbm{1}_{S_{1}}-\mathbb{E}[\mathbbm{1}_{S_{1}}]\right\rvert^{q}\right]^{1/q}$ $=\left\|\mathbbm{1}_{S_{1}}-\mathbb{E}[\mathbbm{1}_{S_{1}}]\right\|_{q}$
$\leq\left\|f-\mathbbm{1}_{S_{1}}\right\|_{q}+\left\|f-\mathbb{E}[f]\right\|_{q}+\left\lvert\mathbb{E}[f]-\mathbb{E}[\mathbbm{1}_{S_{1}}]\right\rvert$
$\leq 2\left\|f-\mathbbm{1}_{S_{1}}\right\|_{q}+\left\|f-\mathbb{E}[f]\right\|_{q}.$
Using the inequality $\left(a+b\right)^{q}\leq 2^{q}\left(a^{q}+b^{q}\right)$ for $a,b\geq 0$, it follows that
$\mathbb{E}\left[\left\lvert\mathbbm{1}_{S_{1}}-\mathbb{E}[\mathbbm{1}_{S_{1}}]\right\rvert^{q}\right]\leq 2^{q}\left(2^{q}\left\|f-\mathbbm{1}_{S_{1}}\right\|_{q}^{q}+\left\|f-\mathbb{E}[f]\right\|_{q}^{q}\right).$ (5)
Since $f-\mathbbm{1}_{S_{1}}$ vanishes on $S_{1}\sqcup S_{2}$ and takes values in $[0,1]$ on $S_{3}$, we have
$\left\|f-\mathbbm{1}_{S_{1}}\right\|_{q}^{q}=\mathbb{E}\left[\left\lvert f-\mathbbm{1}_{S_{1}}\right\rvert^{q}\right]\leq\pi(S_{3}).$ (6)
Inserting (4) and (6) into (5), we get
$\mathbb{E}\left[\left\lvert\mathbbm{1}_{S_{1}}-\mathbb{E}[\mathbbm{1}_{S_{1}}]\right\rvert^{q}\right]\leq 2^{q}\left(2^{q}+\frac{1}{\mathtt{p}_{q}\cdot\mathcal{D}(S_{1},S_{2})^{q}}\right)\pi(S_{3}).$
We now derive a lower bound for $\mathbb{E}[\left\lvert\mathbbm{1}_{S_{1}}-\mathbb{E}[\mathbbm{1}_{S_{1}}]\right\rvert^{q}]$:
$\mathbb{E}\left[\left\lvert\mathbbm{1}_{S_{1}}-\mathbb{E}[\mathbbm{1}_{S_{1}}]\right\rvert^{q}\right]$ $=\left(1-\pi(S_{1})\right)^{q}\pi(S_{1})+\pi(S_{1})^{q}\left(1-\pi(S_{1})\right)$
$=\left(1-\pi(S_{1})\right)\pi(S_{1})\left[\left(1-\pi(S_{1})\right)^{q-1}+\pi(S_{1})^{q-1}\right]$
$\geq\frac{1}{2^{q-1}}\left(1-\pi(S_{1})\right)\pi(S_{1})$
$\geq\frac{1}{2^{q}}\cdot\min\left\{\pi(S_{1}),1-\pi(S_{1})\right\}$
$\geq\frac{1}{2^{q}}\cdot\min\left\{\pi(S_{1}),\pi(S_{2})\right\},$
where we have used the inequality $1/2\cdot\min\{x,1-x\}\leq x(1-x)$ for $x\in[0,1]$. Combining these bounds, we obtain
$\min\left\{\pi(S_{1}),\pi(S_{2})\right\}\leq 2^{2q}\left(2^{q}+\frac{1}{\mathtt{p}_{q}\cdot\mathcal{D}(S_{1},S_{2})^{q}}\right)\pi(S_{3}),$
or equivalently,
$\pi(S_{3})\geq\frac{\mathcal{D}(S_{1},S_{2})^{q}}{2^{2q}\left(\mathtt{p}_{q}^{-1}+2^{q}\cdot\mathcal{D}(S_{1},S_{2})^{q}\right)}\cdot\min\left\{\pi(S_{1}),\pi(S_{2})\right\}.$
The $L^{q}$-Poincaré inequality thus implies a three-set isoperimetric inequality
$\pi(S_{3})\geq\Upsilon(\mathcal{D}(S_{1},S_{2}))\cdot\min\{\pi(S_{1}),\pi(S_{2})\},$
where
$\Upsilon(t)=\frac{\mathtt{p}_{q}t^{q}}{2^{2q}\left(1+2^{q}\mathtt{p}_{q}t^{q}\right)}.$
To derive a three-set isoperimetric inequality corresponding to the $L^{q}$-log-Sobolev inequality, it is useful to analyze the properties of the function $x\mapsto x\cdot\log(1/x)$. The proof of the following lemma is provided in Appendix A.
Lemma 4.2.
For $x\in[0,1]$, consider the function $\rho(x)=x\cdot\log(1/x)$. Then:
  • The function $\rho(x)$ is positive and achieves its maximum at $x=e^{-1}$.
  • For all $x,h\in[0,1]$ such that $x+h\leq 1$, $\rho(x+h)\geq\rho(x)-h$.
Proposition 4.3.
Let $q\geq 1$. If the distribution $\pi$ satisfies $\mathrm{LSI}_{q}$ with constant $\mathtt{s}_{q}$, then $\pi$ satisfies a three-set isoperimetric inequality with functions
$\Upsilon(t)=\frac{\mathtt{s}_{q}t^{q}}{1+\left(1+e^{-1}\right)\mathtt{s}_{q}t^{q}},\qquad\Psi(t)=\frac{t}{2}\cdot\log\left(\frac{2}{t}\right).$
Proof.
Let $S_{1}\sqcup S_{2}\sqcup S_{3}=\mathbb{R}^{d}$ be a partition of space with $\mathcal{D}(S_{1},S_{2})>0$. Similarly to Proposition 4.1, define the function
$f(x)=\max\left\{0,1-\frac{\mathcal{D}(x,S_{1})}{\mathcal{D}(S_{1},S_{2})}\right\}.$
As seen before, the map $x\mapsto\mathcal{D}(x,S_{1})$ is Lipschitz continuous with constant 1. It follows that the function $f$ is equal to 1 on $S_{1}$, vanishes on $S_{2}$, and is Lipschitz continuous on a subset of $S_{3}$, with the modulus of its gradient equal to $\mathcal{D}(S_{1},S_{2})^{-1}$; elsewhere on $S_{3}$, $f$ vanishes. In particular, $f(x)\in[0,1]$ for $x\in S_{3}$. Applying (3) to $f$, we obtain
$\mathbb{E}\left[\left\lvert f\right\rvert^{q}\cdot\log\left\lvert f\right\rvert^{q}\right]-\mathbb{E}\left[\left\lvert f\right\rvert^{q}\right]\cdot\log\mathbb{E}\left[\left\lvert f\right\rvert^{q}\right]\leq\frac{1}{\mathtt{s}_{q}}\cdot\mathbb{E}\left[\left\lvert\nabla f(x)\right\rvert^{q}\right]\leq\frac{\pi(S_{3})}{\mathtt{s}_{q}\cdot\mathcal{D}(S_{1},S_{2})^{q}}.$ (7)
For $x\in[0,1]$, consider the function $\rho(x)=x\cdot\log(1/x)$. Using the convention that $00\cdot\log(0)=1\cdot\log(1)=$ 0, we can rewrite
$\mathbb{E}\left[\left\lvert f\right\rvert^{q}\cdot\log\left\lvert f\right\rvert^{q}\right]={\int_{S_{3}}}\left\lvert f(x)\right\rvert^{q}\cdot\log\left(\left\lvert f(x)\right\rvert^{q}\right)\pi(\mathrm{d}x)=-{\int_{S_{3}}}\rho\left(\left\lvert f(x)\right\rvert^{q}\right)\pi(\mathrm{d}x)$
and
$\mathbb{E}\left[\left\lvert f\right\rvert^{q}\right]\cdot\log\mathbb{E}\left[\left\lvert f\right\rvert^{q}\right]$ $=\left(\pi(S_{1})+\int_{S_{3}}\left\lvert f(x)\right\rvert^{q}\pi(\mathrm{d}x)\right)\cdot\log\left(\pi(S_{1})+\int_{S_{3}}\left\lvert f(x)\right\rvert^{q}\pi(\mathrm{d}x)\right)$
$=-\rho\left(\pi(S_{1})+\int_{S_{3}}\left\lvert f(x)\right\rvert^{q}\pi(\mathrm{d}x)\right).$
By the first part of Lemma 4.2,
$\mathbb{E}\left[\left\lvert f\right\rvert^{q}\cdot\log\left\lvert f\right\rvert^{q}\right]\geq-e^{-1}\cdot\pi(S_{3}).$ (8)
Noting that $\mathbb{E}[\left\lvert f\right\rvert^{q}]\leq 1$, the second part of Lemma 4.2 gives
$\rho\left(\pi(S_{1})+\int_{S_{3}}\left\lvert f(x)\right\rvert^{q}\pi(\mathrm{d}x)\right)$ $\geq\pi(S_{1})\cdot\log\left(\frac{1}{\pi(S_{1})}\right)-\int_{S_{3}}\left\lvert f(x)\right\rvert^{q}\pi(\mathrm{d}x)$
$\geq\pi(S_{1})\cdot\log\left(\frac{1}{\pi(S_{1})}\right)-\pi(S_{3}),$
and thus
$-\mathbb{E}\left[\left\lvert f\right\rvert^{q}\right]\cdot\log\mathbb{E}\left[\left\lvert f\right\rvert^{q}\right]\geq\pi(S_{1})\cdot\log\left(\frac{1}{\pi(S_{1})}\right)-\pi(S_{3}).$ (9)
Inserting (8) and (9) into (7), we obtain
$\frac{\pi(S_{3})}{\mathtt{s}_{q}\cdot\mathcal{D}(S_{1},S_{2})^{q}}\geq\pi(S_{1})\cdot\log\left(\frac{1}{\pi(S_{1})}\right)-(1+e^{-1})\pi(S_{3}),$
or equivalently,
$\pi(S_{3})\geq\frac{\mathcal{D}(S_{1},S_{2})^{q}}{\mathtt{s}_{q}^{-1}+(1+e^{-1})\cdot\mathcal{D}(S_{1},S_{2})^{q}}\cdot\pi(S_{1})\cdot\log\left(\frac{1}{\pi(S_{1})}\right).$
By swapping $S_{1}$ with $S_{2}$ in the definition of the function $f$, an identical argument yields
$\pi(S_{3})\geq\frac{\mathcal{D}(S_{1},S_{2})^{q}}{\mathtt{s}_{q}^{-1}+(1+e^{-1})\cdot\mathcal{D}(S_{1},S_{2})^{q}}\cdot\pi(S_{2})\cdot\log\left(\frac{1}{\pi(S_{2})}\right).$
Combing these results, we have
$\pi(S_{3})$ $\geq\frac{\mathcal{D}(S_{1},S_{2})^{q}}{\mathtt{s}_{q}^{-1}+(1+e^{-1})\cdot\mathcal{D}(S_{1},S_{2})^{q}}\cdot\max\left\{\pi(S_{1})\cdot\log\left(\frac{1}{\pi(S_{1})}\right),~{}\pi(S_{2})\cdot\log\left(\frac{1}{\pi(S_{2})}\right)\right\}$
$\geq\frac{\mathcal{D}(S_{1},S_{2})^{q}}{\mathtt{s}_{q}^{-1}+(1+e^{-1})\cdot\mathcal{D}(S_{1},S_{2})^{q}}\cdot\min\left\{\pi(S_{1}),\pi(S_{2})\right\}\cdot\log\left(\frac{1}{\min\left\{\pi(S_{1}),\pi(S_{2})\right\}}\right)$
$\geq\frac{\mathcal{D}(S_{1},S_{2})^{q}}{\mathtt{s}_{q}^{-1}+(1+e^{-1})\cdot\mathcal{D}(S_{1},S_{2})^{q}}\cdot\frac{\min\left\{\pi(S_{1}),\pi(S_{2})\right\}}{2}\cdot\log\left(\frac{2}{\min\left\{\pi(S_{1}),\pi(S_{2})\right\}}\right)$
The advantage of this last expression is that the function $x\mapsto x/2\cdot\log(2/x)$ is increasing everywhere on $(0,1/2)$, contrary to the function $x\mapsto x\cdot\log(1/x)$. The $L^{q}$-log-Sobolev inequality thus implies a three-set isoperimetric inequality
$\pi(S_{3})\geq\Upsilon(\mathcal{D}(S_{1},S_{2}))\cdot\Psi(\min\{\pi(S_{1}),\pi(S_{2})\}),$
where
$\Upsilon(t)=\frac{\mathtt{s}_{q}t^{q}}{1+(1+e^{-1})\mathtt{s}_{q}t^{q}},\qquad\Psi(t)=\frac{t}{2}\cdot\log\left(\frac{2}{t}\right).$
Notably, the $L^{q}$-Poincaré and $L^{q}$-log-Sobolev inequalities both lead to three-set isoperimetric inequalities with $\Upsilon(t)\sim\mathtt{p}_{q}t^{q}$ as $t\to 0$. Here the sign $\sim$ means that the two quantities are asymptotic, meaning their ratio converges to a constant as $t\to 0$. This regime is particularly relevant in high-dimensional settings, where $t$ typically scales inversely with the dimension, as per the discussion following Corollary 3.11. In contrast, when $\pi$ is log-concave, we have $\Upsilon(t)\sim t$ as $t\to 0$ (Andrieu et al., 2024, Subsection 3.2).
4.2 Close coupling conditions for the Gibbs kernels
Here, we establish close coupling conditions for the systematic-scan and random-scan Gibbs kernels, assuming that the conditional distributions of the target distribution $\pi$ are TV continuous (as per Definition 2.5). The following results rely on sequential coupling arguments; see Roberts and Rosenthal (2013) for a formal construction of such couplings, and Marton (2004); Wang and Wu (2014) for related coupling techniques in Wasserstein distance. We begin with the systematic-scan Gibbs sampler.
Proposition 4.4.
Suppose that the conditional distributions of $\pi$ are $(M,\beta)$-TV continuous. Then,
$\left\|P_{\mathrm{SS}}(x,\cdot)-P_{\mathrm{SS}}(y,\cdot)\right\|_{\textsc{TV}}\leq 1-\left(1-M\left\lvert x-y\right\rvert^{\beta}\right)^{d}\qquad\forall x,y\in\mathbb{R}^{d}.$
Hence, $P_{\mathrm{SS}}$ satisfies a $(\delta,\varepsilon)$-close coupling condition with
$\delta=\left(Md\right)^{-1/\beta},\qquad\varepsilon=e^{-1}.$
Proof.
The coupling inequality for the total variation distance can be expressed as
$\left\|P_{\mathrm{SS}}(x,\cdot)-P_{\mathrm{SS}}(y,\cdot)\right\|_{\textsc{TV}}=\inf\left\{\mathbb{P}(U\neq V):U\sim P_{\mathrm{SS}}(x,\cdot),V\sim P_{\mathrm{SS}}(y,\cdot)\right\}.$
To construct a candidate coupling of $P_{\mathrm{SS}}(x,\cdot)$ and $P_{\mathrm{SS}}(y,\cdot)$, we introduce intermediate states $U^{(k)}$ and $V^{(k)}$, for $k=0,\dots,d$, which represent the configurations after updating the first $k$ coordinates of two chains initialized at $x$ and $y$, respectively. The coupling proceeds as follows:
  • Initialize $U^{\left(0\right)}=x$ and $V^{\left(0\right)}=y$.
  • For $k=1,\dots,d$, draw a maximal coupling
    $W\sim\pi\left(\cdot\mid U^{\left(k-1\right)}_{-k}\right),\qquad W^{\prime}\sim\pi\left(\cdot\mid V^{\left(k-1\right)}_{-k}\right),$
    such that
    $\mathbb{P}\left(W\neq W^{\prime}\mid U^{\left(k-1\right)}_{-k},V^{\left(k-1\right)}_{-k}\right)=\left\|\pi\left(\cdot\mid U^{\left(k-1\right)}_{-k}\right)-\pi\left(\cdot\mid V^{\left(k-1\right)}_{-k}\right)\right\|_{\textsc{TV}}.$ (10)
    See Lindvall (2002, Theorem 5.2) for an explicit construction of a maximal coupling. Note that both sides of (10) are random quantities depending on $U^{\left(k-1\right)}$, $V^{\left(k-1\right)}$.
  • Update the variables by setting
    $U^{(k)}=\left(U^{(k-1)}_{-k},W\right),\qquad V^{(k)}=\left(V^{(k-1)}_{-k},W^{\prime}\right).$
    In words, only the $k$’th coordinate is updated when transitioning from $U^{(k-1)}$ to $U^{(k)}$ (and from $V^{(k-1)}$ to $V^{(k)}$). The notation used here for indexing $U^{(k)}$ and $V^{(k)}$ follows the convention introduced in (1).
  • Finally, set $U=U^{\left(d\right)}$, $V=V^{\left(d\right)}$.
By construction, we have $U\sim P_{\mathrm{SS}}(x,\cdot)$ and $V\sim P_{\mathrm{SS}}(y,\cdot)$. Moreover, observe that for all $k\in[d]$, we have $U^{\left(k\right)}=\left(U_{1:k},x_{k+1:d}\right)$ and $V^{\left(k\right)}=\left(V_{1:k},y_{k+1:d}\right)$. Using this, (10) can be rewritten as
$\mathbb{P}(U_{k}\neq V_{k}\mid U_{1:k-1},V_{1:k-1})=\left\|\pi\left(\cdot\mid U_{1:k-1},x_{k+1:d}\right)-\pi\left(\cdot\mid V_{1:k-1},y_{k+1:d}\right)\right\|_{\textsc{TV}}.$ (11)
Conditional on the event $\left\{U_{1:k-1}=V_{1:k-1}\right\}$, we have
$\left\lvert\left(U_{1:k-1},x_{k+1:d}\right)-\left(V_{1:k-1},y_{k+1:d}\right)\right\rvert=\left\lvert x_{k+1:d}-y_{k+1:d}\right\rvert.$
Thus, by (11) and the assumption that the conditional distributions of $\pi$ are $(M,\beta)$-TV continuous, it follows that
$\mathbb{P}$ $(U_{k}\neq V_{k}\mid U_{1:k-1}=V_{1:k-1})$
$=\mathbb{E}_{U_{1:k-1},V_{1:k-1}}\left[\left\|\pi\left(\cdot\mid U_{1:k-1},x_{k+1:d}\right)-\pi\left(\cdot\mid V_{1:k-1},y_{k+1:d}\right)\right\|_{\textsc{TV}}\mid U_{1:k-1}=V_{1:k-1}\right]$
$\leq M\left\lvert x_{k+1:d}-y_{k+1:d}\right\rvert^{\beta},$
where the expectation is taken over the joint distribution of $U_{1:k-1},V_{1:k-1}$. We can now bound the total variation distance between $P_{\mathrm{SS}}(x,\cdot)$ and $P_{\mathrm{SS}}(y,\cdot)$ as follows:
$\left\|P_{\mathrm{SS}}(x,\cdot)-P_{\mathrm{SS}}(y,\cdot)\right\|_{\textsc{TV}}$ $\leq\mathbb{P}(U\neq V)$
$=1-\mathbb{P}(U=V)$
$=1-\mathbb{P}\left(\bigcap_{k=1}^{d}\left\{U_{k}=V_{k}\right\}\right)$
$=1-\prod_{k=1}^{d}\mathbb{P}(U_{k}=V_{k}\mid U_{1:k-1}=V_{1:k-1})$
$=1-\prod_{k=1}^{d}\left[1-\mathbb{P}(U_{k}\neq V_{k}\mid U_{1:k-1}=V_{1:k-1})\right]$
$\leq 1-\prod_{k=1}^{d}\left(1-M\left\lvert x_{k+1:d}-y_{k+1:d}\right\rvert^{\beta}\right)$
$\leq 1-\left(1-M\left\lvert x-y\right\rvert^{\beta}\right)^{d}.$
We now turn our attention towards obtaining a close coupling for the random-scan Gibbs sampler. However, one issue immediately becomes apparent: when $x,y\in\mathbb{R}^{d}$ do not share a coordinate (i.e., when $x-y$ has a non-zero component in each coordinate direction), the supports of $P_{\mathrm{RS}}(x,\cdot)$ and $P_{\mathrm{RS}}(y,\cdot)$ are disjoint. In this case,
$\left\|P_{\mathrm{RS}}(x,\cdot)-P_{\mathrm{RS}}(y,\cdot)\right\|_{\textsc{TV}}=1.$
[…]
Lemma 4.5.
Let $I_{1},\dots,I_{N}\sim\mathrm{Unif}[d]$ be independent and identically distributed. Denote by $\Xi$ the set of unique values attained by $I_{1},\dots,I_{N}$. If $N=\lceil 4d\cdot\log d\rceil$, then
$\mathbb{P}(\Xi=[d])\geq\frac{1}{2}.$
Lemma 4.5 is closely related to the classic coupon collector problem in probability (Erdős and Rényi, 1961). Using the same notation, observe that when $N=d$, we have
$\mathbb{P}(\Xi=[d])=\frac{d!}{d^{d}}\sim\frac{\sqrt{d}}{e^{d}}$
as $d\to\infty$, by Stirling’s formula. The probability of observing all $d$ unique values in exactly $d$ draws thus decreases exponentially with the dimension $d$.
Proposition 4.6.
Suppose that the conditional distributions of $\pi$ are $(M,\beta)$-TV continuous. Let $N_{d}=\lceil 4d\cdot\log d\rceil$. Then,
$\left\|P_{\mathrm{RS}}^{N_{d}}(x,\cdot)-P_{\mathrm{RS}}^{N_{d}}(y,\cdot)\right\|_{\textsc{TV}}\leq 1-\frac{1}{2}(1-M\left\lvert x-y\right\rvert^{\beta})^{N_{d}}\qquad\forall x,y\in\mathbb{R}^{d}.$
Hence, $P_{\mathrm{RS}}^{N_{d}}$ satisfies a $(\delta,\varepsilon)$-close coupling condition with
$\delta=\left(MN_{d}\right)^{-1/\beta},\qquad\varepsilon=\left(2e\right)^{-1}.$
In particular, $\delta\geq(5Md\cdot\log d)^{-1/\beta}$.
Proof.
Similarly to Theorem 4.4, we construct a sequential coupling of $P_{\mathrm{RS}}^{N_{d}}(x,\cdot)$ and $P_{\mathrm{RS}}^{N_{d}}(y,\cdot)$ as follows:
  • Draw $I_{1},\dots,I_{N_{d}}\sim\mathrm{Unif}[d]$ independent and identically distributed.
  • Initialize $U^{\left(0\right)}=x$ and $V^{\left(0\right)}=y$.
  • For $k=1,\dots,d$, draw a maximal coupling
    $W_{k}\sim\pi\left(\cdot\mid U^{\left(k-1\right)}_{-{I_{k}}}\right),\qquad W_{k}^{\prime}\sim\pi\left(\cdot\mid V^{\left(k-1\right)}_{-{I_{k}}}\right),$
    such that
    $\mathbb{P}\left(W_{k}\neq W_{k}^{\prime}\mid U^{\left(k-1\right)}_{-{I_{k}}},V^{\left(k-1\right)}_{-{I_{k}}}\right)=\left\|\pi\left(\cdot\mid U^{\left(k-1\right)}_{-{I_{k}}}\right)-\pi\left(\cdot\mid V^{\left(k-1\right)}_{-{I_{k}}}\right)\right\|_{\textsc{TV}}.$ (12)
    Note that both sides of (12) are random quantities depending on $U^{\left(k-1\right)}$, $V^{\left(k-1\right)},I_{k}$.
  • Update the variables by setting
    $U^{(k)}=\left(U^{(k-1)}_{-I_{k}},W\right),\qquad V^{(k)}=\left(V^{(k-1)}_{-I_{k}},W^{\prime}\right).$
    In words, only the $I_{k}$’th coordinate is updated when transitioning from $U^{(k-1)}$ to $U^{(k)}$ (and from $V^{(k-1)}$ to $V^{(k)}$). The notation used here for indexing $U^{(k)}$ and $V^{(k)}$ follows the convention introduced in (1).
  • Finally, set $U=U^{(N_{d})}$, $V=V^{\left(N_{d}\right)}$.
By construction, $U\sim P_{\mathrm{RS}}^{N_{d}}(x,\cdot)$ and $V\sim P_{\mathrm{RS}}^{N_{d}}(y,\cdot)$. As before, denote by $\Xi$ the set of unique values attained by $I_{1},\dots,I_{N_{d}}$. When $x$ and $y$ differ in at least two coordinates, the event $\left\{U=V\right\}$ has a non-zero probability of occurring only if $\Xi=[d]$, i.e., when all coordinates have been updated at least once by $P_{\mathrm{RS}}^{N_{d}}$. More generally, for all $x,y\in\mathbb{R}^{d}$, we have
$\mathbb{P}(U=V)\geq\mathbb{P}(\{U=V\}\cap\{\Xi=\left[d\right]\})=\mathbb{P}(\Xi=[d])\cdot\mathbb{P}(U=V\mid\Xi=\left[d\right]).$
By Lemma 4.5, $\mathbb{P}(\Xi=\left[d\right])\geq 1/2$ when $N_{d}=\left\lceil 4d\cdot\log d\right\rceil$. Hence, for this value of $N_{d}$, we get
$\left\|P_{\mathrm{RS}}^{N_{d}}(x,\cdot)-P_{\mathrm{RS}}^{N_{d}}(y,\cdot)\right\|_{\textsc{TV}}$ $\leq\mathbb{P}(U\neq V)$
$=1-\mathbb{P}(U=V)$
$\leq 1-\mathbb{P}(\Xi=[d])\cdot\mathbb{P}(U=V\mid\Xi=[d])$
$\leq 1-\frac{1}{2}\cdot\mathbb{P}(U=V\mid\Xi=[d]).$ (13)
By matching each step of the sequential coupling described above, the conditional probability in (13) can be lower bounded by
$\mathbb{P}(U=V\mid\Xi=[d])$ $\geq\mathbb{P}\left(\bigcap_{k=1}^{N_{d}}\left\{W_{k}=W_{k}^{\prime}\right\}~{}\middle|~{}\Xi=\left[d\right]\right)$
$=\prod_{k=1}^{N_{d}}\mathbb{P}\left(W_{k}=W_{k}^{\prime}~{}\middle|~{}\bigcap_{l=1}^{k-1}\left\{W_{l}=W_{l}^{\prime}\right\}\cap\left\{\Xi=\left[d\right]\right\}\right)$
$=\prod_{k=1}^{N_{d}}\left[1-\mathbb{P}\left(W_{k}\neq W_{k}^{\prime}~{}\middle|~{}\bigcap_{l=1}^{k-1}\left\{W_{l}=W_{l}^{\prime}\right\}\cap\left\{\Xi=\left[d\right]\right\}\right)\right].$ (14)
Consider a fixed a sequence of indices $I_{1},\dots,I_{N_{d}}$ and $k\in\left[N_{d}\right]$. Observe that up to the $k$’th step of our sequential coupling, the coordinates of $U^{\left(k\right)}$ that differ from those of $U^{\left(0\right)}=x$ are necessarily equal to some $W_{l}$ with $l\leq k$. Identically, the coordinates of $V^{\left(k\right)}$ that differ from those of $V^{\left(0\right)}=y$ are necessarily equal to some $W_{l}^{\prime}$ with $l\leq k$. Therefore, conditional on the event $\bigcap_{l=1}^{k-1}\left\{W_{l}=W_{l}^{\prime}\right\},$ we have
$\left\lvert U^{\left(k-1\right)}-V^{\left(k-1\right)}\right\rvert\leq\left\lvert x-y\right\rvert.$
As $\left(W_{k},W_{k}^{\prime}\right)$ is a maximal coupling, it follows that
$\mathbb{P}$ $\left(W_{k}\neq W_{k}^{\prime}~{}\middle|~{}\bigcap_{l=1}^{k-1}\left\{W_{l}=W_{l}^{\prime}\right\}\cap\left\{\Xi=\left[d\right]\right\}\right)$
$=\mathbb{E}_{U^{\left(k-1\right)},V^{\left(k-1\right)},I_{k}}\left[\left\|\pi\left(\cdot\mid U^{\left(k-1\right)}_{-{I_{k}}}\right)-\pi\left(\cdot\mid V^{\left(k-1\right)}_{-{I_{k}}}\right)\right\|_{\textsc{TV}}~{}\middle|~{}\bigcap_{l=1}^{k-1}\left\{W_{l}=W_{l}^{\prime}\right\}\cap\left\{\Xi=\left[d\right]\right\}\right]$
$\leq\mathbb{E}_{U^{\left(k-1\right)},V^{\left(k-1\right)},I_{k}}\left[M\left\lvert U^{\left(k-1\right)}_{-I_{k}}-V^{\left(k-1\right)}_{-I_{k}}\right\rvert^{\beta}~{}\middle|~{}\bigcap_{l=1}^{k-1}\left\{W_{l}=W_{l}^{\prime}\right\}\cap\left\{\Xi=\left[d\right]\right\}\right]$
$\leq\mathbb{E}_{U^{\left(k-1\right)},V^{\left(k-1\right)},I_{k}}\left[M\left\lvert U^{\left(k-1\right)}-V^{\left(k-1\right)}\right\rvert^{\beta}~{}\middle|~{}\bigcap_{l=1}^{k-1}\left\{W_{l}=W_{l}^{\prime}\right\}\cap\left\{\Xi=\left[d\right]\right\}\right]$
$\leq M\left\lvert x-y\right\rvert^{\beta},$ (15)
where the subscript in the expectation indicates that it is taken with respect to the joint distribution of $U^{(k-1)},V^{(k-1)},I_{k}$. Combining (13), (14), (15), we finally obtain
$\left\|P_{\mathrm{RS}}^{N_{d}}(x,\cdot)-P_{\mathrm{RS}}^{N_{d}}(y,\cdot)\right\|_{\textsc{TV}}\leq 1-\frac{1}{2}\left(1-M\left\lvert x-y\right\rvert^{\beta}\right)^{N_{d}}.$
When a geometric drift condition holds for a certain Lyapunov function, these close coupling conditions are reminiscent of local Doeblin conditions (Hairer, 2010, Assumption 3.4) restricted to particular level sets of the Lyapunov function; see also (Hairer and Mattingly, 2011, Assumption 2) for an alternative representation in terms of minorization conditions. In this case, one can establish geometric rates of convergence for the Markov process (Hairer, 2010, Theorem 3.6; Hairer and Mattingly, 2011, Theorem 1.2). Nevertheless, a naive application of Propositions 4.4 and 4.6 leads to spectral gap bounds that scale exponentially with the dimension. It is unclear whether this approach can be improved in the present setting.
4.3 Mixing time bounds for the Gibbs sampler
By combining the three-set isoperimetric inequalities from Subsection 4.1 with the close coupling conditions from Subsection 4.2, we can bound the conductance of the systematic-scan and random-scan Gibbs kernels. These results apply to target distributions that satisfy a $L^{q}$-Poincaré or $L^{q}$-log-Sobolev inequality, and whose conditional distributions are TV continuous. Throughout this subsection, we use $\lesssim$ to denote inequalities that hold up to a universal constant factor.
Theorem 4.7.
Consider a target distribution $\pi$ that satisfies $\mathrm{PI}_{q}$ with constant $\mathtt{p}_{q}$. Suppose that the conditional distributions of $\pi$ are $(M,\beta)$-TV continuous. Then, the conductance of the systematic-scan Gibbs kernel is bounded by
$\Phi(P_{\mathrm{SS}})\gtrsim\min\left\{1,\frac{\mathtt{p}_{q}}{\left(Md\right)^{q/\beta}}\right\}.$
If the chain is initialized at a $\Omega$-warm start, then the $\zeta$-mixing time is bounded by
$\tau(\zeta,P_{\mathrm{SS}})\lesssim\max\left\{1,\frac{\left(Md\right)^{2q/\beta}}{\mathtt{p}_{q}^{2}}\right\}\cdot\log\left(\frac{\sqrt{\Omega}}{\zeta}\right).$
Proof.
The bound on the conductance follows from combining Corollary 3.11, Proposition 4.1, and Proposition 4.4. The bound on the mixing time is obtained via Corollary 3.5. ∎
Theorem 4.8.
Consider a target distribution $\pi$ that satisfies $\mathrm{PI}_{q}$ with constant $\mathtt{p}_{q}$. Suppose that the conditional distributions of $\pi$ are $(M,\beta)$-TV continuous. Let $N_{d}=\left\lceil 4d\cdot\log d\right\rceil$. Then, the conductance of the iterated random-scan Gibbs kernel is bounded by
$\Phi\left(P^{N_{d}}_{\mathrm{RS}}\right)\gtrsim\min\left\{1,\frac{\mathtt{p}_{q}}{\left(MN_{d}\right)^{q/\beta}}\right\}.$
The spectral gap satisfies
$\lambda_{2}\left(P^{N_{d}}_{\mathrm{RS}}\right)\gtrsim\min\left\{1,\frac{\mathtt{p}_{q}^{2}}{\left(MN_{d}\right)^{2q/\beta}}\right\}.$
If the chain is initialized at a $\Omega$-warm start, then the $\zeta$-mixing time is bounded by
$\tau\left(\zeta,P^{N_{d}}_{\mathrm{RS}}\right)\lesssim\max\left\{1,\frac{\left(MN_{d}\right)^{2q/\beta}}{\mathtt{p}_{q}^{2}}\right\}\cdot\log\left(\frac{\sqrt{\Omega}}{\zeta}\right).$
Proof.
The conductance bound follows from Corollary 3.11, Proposition 4.1, and Proposition 4.6. The spectral gap is then bounded using Theorem 3.7, while the mixing time bound is obtained via Corollary 3.5. ∎
We now focus on bounding the conductance and spectral gap of the random-scan Gibbs kernel $P_{\mathrm{RS}}$. As noted by Narayanan and Srivastava (2022, Lemma 5.3), for any reversible Markov kernel $P$ and for all $N\geq 1$, we have
$\Phi(P)\geq\frac{1}{N}\cdot\Phi\left(P^{N}\right).$
Let $\sigma(P)$ denote the second largest eigenvalue of $P$, so that $\lambda_{2}(P)=1-\sigma(P)$. Notably, $\sigma\left(P^{N}\right)=\sigma(P)^{N}$, which implies that
$\lambda_{2}(P)$ $=1-\sigma(P)$
$=1-\sigma\left(P^{N}\right)^{1/N}$
$=1-\left(1-\lambda_{2}\left(P^{N}\right)\right)^{1/N}$
$\geq 1-\left(1-\frac{1}{N}\cdot\lambda_{2}\left(P^{N}\right)\right)$
$=\frac{1}{N}\cdot\lambda_{2}\left(P^{N}\right),$
where we have used the inequality $1-t\leq\left(1-t/N\right)^{N}$ for $0\leq t\leq 1$. Regarding mixing times, observe that $\tau(\zeta,P)\leq N\cdot\tau\left(\zeta,P^{N}\right)$.
Corollary 4.9.
Under the conditions of Theorem 4.8, the conductance of the random-scan Gibbs kernel is bounded by
$\Phi(P_{\mathrm{RS}})\gtrsim\frac{1}{d\cdot\log d}\cdot\min\left\{1,\frac{\mathtt{p}_{q}}{M^{q/\beta}(d\cdot\log d)^{q/\beta}}\right\}.$
The spectral gap satisfies
$\lambda_{2}(P_{\mathrm{RS}})\gtrsim\frac{1}{d\cdot\log d}\cdot\min\left\{1,\frac{\mathtt{p}_{q}^{2}}{M^{2q/\beta}(d\cdot\log d)^{2q/\beta}}\right\}.$
The $\zeta$-mixing time is bounded by
$\tau(\zeta,P_{\mathrm{RS}})\lesssim d\cdot\log d\cdot\max\left\{1,\frac{M^{2q/\beta}(d\cdot\log d)^{2q/\beta}}{\mathtt{p}_{q}^{2}}\right\}\cdot\log\left(\frac{\sqrt{\Omega}}{\zeta}\right).$
These results demonstrate that the systematic-scan and random-scan Gibbs samplers exhibit polynomial mixing time dependence on the dimension $d$ for a large class of target distributions. Note that a single step of the systematic-scan kernel $P_{\mathrm{SS}}$ updates all $d$ coordinates, whereas the random-scan kernel $P_{\mathrm{RS}}$ only updates one. Even after accounting for this difference, a residual $\log d$ factor remains in the mixing time bounds for random-scan. This additional $\log d$ factor highlights a modest penalty compared to systematic scan for well-behaved target distributions. This aligns with prior heuristic observations about the efficiency of these sampling schemes (Diaconis, 2013) and arises naturally from spectral bounds on mixing times (Gaitonde and Mossel, 2024).
By Proposition 4.3, all results in this subsection extend to distributions satisfying $\mathrm{LSI}_{q}$ instead of $\mathrm{PI}_{q}$, with the conductance bounds remaining equivalent up to constant factors. Consequently, the bounds on the mixing time in total variation distance remain unchanged when replacing a Poincaré inequality with a log-Sobolev inequality of the same order. However, improvements may be possible when using a different norm to define the mixing time. For instance, in $\chi^{2}$-divergence, Chen et al. (2020, Theorem 1) demonstrated that the logarithmic factor in $\Psi$ from Proposition 4.3 can be exploited to improve the mixing time’s dependence on the warm start parameter (see Corollary 3.5), reducing it from $\log$ to $\log\log$.
5 Extensions and applications
5.1 Implications for Lipschitz and smooth potentials
We now examine the regularity assumption on the conditional distributions of the target $\pi$ introduced in Definition 2.5. The density of $\pi$ with respect to the Lebesgue measure can be written as
$\pi(\mathrm{d}x)\propto e^{-U\left(x\right)}\mathrm{d}x,$
where the function $U:\mathbb{R}^{d}\to\mathbb{R}$ is known as the potential. Many previous works assume that the potential function is either Lipschitz continuous (Lehec, 2023) or smooth (see, e.g., Ascolani et al., 2024a). We show that TV continuity of the conditional distributions of $\pi$ (Definition 2.5) is a strictly weaker condition than these regularity assumptions on $U$.
Definition 5.1.
Let $L>0$. The potential $U$ is said to be $L$-Lipschitz continuous if
$\left\lvert U(x)-U(y)\right\rvert\leq L\left\lvert x-y\right\rvert\qquad\forall x,y\in\mathbb{R}^{d}.$
Equivalently, the distribution $\pi$ is said to be $L$-log-Lipschitz continuous.
Definition 5.2.
Let $L>0$. The potential $U$ is said to be $L$-smooth if $U$ is continuously differentiable with
$\left\lvert\nabla U(x)-\nabla U(y)\right\rvert\leq L\left\lvert x-y\right\rvert\qquad\forall x,y\in\mathbb{R}^{d}.$
Equivalently, the distribution $\pi$ is said to be $L$-log-smooth.
To establish our result, we first demonstrate that log-Lipschitz continuity and log-smoothness extend to the marginal distributions of $\pi$. The proof of the following lemma is located in Appendix A.
Lemma 5.3.
The following properties hold:
  • If the distribution $\pi$ is $L$-log-Lipschitz continuous, then all marginal distributions of $\pi$ are also $L$-log-Lipschitz continuous.
  • If the distribution $\pi$ is $L$-log-smooth, then all marginal distributions of $\pi$ are also $L$-log-smooth.
Lemma 5.3 complements Saumard and Wellner (2014, Theorems 3.3 and 3.8), who proved that (strong) log-concavity is also inherited by the marginal distributions of $\pi$.
Lemma 5.4.
Suppose that the distribution $\pi$ is $L$-log-Lipschitz continuous. Then, the conditional distributions of $\pi$ are $(\sqrt{L},1/2)$-TV continuous.
Proof.
For simplicity, we redefine the variables as $x_{-k}=s$, $y_{-k}=t$. The joint distribution can be rewritten as $\pi(\theta,s)\propto e^{-U(\theta,s)}.$ From this, the conditional distribution is expressed as
$\pi(\theta|s)=\frac{\pi(\theta,s)}{\pi(s)}=\frac{e^{-U(\theta,s)}}{\int e^{-U(\psi,s)}\mathrm{d}\psi}=\frac{e^{-U(\theta,s)}}{e^{-V(s)}},$
where
$V(s)=-\log\left(\int e^{-U(\psi,s)}\mathrm{d}\psi\right).$
By Lemma 5.3, the function $V$ is $L$-Lipschitz continuous. We now compute the KL-divergence between $\pi(\cdot\mid s)$ and $\pi(\cdot\mid t)$:
$\mathrm{KL}(\pi(\cdot\mid s),\pi(\cdot\mid t))$ $=\int\log\left(\frac{\pi(\theta\mid s)}{\pi(\theta\mid t)}\right)\pi(\mathrm{d}\theta\mid s)$
$=\int\log\pi(\theta\mid s)\pi(\mathrm{d}\theta\mid s)-\int\log\pi(\theta\mid t)\pi(\mathrm{d}\theta\mid s)$
$=-\int U(\theta,s)\pi(\mathrm{d}\theta\mid s)+V(s)+\int U(\theta,t)\pi(\mathrm{d}\theta\mid s)-V(t)$
$\leq\int\left\lvert U(\theta,s)-U(\theta,t)\right\rvert\pi(\mathrm{d}\theta\mid s)+\left\lvert V(s)-V(t)\right\rvert$
$\leq 2L\left\lvert s-t\right\rvert.$
We conclude by applying Pinsker’s inequality:
$\left\|\pi(\cdot\mid s)-\pi(\cdot\mid t)\right\|_{\textsc{TV}}$ $\leq\sqrt{\frac{1}{2}\cdot\mathrm{KL}(\pi(\cdot\mid s),\pi(\cdot\mid t))}$
$\leq\sqrt{L\left\lvert s-t\right\rvert}.$
Lemma 5.5.
Suppose that the distribution $\pi$ is $L$-log-smooth. Then, the conditional distributions of $\pi$ are $(\sqrt{L},1)$-TV continuous.
Proof.
We proceed similarly to Lemma 5.4. Using the same notation, recall that
$\mathrm{KL}(\pi(\cdot\mid s),\pi(\cdot\mid t))=\int(U(\theta,t)-U(\theta,s))\pi(\mathrm{d}\theta\mid s)+V(s)-V(t).$
Observe that $\nabla V(s)=\int\nabla_{s}U(\theta,s)\pi(\mathrm{d}\theta\mid s).$ It follows that
$\mathrm{KL}(\pi(\cdot\mid s),\pi(\cdot\mid t))$ $=\int(U(\theta,t)-U(\theta,s))\pi(\mathrm{d}\theta\mid s)+V(s)-V(t)$
$\leq\int\left(\nabla_{s}U(\theta,s)^{T}(t-s)+\frac{L}{2}\left\lvert s-t\right\rvert^{2}\right)\pi(\mathrm{d}\theta\mid s)$
$\quad+\nabla V(t)^{T}(s-t)+\frac{L}{2}\left\lvert s-t\right\rvert^{2}$
$\leq\int(\nabla_{s}U(\theta,s)-\nabla_{t}U(\theta,t))^{T}(t-s)\pi(\mathrm{d}\theta\mid s)+L\left\lvert s-t\right\rvert^{2}$
$\leq 2L\left\lvert s-t\right\rvert^{2}.$
We conclude by applying Pinsker’s inequality:
$\left\|\pi(\cdot\mid s)-\pi(\cdot\mid t)\right\|_{\textsc{TV}}$ $\leq\sqrt{\frac{1}{2}\cdot\mathrm{KL}(\pi(\cdot\mid s),\pi(\cdot\mid t))}$
$\leq\sqrt{L}\left\lvert s-t\right\rvert.$
The class of log-smooth and (strongly) log-concave target distributions is widely studied in the analysis of MCMC mixing times; see Ascolani et al. (2024a); Wadia (2024) for results regarding the Gibbs sampler. For comparison, we now establish conductance bounds for the systematic-scan and random-scan Gibbs kernels when the target distribution is log-smooth and satisfies $\mathrm{PI}_{1}$. These results follow directly from Theorem 4.7 and Corollary 4.9, combined with Lemma 5.5.
Corollary 5.6.
Consider a $L$-log-smooth target distribution $\pi$ that satisfies $\mathrm{PI}_{1}$. Then, the conductance of the systematic-scan Gibbs kernel is bounded by
$\Phi(P_{\mathrm{SS}})\gtrsim\min\left\{1,\frac{\mathtt{p}_{1}}{\sqrt{L}d}\right\}.$
If the chain is initialized at a $\Omega$-warm start, then the $\zeta$-mixing time is bounded by
$\tau(\zeta,P_{\mathrm{SS}})\lesssim\max\left\{1,\frac{Ld^{2}}{\mathtt{p}_{1}^{2}}\right\}\cdot\log\left(\frac{\sqrt{\Omega}}{\zeta}\right).$
However, the mixing time bound in Corollary 4.7 appears to be loose when $\pi$ is a Gaussian distribution, as suggested by comparisons with Roberts and Sahu (1997).
Corollary 5.7.
Consider a $L$-log-smooth target distribution $\pi$ that satisfies $\mathrm{PI}_{1}$. Then, the conductance of the random-scan Gibbs kernel is bounded by
$\Phi(P_{\mathrm{RS}})\gtrsim\frac{1}{d\cdot\log d}\cdot\min\left\{1,\frac{\mathtt{p}_{1}}{\sqrt{L}d\cdot\log d}\right\}.$
The spectral gap satisfies
$\lambda_{2}(P_{\mathrm{RS}})\gtrsim\frac{1}{d\cdot\log d}\cdot\min\left\{1,\frac{\mathtt{p}_{1}^{2}}{L(d\cdot\log d)^{2}}\right\}.$
If the chain is initialized at a $\Omega$-warm start, then the $\zeta$-mixing time is bounded by
$\tau(\zeta,P_{\mathrm{RS}})\lesssim d\cdot\log d\cdot\max\left\{1,\frac{L(d\cdot\log d)^{2}}{\mathtt{p}_{1}^{2}}\right\}\cdot\log\left(\frac{\sqrt{\Omega}}{\zeta}\right).$
Unfortunately, the spectral gap bound in Corollary 4.9 is suboptimal when $\pi$ is strongly log-concave and $L$-log-smooth. In such cases, Ascolani et al. (2024a, Corollary 3.6) established that
$\lambda_{2}(P_{\mathrm{RS}})\gtrsim\frac{\mathtt{p}_{1}^{2}}{Ld}.$ (16)
Moreover, Amit (1996, Theorem 1) demonstrated that the bound in (16) is tight when $\pi$ is a Gaussian distribution. The sharp bounds obtained in Ascolani et al. (2024a) are based on contraction arguments in KL-divergence and critically depend on the assumption of strong log-concavity. In particular, it is unclear whether these bounds even extend to bounded perturbations as in Holley and Stroock (1987). In contrast, our approach does not exploit such strong structural assumptions, which limits its capacity to achieve comparably sharp bounds. However, this generality allows it to apply to a broader class of target distributions, including those that are not strongly log-concave.
Furthermore, we believe that the lower bound on the spectral gap in Theorem 3.7 is not tight for the Gibbs sampler. This belief is supported by the case of the random walk on the $d$-dimensional hypercube, where Trevisan (2013, Chapter 5) showed that the conductance and spectral gap are of the same order, making the first bound from Theorem 3.7 loose in this setting. The Gibbs sampler shares a similar local structure to the $d$-dimensional hypercube: each update resamples a single coordinate from its exact conditional, in analogy to the single-bit flips in the hypercube walk.
5.2 Non-log-concave examples
Here, we present examples of $d$-dimensional target distributions where our methods yield explicit mixing time bounds for the Gibbs sampler, in contrast to those of Ascolani et al. (2024a); Wadia (2024).
Continuous Gaussian mixture.
For all $t\in\mathbb{R}^{d}$, let $\varphi_{t}(\mathrm{d}x)\propto\exp(-|x-t|^{2}/2)\mathrm{d}x$ denote a Gaussian distribution with mean $t$ and covariance matrix $I_{d}$. Define the mixing distribution
$\nu(\mathrm{d}t)\propto e^{-\left\lvert\left\lvert t\right\rvert^{2}-4\right\rvert}\mathrm{d}t.$
The resulting Gaussian mixture
$\pi_{1}(\mathrm{d}x)=\int\varphi_{t}(\mathrm{d}x)\nu(\mathrm{d}t)$
is log-smooth and satisfies $\mathrm{LSI}_{2}$ (Chen et al., 2021, Subsection 4.2). However, since $\nu$ is multimodal, $\pi_{1}$ is not log-concave.
Discrete Laplace mixture.
Let $m_{1},m_{2}\in\mathbb{R}^{d}$ with $m_{1}=-m_{2}=(1,1,\dots,1)$. Consider the two-component mixture of Laplace distributions
$\pi_{2}(\mathrm{d}x)\propto\left(\frac{1}{2}\cdot e^{-\left\lvert x-m_{1}\right\rvert}+\frac{1}{2}\cdot e^{-\left\lvert x-m_{2}\right\rvert}\right)\mathrm{d}x.$
This distribution is log-Lipschitz continuous, and satisfies $\mathrm{PI}_{2}$ as shown by Chafaï and Malrieu (2010); Schlichting (2019). However, $\pi_{2}$ is not log-concave.
Perturbed Laplace distribution.
Consider the distribution
$\pi_{3}(\mathrm{d}x)\propto e^{-\left\lvert x\right\rvert+2\cos\left\lvert x\right\rvert}\mathrm{d}x.$
This distribution is log-Lipschitz continuous and satisfies $\mathrm{PI}_{1}$ (Bobkov and Houdré, 1997). However, the added perturbation prevents $\pi_{3}$ from being log-concave.
6 Conclusion
We established verifiable conductance bounds for systematic-scan and random-scan Gibbs samplers over a broad class of target distributions, particularly those satisfying a Poincaré or log-Sobolev inequality and possessing sufficiently regular conditional distributions (Theorem 4.7 and Corollary 4.9). These bounds directly provide estimates on the mixing time in total variation distance and remain valid for either log-Lipschitz or log-smooth target distributions. In doing so, we extend prior results beyond the log-concave setting (Ascolani et al., 2024a; Wadia, 2024) and present several examples.
Our work also complements recent results showing that the spectral gaps and mixing times of systematic-scan and random-scan Gibbs samplers are closely related (Gaitonde and Mossel, 2024; Chlebicka et al., 2025). While our analysis focuses on coordinate-wise updates, our results naturally extend to blocked Gibbs samplers under corresponding block-smoothness conditions, allowing for joint updates of coordinate subsets. Moreover, our work also has implications for Metropolis-within-Gibbs sampling schemes; Ascolani et al. (2024b) examine how the conductance of generic coordinate-wise schemes compares to that of exact Gibbs sampling, while Qin et al. (2025) perform a similar analysis for the spectral gap. Their results, combined with our bounds, provide a framework for studying the convergence of broader coordinate-wise sampling methods.
Regarding dimension dependence, it is clear that our bounds on the spectral gap and mixing time are not tight, particularly for Gaussian targets (Amit, 1996; Roberts and Sahu, 1997) and, more generally, for log-smooth and strongly log-concave distributions (Ascolani et al., 2024a). We attribute this to the generality of our assumptions, which might not fully exploit the underlying structure of these distributions, and to the use of total variation distance in defining the mixing times. A natural direction for future research is therefore to refine these conductance bounds for both variants of the Gibbs sampler, particularly in terms of their dependence on the dimension and warmness parameter, as studied in Chen et al. (2020). One possibility is to study the mixing time of the Gibbs sampler under alternative probability metrics, such as appropriate Wasserstein distances, as in Wang and Wu (2014). Beyond the conductance approach, one could also explore recent techniques that may yield sharper spectral gap estimates for the Gibbs sampler, such as quasi-telescoping inequalities (also known as the spectral telescope), as developed in Qin and Wang (2024).
Acknowledgments
Alexander Goyal is fully funded by the Roth scholarship. George Deligiannidis is supported by the Engineering and Physical Sciences Research Council [grant number EP/Y018273/1].
Appendix A Appendix
Proof of Proposition 2.4.
Define the function $N_{q}(x)=\left\lvert x\right\rvert^{q}\cdot\log(1+\left\lvert x\right\rvert^{q})$ for all $x\in\mathbb{R}^{d}$. Given a function $f:\mathbb{R}^{d}\to\mathbb{R}$, we denote its Orlicz norm associated to $N_{q}$ by
$\left\|f\right\|_{N_{q}}=\inf\left\{\lambda>0:\int N_{q}\left(\frac{f(x)}{\lambda}\right)\pi(\mathrm{d}x)\leq 1\right\}.$
Let $\left\|f\right\|_{q}$ also denote the $L^{q}_{\pi}$-norm of $f$, as defined in Subsection 4.1. We now introduce two functional inequalities in the Orlicz space. Let $\mathtt{c}_{q}>0$ and $\tilde{\mathtt{c}}_{q}>0$ denote the largest constants such that, for all locally Lipschitz functions $f:\mathbb{R}^{d}\to\mathbb{R}$,
$\mathtt{c}_{q}\left\|f-\mathbb{E}[f]\right\|_{N_{q}}^{q}$ $\leq\mathbb{E}\left[\left\lvert\nabla f\right\rvert^{q}\right],$ (17)
$\tilde{\mathtt{c}}_{q}\left\|f-m[f]\right\|_{N_{q}}^{q}$ $\leq\mathbb{E}\left[\left\lvert\nabla f\right\rvert^{q}\right],$ (18)
where $m[f]$ denotes a median of $f$. Bobkov and Zegarlinski (2005, Proposition 3.1) showed that $\mathrm{LSI}_{q}$ is equivalent to the functional inequality (17), with constants satisfying $\mathtt{c}_{q}/16\leq\mathtt{s}_{q}\leq 7\mathtt{c}_{q}$. Moreover, by Milman (2009a, Lemma 2.1), we have $\tilde{\mathtt{c}}_{q}/2\leq\mathtt{c}_{q}\leq 3\tilde{\mathtt{c}}_{q}$. The functional inequality (18) is therefore equivalent to $\mathrm{LSI}_{q}$, with constants satisfying
$\frac{\tilde{\mathtt{c}}_{q}}{32}\leq\mathtt{s}_{q}\leq 21\tilde{\mathtt{c}}_{q}.$ (19)
Now, take any locally Lipschitz function $f$ with $m[f]=0$ and define $g=\mathrm{sgn}(f)\cdot\left\lvert f\right\rvert^{r/q}$, where $\mathrm{sgn}(\cdot)$ denotes the sign function. In particular, we have $m[g]=0$ as well. Applying (18) to $g$ yields
${\tilde{\mathtt{c}}_{q}}^{1/q}\left\|g\right\|_{N_{q}}\leq\left\|\left\lvert\nabla g\right\rvert\right\|_{q}.$ (20)
Since the solution to the equation $1/u+1/r=1/q$ is $u=qr/(r-q)$, an application of Hölder’s inequality gives
$\left\|\left\lvert\nabla g\right\rvert\right\|_{q}$ $\leq\frac{r}{q}\left\|\left\lvert f\right\rvert^{r/q-1}\cdot\left\lvert\nabla f\right\rvert\right\|_{q}$
$=\frac{r}{q}\left\|\left\lvert f\right\rvert^{r/q-1}\right\|_{qr/(r-q)}\cdot\left\|\left\lvert\nabla f\right\rvert\right\|_{r}$
$=\frac{r}{q}\left\|f\right\|_{r}^{r/q-1}\cdot\left\|\left\lvert\nabla f\right\rvert\right\|_{r}$
$\leq\frac{r}{q}\left(\frac{5}{4}\right)^{1/q-1/r}\left\|f\right\|_{N_{r}}^{r/q-1}\cdot\left\|\left\lvert\nabla f\right\rvert\right\|_{r}.$
To justify the transition from the second to the third line, observe that
$\left(\frac{r}{q}-1\right)\frac{qr}{r-q}=\frac{r^{2}-qr}{r-q}=r$
and
$\frac{r-q}{qr}=\frac{1}{r}\left(\frac{r}{q}-1\right).$
The last inequality follows from Bobkov and Zegarlinski (2005, Lemma 3.2). Furthermore, for all $\lambda>0$, we have
$N_{q}\left(\frac{g}{\lambda}\right)$ $=\frac{\left\lvert f\right\rvert^{r}}{\lambda^{q}}\cdot\log\left(1+\frac{\left\lvert f\right\rvert^{r}}{\lambda^{q}}\right)$
$=\left\lvert\frac{f}{\lambda^{q/r}}\right\rvert^{r}\cdot\log\left(1+\left\lvert\frac{f}{\lambda^{q/r}}\right\rvert^{r}\right)$
$=N_{r}\left(\frac{f}{\lambda^{q/r}}\right).$
It follows that
$\left\|g\right\|_{N_{q}}$ $=\inf\left\{\lambda>0:\int N_{q}\left(\frac{g(x)}{\lambda}\right)\pi(\mathrm{d}x)\leq 1\right\}$
$=\inf\left\{\lambda>0:\int N_{r}\left(\frac{f(x)}{\lambda^{q/r}}\right)\pi(\mathrm{d}x)\leq 1\right\}$
$=\inf\left\{\lambda^{q/r}>0:\int N_{r}\left(\frac{f(x)}{\lambda^{q/r}}\right)\pi(\mathrm{d}x)\leq 1\right\}^{r/q}$
$=\inf\left\{\lambda^{\prime}>0:\int N_{r}\left(\frac{f(x)}{\lambda^{\prime}}\right)\pi(\mathrm{d}x)\leq 1\right\}^{r/q}$
$=\left\|f\right\|_{N_{r}}^{r/q}.$
Inserting these identities into (20) gives
${\tilde{\mathtt{c}}_{q}}^{1/q}\left\|f\right\|_{N_{r}}^{r/q}\leq\left(\frac{5}{4}\right)^{1/q-1/r}\frac{r}{q}\left\|f\right\|_{N_{r}}^{r/q-1}\cdot\left\|\left\lvert\nabla f\right\rvert\right\|_{r}.$
Rearranging terms and raising both sides to the power $r$, we obtain
$\left(\frac{4}{5}\right)^{r/q-1}\left(\frac{q}{r}\right)^{r}{\tilde{\mathtt{c}}_{q}}^{r/q}\left\|f\right\|_{N_{r}}^{r}\leq\mathbb{E}\left[\left\lvert\nabla f\right\rvert^{r}\right].$
Hence,
$\left(\frac{4}{5}\right)^{r/q-1}\left(\frac{q}{r}\right)^{r}{\tilde{\mathtt{c}}_{q}}^{r/q}\leq\tilde{\mathtt{c}}_{r}.$
Finally, applying (19) twice yields
$\frac{5}{128}\left(\frac{4}{105}\right)^{r/q}\left(\frac{q}{r}\right)^{r}\mathtt{s}_{q}^{r/q}\leq\mathtt{s}_{r}.$
Proof of Proposition 3.10.
Let $S$ $\subseteq\mathbb{R}^{d}$ with $\pi(S)\leq 1/2$. Note that in this case, $\pi(S^{c})\geq\pi(S)$. Define the sets
$S_{1}$ =$=\left\{x\in S:P(x,S^{c})\leq\frac{\varepsilon}{2}\right\},$
$S_{2}$ =$=\left\{x\in S^{c}:P(x,S)\leq\frac{\varepsilon}{2}\right\},$
$S_{3}$ $=\left(S_{1}\sqcup S_{2}\right)^{c}.$
Here, $S_{1}$ and $S_{2}$ can be viewed as “bad” sets: in $S_{1}$ (respectively $S_{2}$), the probability to escape $S$ (respectively $S^{c}$) is small. Note that $S_{1}\sqcup S_{2}\sqcup S_{3}=\mathbb{R}^{d}$, forming a partition of space. We now separate two different cases.
In the first case, suppose that $\pi(S_{1})\leq\pi(S)/2$ or $\pi(S_{2})\leq\pi(S^{c})/2$, i.e., at least one of the “bad” sets is small. If $\pi(S_{1})\leq\pi(S)/2$, then
$\int_{S}P(x,S^{c})\pi(\mathrm{d}x)$ $\geq\int_{S\setminus S_{1}}P(x,S^{c})\pi(\mathrm{d}x)$
$\geq\frac{\varepsilon}{2}\cdot\pi(S\setminus S_{1})$
$=\frac{\varepsilon}{2}\left[\pi(S)-\pi(S_{1})\right]$
$\geq\frac{\varepsilon}{4}\cdot\pi(S).$
If $\pi(S_{2})\leq\pi(S^{c})/2$, we proceed analogously:
$\int_{S}P(x,S^{c})\pi(\mathrm{d}x)$ $=\int_{S^{c}}P(x,S)\pi(\mathrm{d}x)$
$\geq\int_{S^{c}\setminus S_{2}}P(x,S^{c})\pi(\mathrm{d}x)$
$\geq\frac{\varepsilon}{2}\cdot\pi(S^{c}\setminus S_{2})$
$=\frac{\varepsilon}{2}\left[\pi(S^{c})-\pi(S_{2})\right]$
$\geq\frac{\varepsilon}{4}\cdot\pi(S^{c})$
$\geq\frac{\varepsilon}{4}\cdot\pi(S).$
Importantly, reversibility is not required to establish the first equality, as noted by Dwivedi et al. (2019, Lemma 2).
In the second case, suppose that $\pi(S_{1})>\pi(S)/2$ and $\pi(S_{2})>\pi(S^{c})/2$, i.e., both of the “bad” sets are large. Take $x\in S_{1}$, $y\in S_{2}$. We have
$\left\|P(x,\cdot)-P(y,\cdot)\right\|_{\textsc{TV}}\geq P(x,S)-P(y,S)=1-P(x,S^{c})-P(y,S)\geq 1-\varepsilon.$
By the contrapositive of the $(\delta,\varepsilon)$-close coupling condition, this implies $\left\lvert x-y\right\rvert>\delta$, and consequently, $\mathcal{D}(S_{1},S_{2})\geq\delta$. This leads to
$\int_{S}P(x,S^{c})\pi(\mathrm{d}x)$ $=\frac{1}{2}\left[\int_{S}P(x,S^{c})\pi(\mathrm{d}x)+\int_{S^{c}}P(x,S)\pi(\mathrm{d}x)\right]$
$\geq\frac{1}{2}\left[\int_{S\setminus S_{1}}P(x,S^{c})\pi(\mathrm{d}x)+\int_{S^{c}\setminus S_{2}}P(x,S)\pi(\mathrm{d}x)\right]$
$\geq\frac{\varepsilon}{4}\left[\pi(S\setminus S_{1})+\pi(S^{c}\setminus S_{2})\right]$
$=\frac{\varepsilon}{4}\cdot\pi(S_{3})$
$\geq\frac{\varepsilon}{4}\cdot\Upsilon\left(\mathcal{D}(S_{1},S_{2})\right)\cdot\Psi\left(\min\left\{\pi(S_{1}),\pi(S_{2})\right\}\right)$
$\geq\frac{\varepsilon}{4}\cdot\Upsilon(\delta)\cdot\Psi\left(\frac{\pi(S)}{2}\right).$
The statement follows by taking minimums of both cases. ∎
Proof of Lemma 4.2.
The derivative of $\rho$ is given by $\rho^{\prime}(x)=\log(1/x)-1$, which is positive for $x<e^{-1}$ and negative for $x\geq e^{-1}$. This establishes the first part of the lemma. Furthermore, for $e^{-1}\leq x\leq 1$, observe that $\rho^{\prime}(x)\in[-1,0]$, implying that $\rho$ is Lipschitz continuous on this interval with constant 1.
To prove the second part, we consider the following cases. First, if $x+h<e^{-1}$, then both $x$ and $x+h$ lie in the region where $\rho$ is increasing. Therefore, $\rho(x+h)\geq\rho(x)\geq\rho(x)-h$, as required. Next, consider the case where $x+h\geq e^{-1}$. If $x\geq e^{-1}$, since $\rho$ is decreasing and Lipschitz continuous on this region, we have as desired
$\rho(x)-\rho(x+h)=\left\lvert\rho(x)-\rho(x+h)\right\rvert\leq h.$
If $x<e^{-1}$, we split the integral representation of $\rho(x+h)-\rho(x)$ as follows:
$\rho(x+h)$ $=\rho(x)+\int_{x}^{e^{-1}}\rho^{\prime}(s)\mathrm{d}s+\int_{e^{-1}}^{x+h}\rho^{\prime}(s)\mathrm{d}s$
$\geq\rho(x)+\int_{e^{-1}}^{x+h}\rho^{\prime}(s)\mathrm{d}s$
$\geq\rho(x)-\left(x+h-e^{-1}\right)$
$\geq\rho(x)-h.$
This establishes the second part of the lemma. ∎
Proof of Lemma 4.5.
Let $T$ be the random variable representing the number of i.i.d. draws from the distribution $\mathrm{Unif}[d]$ required to observe all $d$ unique values. Following Mitzenmacher and Upfal (2017, Subsection 2.4.1), we have
$\mathbb{E}[T]=d\sum_{j=1}^{d}\frac{1}{j}.$
Approximating the above sum with the logarithm, we have
$d\cdot\log d\leq\mathbb{E}[T]\leq d\cdot\log d+d.$
An application of Markov’s inequality finally yields
$\mathbb{P}(\Xi=[d])$ $=\mathbb{P}(T\leq N)$
$\geq\mathbb{P}(T\leq 4d\cdot\log d)$
$\geq\mathbb{P}(T\leq 2(d\cdot\log d+d))$
$=1-\mathbb{P}(T>2(d\cdot\log d+d))$
$\geq 1-\frac{\mathbb{E}[T]}{2(d\cdot\log d+d)}$
$\geq\frac{1}{2}.$
Proof of Lemma 5.3.
Consider a partition of the space such that $\pi=\pi\left(\theta,s\right)$ with $\theta\in\mathbb{R}^{d_{1}}$, $s\in\mathbb{R}^{d_{2}}$, where $d_{1}+d_{2}=d$. Write $\pi\left(\theta,s\right)\propto e^{-U\left(\theta,s\right)}$. The marginal distribution of $\pi$ with respect to $s$ can be expressed as $\pi\left(s\right)\propto e^{-V\left(s\right)}$, where
$V(s)=-\log\left(\int e^{-U(\theta,s)}\mathrm{d}\theta\right).$
Additionally, the conditional distribution of $\pi$ is
$\pi(\mathrm{d}\theta\mid s)=\frac{e^{-U(\theta,s)}\mathrm{d}\theta}{\int e^{-U(\eta,s)}\mathrm{d}\eta}.$
If $U$ is Lipschitz continuous, then the gradient $\nabla U$ exists almost everywhere and satisfies $\left\lvert U(\theta,s)\right\rvert\leq L$. The gradient of $V$ can be written as
$\nabla V(s)=\int\nabla_{s}U(\theta,s)\pi(\mathrm{d}\theta\mid s),$
where $\nabla_{s}U(\theta,s)$ represents the gradient of $U(\theta,s)$ with respect to the coordinates of $s$. Therefore,
$\left\lvert\nabla V(s)\right\rvert=\int\left\lvert\nabla_{s}U(\theta,s)\right\rvert\pi(\mathrm{d}\theta\mid s)\leq L$
almost everywhere, which shows that $V$ is also $L$-Lipschitz continuous.
Now suppose that $U$ is $L$-smooth, meaning that the Hessian $\nabla^{2}U$ exists almost everywhere and satisfies $\nabla^{2}U(\theta,s)\preceq LI_{d}$, where $I_{d}$ denotes the $d$-dimensional identity matrix. Writing $w^{2}=ww^{T}$ for a vector $w\in\mathbb{R}^{d_{2}}$, the Hessian of $V$ is
$\nabla^{2}V(s)=\int\nabla^{2}U(\theta,s)\pi(\mathrm{d}\theta\mid s)+\left(\int\nabla_{s}U(\theta,s)\pi(\mathrm{d}\theta\mid s)\right)^{2}-\int\left(\nabla_{s}U(\theta,s)\right)^{2}\pi(\mathrm{d}\theta\mid s).$
By the Cauchy-Schwarz inequality,
$\left(\int\nabla_{s}U(\theta,s)\pi(\mathrm{d}\theta\mid s)\right)^{2}-\int\left(\nabla_{s}U(\theta,s)\right)^{2}\pi(\mathrm{d}\theta\mid s)\preceq 0.$
Thus,
$\nabla^{2}V(s)\preceq\int\nabla^{2}U(\theta,s)\pi(\mathrm{d}\theta\mid s)\preceq LI_{d_{2}},$
implying that $V$ is also $L$-smooth. ∎

### v2

@@
 Moreover, we demonstrate that our results remain valid for log-Lipschitz and log-smooth target distributions.
-Our approach relies on novel three-set isoperimetric inequalities and a sequential coupling argument for the Gibbs sampler.
+Our approach relies on novel isoperimetric inequalities and a sequential coupling argument for the Gibbs sampler.
 1 Introduction
@@
 Markov chain Monte Carlo (MCMC) methods are among the most popular algorithms for generating samples from a potentially high-dimensional target distribution $\pi$.
-The idea is to simulate an ergodic Markov chain that admits $\pi$ as its stationary distribution, and use the states of the chain as approximate samples from $\pi$.
+The idea is to simulate an ergodic Markov chain that admits $\pi$ as its invariant distribution, and use the states of the chain as approximate samples from $\pi$.
 The Gibbs sampler (Geman and Geman, 1984) is one of the most widely used MCMC methods in practical applications.
@@
 For such models, Ascolani and Zanella (2024) established further bounds on the convergence rate of the Gibbs sampler, leveraging Bayesian asymptotics under specific random data-generating assumptions.
-In this paper, we characterize the convergence of Gibbs sampling using the so-called mixing time, which quantifies the number of iterations needed for the Markov chain to be within a specified distance of its stationary distribution.
+In this paper, we characterize the convergence of Gibbs sampling using the so-called mixing time, which quantifies the number of iterations needed for the Markov chain to be within a specified distance of its invariant distribution.
 A common technique for bounding the mixing time is to analyze a quantity known as the conductance of the Markov chain.
@@
 This is often formalized via Cheeger’s isoperimetric inequality (Cheeger, 1970), which holds for log-concave target distributions (Bobkov, 1999) and remains valid under bounded perturbations, as characterized by the criterion of Holley and Stroock (1987).
-The conductance approach has proven successful for the Hit-and-Run algorithm (Lovász and Simonovits, 1993; Lovász, 1999; Lovász and Vempala, 2004; Vempala, 2005), as well as for random-walk Metropolis-type algorithms (Dwivedi et al., 2019; Mangoubi and Vishnoi, 2019; Chen et al., 2020; Chewi et al., 2021; Zou et al., 2021; Wu et al., 2022; Mou et al., 2022; Chen and Gatmiry, 2023), among others.
-However, conductance-based proof techniques have seen limited use in the context of Gibbs sampling, with notable exceptions including sampling from uniform distributions over convex bodies (Narayanan and Srivastava, 2022; Laddha and Vempala, 2023), posterior distributions in Bayesian regression models (Lee and Zhang, 2024), and strongly log-concave targets (Wadia, 2024).
+The conductance approach has proven successful for the Hit-and-Run algorithm (Lovász and Simonovits, 1993; Lovász, 1999; Lovász and Vempala, 2004; Vempala, 2005), as well as for random-walk Metropolis-type algorithms (Dwivedi et al., 2019; Mangoubi and Vishnoi, 2019; Chen et al., 2020; Chewi et al., 2021; Zou et al., 2021; Wu et al., 2022; Mou et al., 2022a; Chen and Gatmiry, 2023), among others.
+However, conductance-based proof techniques have seen limited use in the context of Gibbs sampling, with notable exceptions including sampling from uniform distributions over convex bodies (Narayanan and Srivastava, 2022; Laddha and Vempala, 2023; Narayanan et al., 2024), posterior distributions in Bayesian regression models (Lee and Zhang, 2024), and strongly log-concave targets (Wadia, 2024).
 In the recent seminal work by Andrieu et al.
@@
 Nevertheless, when deriving explicit bounds on the mixing time of the random-walk Metropolis algorithm, even these authors exclusively focus on the strongly log-concave case.
-To the best of our knowledge, the work of Vempala and Wibisono (2019) on the unadjusted Langevin algorithm is the only existing result that establishes convergence bounds under a Poincaré or log-Sobolev inequality rather than relying on Cheeger’s isoperimetric inequality.
+For several other classes of sampling algorithms, convergence bounds under a Poincaré or log-Sobolev inequality have been established without relying on Cheeger’s isoperimetric inequality.
+These include discretizations of the overdamped Langevin diffusion (Altschuler and Chewi, 2024b; Chewi et al., 2025; Mou et al., 2022b; Balasubramanian et al., 2022; Kandasamy and Nagaraj, 2024; Mitra and Wibisono, 2025; Vempala and Wibisono, 2019; Anari et al., 2024), discretizations of the underdamped Langevin diffusion (Altschuler et al., 2025; Lehec, 2025; Ma et al., 2021; Zhang et al., 2023; Camrud et al., 2023; Kandasamy and Nagaraj, 2024; Anari et al., 2024), as well as proximal samplers (Wibisono, 2019; Chen et al., 2022; Fan et al., 2023; Liang and Chen, 2022; Mitra and Wibisono, 2025), among others.
+In these works, convergence is typically established by comparing the discrete-time sampling algorithm to an associated continuous-time process, such as the Langevin diffusion, whose convergence properties are well understood in terms of functional inequalities.
+In contrast, the analysis of Andrieu et al.
+(2024) is carried out directly at the level of the discrete-time Markov chain, without relying on an explicit comparison with a continuous-time diffusion.
+For the Gibbs sampler, this distinction is especially relevant, as there does not appear to exist a natural continuous-time process against which the algorithm can be compared.
 It is therefore interesting to investigate whether the framework developed by Andrieu et al.
 (2024) can be applied to the Gibbs sampler targeting non-log-concave distributions, specifically those satisfying either a Poincaré or log-Sobolev inequality.
+Another central notion in Andrieu et al.
+(2024) is that of a so-called close coupling condition for a Markov chain, which provides a quantitative link between geometry and probability.
+Informally, it requires that whenever two points are “close” in an appropriate metric, the distributions obtained after one step of the Markov chain from these points are “close” in total variation distance.
+The importance of such connections between geometric and probabilistic distances in the context of mixing times goes back at least to the work of Lovász (1999, Section 5), which studies them in the context of the Hit-and-Run algorithm on convex bodies.
+In Andrieu et al.
+(2024), this idea is formalized into a condition that, together with isoperimetry, yields explicit mixing time bounds directly at the level of the discrete-time Markov chain.
 1.1 Our contributions
 In this paper, we establish conductance bounds for the systematic-scan and random-scan Gibbs samplers over a broad class of target distributions supported on $\mathbb{R}^{d}$.
-In particular, consider a target distribution $\pi$ satisfying a $L^{q}$-Poincaré inequality with constant $\mathtt{p}_{q}$ and whose one-dimensional conditional distributions are Hölder continuous in total variation distance with constant $M$ and exponent $\beta$ (see Subsection 2.2 for formal definitions).
+In particular, consider a target distribution $\pi$ satisfying an $L^{q}$-Poincaré inequality with constant $\mathtt{p}_{q}$ and whose one-dimensional conditional distributions are Hölder continuous in total variation distance with constant $M$ and exponent $\beta$ (see Subsection 2.2 for formal definitions).
 Denote by $P_{\mathrm{SS}}$ and $P_{\mathrm{RS}}$ the Markov kernels corresponding to the systematic-scan and random-scan Gibbs samplers, respectively.
-Then, for sufficiently large $d$, our main results (Theorem 4.7 and Corollary 4.9) provide the following bounds on the conductance of the systematic-scan Gibbs kernel, $\Phi(P_{\mathrm{SS}})$, and the random-scan Gibbs kernel, $\Phi(P_{\mathrm{RS}})$:
-$\Phi(P_{\mathrm{SS}})\gtrsim\frac{\mathtt{p}_{q}}{\left(Md\right)^{q/\beta}},$
-$\Phi(P_{\mathrm{RS}})\gtrsim\frac{\mathtt{p}_{q}}{M^{q/\beta}(d\cdot\log d)^{q/\beta+1}},$
-where $\gtrsim$ denotes inequalities that hold up to a universal constant. Bounds on the mixing time in total variation distance then follow directly from Lovász and Simonovits (1993, Corollary 1.5).
+Then, for sufficiently large $d$, our main results (Theorem 4.7 and Corollary 4.9) provide the following bounds on the conductance of the lazy systematic-scan Gibbs kernel, $\Phi(\widetilde{P}_{\mathrm{SS}})$, and the random-scan Gibbs kernel, $\Phi(P_{\mathrm{RS}})$:
+$\Phi(\widetilde{P}_{\mathrm{SS}})\gtrsim\frac{\mathtt{p}_{q}}{2^{2q}(Md)^{q/\beta}},$
+$\Phi(P_{\mathrm{RS}})\gtrsim\frac{\mathtt{p}_{q}}{2^{2q}(3M)^{q/\beta}(d\cdot\log d)^{q/\beta+1}},$
+where $\gtrsim$ denotes inequalities that hold up to a universal constant. Bounds on the mixing time in total variation distance then follow from results of Lovász and Simonovits (1993).
 To prove the conductance bounds, we establish novel intermediate results that may also be of independent interest.
-  • We prove new three-set isoperimetric inequalities for distributions satisfying a $L^{q}$-Poincaré (or log-Sobolev) inequality, without assuming log-concavity (Propositions 4.1 and 4.3).
+  • We prove new three-set isoperimetric inequalities for distributions satisfying an $L^{q}$-Poincaré or $L^{q}$-log-Sobolev inequality, without assuming log-concavity (Propositions 4.1 and 4.3).
 This is achieved by applying these functional inequalities to a carefully chosen test function.
@@
   • We show that if $\pi$ is log-Lipschitz continuous or log-smooth, then the one-dimensional conditional distributions under $\pi$ are Hölder continuous in total variation distance (Lemmas 5.4 and 5.5).
-The proof relies on the fact that log-Lipschitz continuity and log-smoothness are inherited by the marginal distributions of $\pi$, as shown in Lemma 5.3.
+In the log-smooth case, this is shown under the additional assumption that the one-dimensional conditional distributions of $\pi$ satisfy a uniform $L^{2}$-Poincaré inequality.
+The proof relies on the fact that log-Lipschitz continuity and log-smoothness (under the above assumption) are inherited by the marginal distributions of $\pi$, as shown in Lemma 5.3.
 After establishing the three-set isoperimetric inequalities and close coupling conditions, bounds on the conductance follow from Proposition 3.11 below, which is a slight modification of Andrieu et al.
@@
 1.2 Connections with recent work
-As a special case, the conductance approach has been successfully applied to the random-scan Gibbs sampler targeting uniform distributions over convex bodies, where quantitative convergence guarantees have been established (Narayanan and Srivastava, 2022; Laddha and Vempala, 2023).
+As a special case, the conductance approach has been successfully applied to the random-scan Gibbs sampler targeting uniform distributions over convex bodies, where quantitative convergence guarantees have been established (Narayanan and Srivastava, 2022; Laddha and Vempala, 2023; Narayanan et al., 2024).
 Narayanan and Srivastava (2022) establish a close coupling condition for a multi-step Metropolis-within-Gibbs scheme, where the corresponding Markov kernel updates a randomly chosen coordinate using a single step of random-walk Metropolis while keeping the others fixed.
@@
 This allows them to control the mixing behavior without relying on auxiliary chains.
-In our work, we extend these ideas to general target distributions and establish three-set isoperimetric inequalities that hold for arbitrary partitions.
+More recently, Narayanan et al.
+(2024) refine these results by reducing the dependence on the initialization (the “warm start” parameter) of the mixing time bounds from polynomial to logarithmic.
+Their approach decomposes the convex body into axis-aligned dyadic cubes and establishes an isoperimetric inequality with respect to a metric that magnifies distances near the boundary.
+The dimension dependence of their mixing time bounds can be further improved using the sharper isoperimetric inequalities of Fernandez (2025).
+In this work, we develop a complementary approach for target distributions supported on $\mathbb{R}^{d}$ and establish three-set isoperimetric inequalities with respect to the Euclidean metric that hold for arbitrary partitions.
+However, it is unclear whether the results in Narayanan and Srivastava (2022); Laddha and Vempala (2023); Narayanan et al.
+(2024) for sampling from convex bodies can be recovered from our framework.
 Additionally, the conductance approach has been used to establish rapid mixing of the Gibbs sampler targeting the posterior distributions in Bayesian probit, logit, and lasso regression models (Lee and Zhang, 2024).
@@
 For such data-augmented chains, Lee and Zhang (2024) show that close coupling conditions follow directly from the conditional independence structure of the posterior distribution.
-Isoperimetric inequalities are obtained via transference inequalities to a transformed Markov chain with a log-concave target.
+Isoperimetric inequalities are obtained via transference inequalities to a transformed Markov chain with a log-concave target distribution.
 Under suitable assumptions on the data-generating process, these authors then establish bounds on the mixing time that hold with high probability.
@@
 Finally, Section 5 extends these bounds to log-Lipschitz continuous and log-smooth target distributions.
-We also provide examples of non-log-concave target distributions for which our results can be applied to establish mixing time bounds for both variants of the Gibbs sampler, filling a gap not addressed by any other existing literature.
+We also provide examples of non-log-concave target distributions for which our results can be applied to establish mixing time bounds for both variants of the Gibbs sampler, filling a gap in the existing literature.
 2 Preliminaries
@@
 We consider a probability distribution $\pi(\mathrm{d}x)$ supported on $\mathbb{R}^{d}$, equipped with its standard Borel $\sigma$-algebra.
-Let $P(x,\mathrm{d}y)$ be the Markov transition kernel of a discrete-time Markov chain on $\mathbb{R}^{d}$ with stationary (or invariant) distribution $\pi$.
+Let $P(x,\mathrm{d}y)$ be the Markov transition kernel of a discrete-time Markov chain on $\mathbb{R}^{d}$ with invariant distribution $\pi$.
 Here, the Markov chain is interpreted as the MCMC method for generating approximate samples from $\pi$.
@@
 $\pi(\mathrm{d}x)P(x,\mathrm{d}y)=\pi(\mathrm{d}y)P(y,\mathrm{d}x)\qquad\forall x,y\in\mathbb{R}^{d}.$
-Reversibility with respect to $\pi$ implies stationarity of $\pi$, however the reverse implication is in general false.
+Reversibility with respect to $\pi$ implies that $\pi$ is invariant, however the reverse implication is in general false.
 Among MCMC methods, we focus on the Gibbs sampler.
@@
 We additionally use the notation $\left[d\right]=\left\{1,\dots,d\right\}$.
-Let $X\sim\pi$.
-Denote the conditional distribution of $X_{k}$ given $X_{-k}=x_{-k}$ as $\pi(\cdot\mid x_{-k})$.
+Let $X\sim\pi$, that is, $X$ is distributed according to $\pi$.
+Denote the conditional distribution of $X_{k}$ given $X_{-k}=x_{-k}$ by $\pi(\cdot\mid x_{-k})$.
 For all $k\in[d]$, define the Markov kernel
-$P_{k}(x,\mathrm{d}y)=\pi(\mathrm{d}y_{k}\mid x_{-k})\delta_{x_{-k}}(\mathrm{d}y_{-k}),$
-where $\delta_{x_{-k}}$ denotes the $(d-1)$-dimensional Dirac delta distribution: for all measurable $S\subseteq\mathbb{R}^{d-1}$,
-$\delta_{x_{-k}}(S)=\begin{cases}1&\text{if }x_{-k}\in S,\\ 0&\text{if }x_{-k}\notin S.\end{cases}$
-Using the kernels $P_{k}$, we now define the Markov kernels corresponding to the random-scan and systematic-scan Gibbs samplers.
-For two Markov kernels $P$ and $Q$, we denote their composition as $PQ(x,\mathrm{d}y)=\int P(u,\mathrm{d}y)Q(x,\mathrm{d}u)$.
+$P_{k}(x,\mathrm{d}y)=\pi(\mathrm{d}y_{k}\mid x_{-k})\delta_{x_{-k}}(\mathrm{d}y_{-k}),$ (2)
+where $\delta_{x_{-k}}$ denotes the (d1)(d-1)-dimensional Dirac delta distribution: for all measurable $S\subseteq\mathbb{R}^{d-1}$,
+$\delta_{x_{-k}}(S)=\begin{cases}1&\mathrm{if}~x_{-k}\in S,\\ 0&\mathrm{if}~x_{-k}\notin S.\end{cases}$
+Using the kernels $(P_{k})_{k\in[d]}$, we now define the Markov kernels corresponding to the random-scan and systematic-scan Gibbs samplers.
+For two Markov kernels $P$ and $Q$, we denote their composition by $PQ(x,\mathrm{d}y)=\int Q(u,\mathrm{d}y)P(x,\mathrm{d}u)$.
+With this convention, $PQ$ corresponds to first updating the chain according to $P$ and then according to $Q$.
 Definition 2.1.
-The Markov kernel associated with the random-scan Gibbs sampler is defined as the mixture of the Markov kernels $P_{k}$:
+The Markov kernel associated with the random-scan Gibbs sampler is defined as the mixture of the Markov kernels $(P_{k})_{k\in[d]}$:
 $P_{\mathrm{RS}}=\frac{1}{d}\sum_{k=1}^{d}P_{k}.$
-The Markov kernel associated with the systematic-scan Gibbs sampler is defined as the composition of the Markov kernels $P_{k}$:
-$P_{\mathrm{SS}}=P_{d}P_{d-1}\dots P_{2}P_{1}.$
+The Markov kernel associated with the systematic-scan Gibbs sampler is defined as the composition of the Markov kernels $(P_{k})_{k\in[d]}$:
+$P_{\mathrm{SS}}=P_{1}P_{2}\dots P_{d-1}P_{d}.$
 In words, the random-scan Gibbs sampler selects a single coordinate uniformly at random and resamples it from its conditional distribution, while leaving the other coordinates unchanged.
 In contrast, the systematic-scan Gibbs sampler sequentially updates each coordinate in a predetermined order, from 1 to $d$, resampling each from its respective conditional distribution.
-It is worth noting that the Markov kernel associated with the random-scan Gibbs sampler is $\pi$-reversible, whereas the Markov kernel associated with the systematic-scan Gibbs sampler is generally not.
+Naturally, both variants of the Gibbs sampler are $\pi$-invariant.
+However, the Markov kernel associated with the random-scan Gibbs sampler is reversible with respect to $\pi$, whereas the Markov kernel associated with the systematic-scan Gibbs sampler is generally not.
 To address this, one approach is to use the multiplicative reversibilization of the systematic-scan Gibbs sampler.
 The kernel for this version is defined as
-$P_{\mathrm{SS}}^{*}P_{\mathrm{SS}}=P_{1}P_{2}\dots P_{d-1}P_{d}P_{d-1}\dots P_{2}P_{1},$
-where $P^{*}$ represents the adjoint operator of a Markov kernel $P$.
+$P_{\mathrm{SS}}P_{\mathrm{SS}}^{*}=P_{1}P_{2}\dots P_{d-1}P_{d}^{2}P_{d-1}\dots P_{2}P_{1},$
+where $P^{*}$ represents the adjoint of a Markov kernel $P$ (see Subsection 3.1).
 In this version, the coordinates are updated first in increasing order, from $1$ to $d$, and then in decreasing order, from $d$ to $1$.
 2.2 Assumptions on the target distribution
-[…]
+In the following, let $\left\lvert\cdot\right\rvert$ denote the standard Euclidean norm on $\mathbb{R}^{d}$.
+Unless stated otherwise, all expectations and variances are taken with respect to the target distribution $\pi$.
+We say that a function $f:\mathbb{R}^{d}\to\mathbb{R}$ is locally Lipschitz if its restriction to any open ball in $\mathbb{R}^{d}$ is Lipschitz continuous.
+With these conventions in mind, we now consider two functional inequalities for $\pi$, namely the Poincaré and log-Sobolev inequalities.
 Definition 2.2.
 Let $q\geq 1$.
-The distribution $\pi$ is said to satisfy a $L^{q}$-Poincaré inequality ($\mathrm{PI}_{q}$) with constant $\mathtt{p}_{q}>0$ if for any locally Lipschitz function $f:\mathbb{R}^{d}\to\mathbb{R}$,
+The distribution $\pi$ is said to satisfy an $L^{q}$-Poincaré inequality ($\mathrm{PI}_{q}$) with constant $\mathtt{p}_{q}>0$ if for any locally Lipschitz function $f:\mathbb{R}^{d}\to\mathbb{R}$,
 $\mathbb{E}\left[\left\lvert f-\mathbb{E}[f]\right\rvert^{q}\right]\leq\frac{1}{\mathtt{p}_{q}}\cdot\mathbb{E}\left[\left\lvert\nabla f\right\rvert^{q}\right].$ (3)
@@
 Let $q\geq 1$.
-The distribution $\pi$ is said to satisfy a $L^{q}$-log-Sobolev inequality ($\mathrm{LSI}_{q}$) with constant $\mathtt{s}_{q}>0$ if for any locally Lipschitz function $f:\mathbb{R}^{d}\to\mathbb{R}$,
+The distribution $\pi$ is said to satisfy an $L^{q}$-log-Sobolev inequality ($\mathrm{LSI}_{q}$) with constant $\mathtt{s}_{q}>0$ if for any locally Lipschitz function $f:\mathbb{R}^{d}\to\mathbb{R}$,
 $\mathbb{E}\left[\left\lvert f\right\rvert^{q}\cdot\log\left\lvert f\right\rvert^{q}\right]-\mathbb{E}\left[\left\lvert f\right\rvert^{q}\right]\cdot\log\mathbb{E}\left[\left\lvert f\right\rvert^{q}\right]\leq\frac{1}{\mathtt{s}_{q}}\cdot\mathbb{E}\left[\left\lvert\nabla f\right\rvert^{q}\right].$ (4)
@@
 As shown in Bobkov and Zegarlinski (2005, Proposition 6.2), $\mathrm{LSI}_{q}$ is also preserved by bounded perturbation à la Holley and Stroock (1987).
+We recall that a distribution $\nu$ is said to be a bounded perturbation of another distribution $\mu$ if there exist constants $0<c_{1}\leq c_{2}<\infty$ such that the Radon-Nikodym derivative satisfies $c_{1}\leq\mathrm{d}\nu/\mathrm{d}\mu\leq c_{2}$.
 More generally, Ledoux (1999) proved that $\mathrm{PI}_{2}$ and $\mathrm{LSI}_{2}$ respectively imply sub-exponential and sub-Gaussian tail bounds for the distribution $\pi$.
-Log-Sobolev inequalites have also been proven for mixture distributions (Chen et al., 2021).
+Poincaré inequalities have been established for a broad class of distributions satisfying a Lyapunov condition (Bakry et al., 2008) and log-Sobolev inequalities have also been proven for mixture distributions (Chen et al., 2021).
 Furthermore, if $\pi$ satisfies $\mathrm{LSI}_{q}$ with constant $\mathtt{s}_{q}$, then Bobkov and Zegarlinski (2005, Theorem 2.1) showed that $\pi$ satisfies $\mathrm{PI}_{q}$ with constant $\mathtt{p}_{q}\geq 4\mathtt{s}_{q}/\log 2$.
@@
 $\frac{5}{128}\left(\frac{4}{105}\right)^{r/q}\left(\frac{q}{r}\right)^{r}\mathtt{s}_{q}^{r/q}\leq\mathtt{s}_{r}.$
-To the best of our knowledge, Proposition 2.4 does not appear explicitly in the literature. For completeness, we include a proof in Appendix A, which may also be of independent interest.
+To the best of our knowledge, Proposition 2.4 does not appear explicitly in the literature. For completeness, we include a proof in Appendix A, which may be of independent interest.
 We now introduce one final regularity assumption on the conditional distributions of $\pi$.
 To formalize this, recall that the total variation (TV) distance between two probability distributions $\mu$ and $\nu$ supported on $\mathbb{R}^{d}$ is given by
-$\left\|\mu-\nu\right\|_{\textsc{TV}}=\sup\left\{\left\lvert\mu(S)-\nu(S)\right\rvert:S\subseteq\mathbb{R}^{d}\text{ measurable}\right\}.$
+$\left\|\mu-\nu\right\|_{\textsc{TV}}=\sup\left\{\left\lvert\mu(S)-\nu(S)\right\rvert:S\subseteq\mathbb{R}^{d}~\mathrm{measurable}\right\}.$
 Definition 2.5.
 Let $M>0$ and $\beta\in(0,1]$. The conditional distributions of $\pi$ are said to be $(M,\beta)$-TV continuous if
-$\left\|\pi(\cdot\mid x_{-k})-\pi(\cdot\mid y_{-k})\right\|_{\textsc{TV}}\leq M\left\lvert x_{-k}-y_{-k}\right\rvert^{\beta}\qquad\forall k\in[d],~{}\forall x,y\in\mathbb{R}^{d}.$
+$\left\|\pi(\cdot\mid x_{-k})-\pi(\cdot\mid y_{-k})\right\|_{\textsc{TV}}\leq M\left\lvert x_{-k}-y_{-k}\right\rvert^{\beta}\qquad\forall k\in[d],\,\forall x,y\in\mathbb{R}^{d}.$
 This type of continuity has been shown to hold for posterior distributions in Bayesian models, under assumptions such as Lipschitz continuity of the likelihood and sufficient integrability of the prior (Dolera and Mainini, 2020, 2023).
-Moreover, in Subsection 5.1, we will show how to verify that the conditional distributions of $\pi$ are TV-continuous under standard conditions on the potential function $U=-\log\pi$.
+Moreover, in Subsection 5.1, we will show how to verify that the conditional distributions of $\pi$ are TV continuous under standard conditions on the potential function $U=-\log\pi$.
 3 Convergence analysis of Markov chains
+3.1 Operator representation and spectral gap
+Let
+$L^{2}_{\pi}=\left\{f:\mathbb{R}^{d}\to\mathbb{R}:\int f(x)^{2}\pi(\mathrm{d}x)<\infty\right\}$
+denote the Hilbert space of square $\pi$-integrable functions. For all $f,g\in L^{2}_{\pi}$, the scalar product associated to this space is
+$\left\langle f,g\right\rangle=\int f(x)g(x)\pi(\mathrm{d}x),$
+with corresponding $L^{2}_{\pi}$-norm $\left\|f\right\|_{2}=\sqrt{\left\langle f,f\right\rangle}$.
+Note that a $\pi$-invariant Markov kernel $P$ can be seen as an operator that acts on functions $f\in L^{2}_{\pi}$ via
+$Pf(x)=\int f(y)P(x,\mathrm{d}y).$
+With this identification, reversibility of $P$ is equivalent to self-adjointness of the associated operator on $L^{2}_{\pi}$.
+In the reversible setting, the Markov kernel $P$ is said to be positive semi-definite if for all $f\in L^{2}_{\pi}$,
+$\left\langle f,Pf\right\rangle\geq 0.$
+In particular, the random-scan Gibbs kernel $P_{\mathrm{RS}}$ is positive semi-definite.
+Indeed, each kernel $P_{k}$ defined in (2) satisfies $P_{k}f(x)=\mathbb{E}[f(X)\mid X_{-k}=x_{-k}]$ and is positive semi-definite since
+$\left\langle f,P_{k}f\right\rangle=\mathbb{E}[f(X)\cdot\mathbb{E}[f(X)\mid X_{-k}]]=\mathbb{E}\left[\mathbb{E}\left[f(X)\mid X_{-k}\right]^{2}\right]\geq 0.$
+It follows that
+$\left\langle f,P_{\mathrm{RS}}f\right\rangle=\frac{1}{d}\sum_{k=1}^{d}\left\langle f,P_{k}f\right\rangle\geq 0.$
+A common way to ensure positive semi-definiteness is to work with lazy Markov kernels (see, for example, Lovász and Simonovits, 1993, Subsection 1.b). We say that the kernel $P$ is lazy if
+$P(x,\{x\})\geq\frac{1}{2}\qquad\forall x\in\mathbb{R}^{d}.$
+Any (reversible or non-reversible) Markov kernel can be made lazy as follows. Let $I$ denote the identity operator, defined by $If=f$.
+Definition 3.1.
+The lazy version of a Markov kernel $P$ is defined as
+$\widetilde{P}(x,\mathrm{d}y)=\frac{1}{2}\delta_{x}(\mathrm{d}y)+\frac{1}{2}P(x,\mathrm{d}y).$
+The associated operator is $\widetilde{P}=(I+P)/2$.
+In particular, the lazy version of a $\pi$-invariant Markov kernel $P$ is also $\pi$-invariant.
+An important quantity for analyzing the convergence of reversible Markov chains is the spectral gap, which we now recall.
+Definition 3.2.
+Consider a $\pi$-invariant Markov chain with transition kernel $P$. Suppose that $P$ is reversible. The spectral gap of $P$ is defined as
+$\lambda_{2}(P)=\inf\left\{\frac{\left\langle f,(I-P)f\right\rangle}{\left\|f\right\|^{2}_{2}}:f\in L^{2}_{\pi},\,\mathbb{E}[f]=0,\,f\neq 0\right\}.$
+When $P$ is reversible, the spectral gap provides a bound on the $L^{2}_{\pi}$-convergence rate of the Markov chain: for any $f\in L^{2}_{\pi}$ with $\mathbb{E}[f]=0$,
+$\left\|Pf\right\|_{2}\leq\left(1-\lambda_{2}(P)\right)\left\|f\right\|_{2}.$
+See (Levin and Peres, 2017, Exercise 12.4) for a proof.
+For non-reversible chains, various notions of spectral gaps have been proposed; see, for instance, Kontoyiannis and Meyn (2012) for definitions based on suitably weighted Banach spaces, and Chatterjee (2025) for finite state spaces.
+Other approaches (Fill, 1991; Paulin, 2015; Choi, 2020) use reversibilized Markov kernels to derive similar bounds on the convergence rate.
+However, the sharpness of such bounds in the non-reversible setting remains unclear, and a detailed investigation lies beyond the scope of this work.
 3.2 Mixing times of Markov chains
@@
 Let $\mu$ denote the initial distribution of the Markov chain associated to a kernel $P$.
-After $k$ steps, the distribution of the chain is given by $\mu P^{k}$, where
+After $k\geq 0$ steps, the distribution of the chain is given by $\mu P^{k}$, where
 $\mu P^{k}(\mathrm{d}x)=\int P^{k}(y,\mathrm{d}x)\mu(\mathrm{d}y).$
-It is often assumed that the initial distribution is “not too far” from the stationary distribution, a condition referred to as a warm start.
+It is often assumed that the initial distribution is “not too far” from the invariant distribution, a condition referred to as a warm start.
 Definition 3.3.
 Let $\Omega>0$. A $\pi$-invariant Markov chain with initial distribution $\mu$ is said to have a $\Omega$-warm start if
-$\sup\left\{\frac{\mu(S)}{\pi(S)}:S\subseteq\mathbb{R}^{d}\text{ measurable}\right\}\leq\Omega.$
+$\sup\left\{\frac{\mu(S)}{\pi(S)}:S\subseteq\mathbb{R}^{d}~\mathrm{measurable},\,\pi(S)>0\right\}\leq\Omega.$
 Note that the constant $\Omega$ can grow exponentially with the dimension $d$ even in fairly innocent examples.
 For instance, this occurs when the initial distribution is a Gaussian centered at the mode of a strongly log-concave target (Dwivedi et al., 2019).
-Nevertheless, when $\pi$ is strongly log-concave, Altschuler and Chewi (2024) showed that a distribution with a $\mathcal{O}(1)$-warm start can be efficiently obtained using the underdamped Langevin dynamics in $\mathcal{O}(\sqrt{d})$ iterations.
-We now define the mixing time of a Markov chain as the minimum number of iterations required for the chain to be “close” to its stationary distribution in total variation distance.
+Nevertheless, when $\pi$ is strongly log-concave, Altschuler and Chewi (2024a) showed that an initial distribution that is within $\mathcal{O}(1)$ $\chi^{2}$-distance from $\pi$ can be efficiently obtained in $\mathcal{O}(\sqrt{d})$ iterations using the underdamped Langevin dynamics.
+We note, however, that this notion of proximity for the initial distribution is weaker than the warm start in Definition 3.3.
+We now define the mixing time of a Markov chain as the minimum number of iterations required for the chain to be “close” to its invariant distribution in total variation distance.
 Definition 3.4.
@@
 Consider a $\pi$-invariant Markov chain with transition kernel $P$. The conductance of $P$ is defined as
-$\Phi(P)=\inf\left\{\frac{\int_{S}P(x,S^{c})\pi(\mathrm{d}x)}{\pi(S)}:S\subseteq\mathbb{R}^{d}\text{ measurable},\text{ }\pi(S)\leq\frac{1}{2}\right\}.$
+$\Phi(P)=\inf\left\{\frac{\int_{S}P(x,S^{c})\pi(\mathrm{d}x)}{\pi(S)}:S\subseteq\mathbb{R}^{d}~\mathrm{measurable},\,0<\pi(S)\leq\frac{1}{2}\right\}.$
+In particular, the conductance of a Markov kernel $P$ and that of its lazy version $\widetilde{P}$ are related by
+$\Phi(\widetilde{P})=\frac{1}{2}\cdot\Phi(P).$ (5)
 The connection between the mixing time of a Markov chain and its conductance was established by Lovász and Simonovits (1993), who extended the earlier work of Jerrum and Sinclair (1988) to general state spaces.
-Importantly, this result does not require the transition kernel $P$ to be reversible.
-Theorem 3.4 (Lovász and Simonovits, 1993, Corollary 1.5).
-Consider a $\pi$-invariant Markov chain with transition kernel $P$ initialized at a $\Omega$-warm start. Then, for all $k\geq 0$,
+Theorem 3.6.
+Consider a $\pi$-invariant Markov chain with transition kernel $P$ initialized at a $\Omega$-warm start.
+Suppose that $P$ is either lazy, or both reversible and positive semi-definite.
+Then, for all $k\geq 0$,
 $\left\|\mu P^{k}-\pi\right\|_{\textsc{TV}}\leq\sqrt{\Omega}\cdot\exp\left(-\frac{k}{2}\cdot\Phi(P)^{2}\right).$
+Proof.
+If $P$ is lazy, the bound follows from Lovász and Simonovits (1993, Corollary 1.5).
+If $P$ is reversible and positive semi-definite, the same bound follows from Lee and Zhang (2024, Lemma 5.3), which is a modified version of the former result.
+∎
 Corollary 3.7.
-Consider a $\pi$-invariant Markov chain with transition kernel $P$ initialized at a $\Omega$-warm start. Then, its $\zeta$-mixing time can be bounded by
+Let $\zeta>0$. Under the conditions of Theorem 3.6, the $\zeta$-mixing time of the Markov chain is bounded by
 $\tau(\zeta,P)\leq\frac{2}{\Phi(P)^{2}}\cdot\log\left(\frac{\sqrt{\Omega}}{\zeta}\right).$
 The problem of upper bounding the mixing time is thus reduced to proving a lower bound on the conductance.
-3.2 The spectral gap
-Let
-$L^{2}_{\pi}=\left\{f:\mathbb{R}^{d}\to\mathbb{R}:\int f(x)^{2}\pi(\mathrm{d}x)<\infty\right\}$
-denote the Hilbert space of square $\pi$-integrable functions. For all $f,g\in L^{2}_{\pi}$, the scalar product associated to this space is
-$\langle f,g\rangle=\int f(x)g(x)\pi(\mathrm{d}x),$
-with corresponding $L^{2}_{\pi}$-norm $\left\|f\right\|_{2}=\sqrt{\langle f,f\rangle}$. Note that a Markov kernel $P$ can be seen as an operator that acts on functions $f\in L^{2}_{\pi}$ via
-$Pf(x)=\int f(y)P(x,\mathrm{d}y).$
-Let also $I$ denote the identity operator: $If=f$.
-Definition 3.6.
-Consider a $\pi$-invariant Markov chain with transition kernel $P$. Suppose that $P$ is $\pi$-reversible. The spectral gap of $P$ is defined as
-$\lambda_{2}(P)=\inf\left\{\frac{\langle f,(I-P)f\rangle}{\left\|f\right\|_{2}}:f\in L^{2}_{\pi},\text{ }\mathbb{E}[f]=0\right\}.$
-When $P$ is reversible, the spectral gap provides a bound on the $L^{2}_{\pi}$-convergence rate of the Markov chain: for any $f\in L^{2}_{\pi}$ with $\mathbb{E}\left[f\right]=0$, we have
-$\left\|Pf\right\|_{2}\leq\left(1-\lambda_{2}\right)\left\|f\right\|_{2}.$
-See (Levin and Peres, 2017, Exercise 12.4) for a proof.
-The spectral gap of a reversible Markov kernel $P$ is related to its conductance via the following bounds, also known as Cheeger’s inequalities (Cheeger, 1970).
+The spectral gap of a reversible Markov kernel is also related to its conductance via the following bounds, also known as Cheeger’s inequalities (Cheeger, 1970).
 Theorem 3.8 (Lawler and Sokal, 1988, Theorem 2.1).
@@
 $\frac{1}{2}\cdot\Phi^{2}(P)\leq\lambda_{2}(P)\leq\Phi(P).$
-A lower bound on the conductance thus translates to a lower bound on the spectral gap.
-Theorem 3.7 applies specifically to the reversible setting and covers the random-scan Gibbs sampler as well as reversibilized forms of the systematic-scan Gibbs sampler, such as the variant with kernel $P_{\mathrm{SS}}^{*}P_{\mathrm{SS}}$ discussed in Subsection 2.1.
-In this paper, however, we do not directly assume reversibility and instead focus on mixing time bounds as in Corollary 3.5.
-For non-reversible chains, various notions of spectral gaps have been proposed; see, for instance, Kontoyiannis and Meyn (2012) for definitions based on suitably weighted Banach spaces, and Chatterjee (2023) for finite state spaces.
-Other approaches (Fill, 1991; Paulin, 2015; Choi, 2020) use reversibilized Markov kernels to derive similar bounds on the convergence rate.
-However, the sharpness of such bounds in the non-reversible setting remains unclear, and a detailed investigation lies beyond the scope of this work.
+Theorem 3.8 applies specifically to the reversible setting and covers the random-scan Gibbs sampler as well as reversibilized forms of the systematic-scan Gibbs sampler, such as the variant with kernel $P_{\mathrm{SS}}P_{\mathrm{SS}}^{*}$ discussed in Subsection 2.1.
 3.3 Bounding the conductance
-In order to bound the conductance of a $\pi$-invariant Markov kernel $P$, two main components are required: a three-set isoperimetric inequality for the distribution $\pi$, and a close coupling condition for the kernel $P$.
+In order to bound the conductance of a $\pi$-invariant Markov kernel $P$, two main components are required: an isoperimetric inequality for the distribution $\pi$, and a close coupling condition for the kernel $P$.
+Following the terminology of Andrieu et al.
+(2024), we refer to the particular isoperimetric inequalities used here as “three-set” isoperimetric inequalities.
 In what follows, we write $\mathrm{Id}:\mathbb{R}\to\mathbb{R}$ for the identity function and use $\sqcup$ to denote the union of disjoint sets.
@@
 Let $\Upsilon:(0,\infty)\to(0,\infty)$ and $\Psi:(0,1/2)\to(0,\infty)$ be two non-decreasing functions.
-Additionaly, assume that $\Psi/\mathrm{Id}$ is non-increasing.
+Additionally, suppose that $\Psi/\mathrm{Id}$ is non-increasing.
 The distribution $\pi$ is said to satisfy a three-set isoperimetric inequality with functions $\Upsilon,\Psi$, if for all measurable partitions $S_{1}\sqcup S_{2}\sqcup S_{3}=\mathbb{R}^{d}$,
@@
 When $\pi$ is log-concave (or a bounded perturbation thereof), a three-set inequality can be derived by controlling a quantity known as the isoperimetric profile of $\pi$ Milman, 2009b, Theorem 1.2; Andrieu et al., 2024, Section 3.
-Bounds on the isoperimetric profile have also been established when $\pi$ satisfies $\mathrm{PI}_{1}$ and $\mathrm{LSI}_{1}$ (Rothaus, 1985; Bobkov and Houdré, 1997).
+Appropriate bounds on the isoperimetric profile are known to be equivalent to $\pi$ satisfying $\mathrm{PI}_{1}$ and, similarly, $\mathrm{LSI}_{1}$ (Rothaus, 1985; Bobkov and Houdré, 1997; Kolesnikov, 2007).
 In all these cases, $\Upsilon$ is taken to be a multiple of the identity function.
@@
 Notably, Proposition 3.11 and Corollary 3.12 do not require the Markov kernel $P$ to be reversible.
-Conventionally, $\delta$ is chosen to scale inversely with the dimension $d$ to ensure that $\varepsilon$ remains of constant order.
+Conventionally, close coupling conditions are established with the parameter $\delta$ scaling inversely with the dimension $d$ and the parameter $\varepsilon$ being of constant order.
 Under this regime, Corollary 3.12 yields
@@
 To overcome these challenges, we take a more direct approach to derive three-set isoperimetric inequalities for distributions satisfying $\mathrm{PI}_{q}$ or $\mathrm{LSI}_{q}$ with $q\geq 1$.
-In this setting, $\Upsilon$ is no longer restricted to be proportional to the identity function.
+In these settings, $\Upsilon$ is no longer restricted to be proportional to the identity function.
 In what follows, denote the indicator function of a measurable set $S\subseteq\mathbb{R}^{d}$ as $\mathbbm{1}_{S}$.
@@
 Let $q\geq 1$. If the distribution $\pi$ satisfies $\mathrm{PI}_{q}$ with constant $\mathtt{p}_{q}$, then $\pi$ satisfies a three-set isoperimetric inequality with functions
-$\Upsilon(t)=\frac{\mathtt{p}_{q}t^{q}}{2^{2q}\left(1+2^{q}\mathtt{p}_{q}t^{q}\right)},\qquad\Psi(t)=t.$
+$\Upsilon(t)=\frac{\mathtt{p}_{q}t^{q}}{2^{2q-1}(1+2^{q}\mathtt{p}_{q}t^{q})},\qquad\Psi(t)=t.$
 Proof.
@@
 Swapping the roles of $x$ and $y$ and letting $\xi\to 0$, we obtain $\left\lvert\mathcal{D}(x,S_{1})-\mathcal{D}(y,S_{1})\right\rvert\leq\left\lvert x-y\right\rvert$.
-It follows that the function $f$ is equal to 1 on $S_{1}$, vanishes on $S_{2}$, and is Lipschitz continuous on a subset of $S_{3}$, with the modulus of its gradient equal to $\mathcal{D}(S_{1},S_{2})^{-1}$; elsewhere on $S_{3}$, $f$ vanishes.
-In particular, $f(x)\in[0,1]$ for $x\in S_{3}$.
+Hence, the function $f$ is Lipschitz continuous on $\mathbb{R}^{d}$.
+By Rademacher’s theorem, $f$ is differentiable almost everywhere and satisfies
+$\left\lvert\nabla f(x)\right\rvert\leq\frac{1}{\mathcal{D}(S_{1},S_{2})}$
+for almost every $x\in\mathbb{R}^{d}$.
+Moreover, the function $f$ takes the constant value 1 on $S_{1}$, vanishes on $S_{2}$, and satisfies $f(x)\in[0,1]$ for all $x\in S_{3}$.
 Applying (3) to $f$, we obtain
-$\left\|f-\mathbb{E}[f]\right\|_{q}^{q}=\mathbb{E}\left[\left\lvert f-\mathbb{E}[f]\right\rvert^{q}\right]\leq\frac{1}{\mathtt{p}_{q}}\cdot\mathbb{E}\left[\left\lvert\nabla f\right\rvert^{q}\right]\leq\frac{\pi(S_{3})}{\mathtt{p}_{q}\cdot\mathcal{D}(S_{1},S_{2})^{q}}.$ (4)
+$\left\|f-\mathbb{E}[f]\right\|_{q}^{q}=\mathbb{E}[\left\lvert f-\mathbb{E}[f]\right\rvert^{q}]\leq\frac{1}{\mathtt{p}_{q}}\cdot\mathbb{E}[\left\lvert\nabla f\right\rvert^{q}]\leq\frac{\pi(S_{3})}{\mathtt{p}_{q}\cdot\mathcal{D}(S_{1},S_{2})^{q}}.$ (6)
 We now compare $\mathbb{E}[\left\lvert f-\mathbb{E}[f]\right\rvert^{q}]$ to the corresponding quantity for the indicator function $\mathbbm{1}_{S_{1}}$, namely $\mathbb{E}\left[\left\lvert\mathbbm{1}_{S_{1}}-\mathbb{E}[\mathbbm{1}_{S_{1}}]\right\rvert^{q}\right]$.
 By the triangle inequality and Jensen’s inequality, we get
-$\mathbb{E}\left[\left\lvert\mathbbm{1}_{S_{1}}-\mathbb{E}[\mathbbm{1}_{S_{1}}]\right\rvert^{q}\right]^{1/q}$ $=\left\|\mathbbm{1}_{S_{1}}-\mathbb{E}[\mathbbm{1}_{S_{1}}]\right\|_{q}$
+$\mathbb{E}[\left\lvert\mathbbm{1}_{S_{1}}-\mathbb{E}[\mathbbm{1}_{S_{1}}]\right\rvert^{q}]^{1/q}$ $=\left\|\mathbbm{1}_{S_{1}}-\mathbb{E}[\mathbbm{1}_{S_{1}}]\right\|_{q}$
 $\leq\left\|f-\mathbbm{1}_{S_{1}}\right\|_{q}+\left\|f-\mathbb{E}[f]\right\|_{q}+\left\lvert\mathbb{E}[f]-\mathbb{E}[\mathbbm{1}_{S_{1}}]\right\rvert$
 $\leq 2\left\|f-\mathbbm{1}_{S_{1}}\right\|_{q}+\left\|f-\mathbb{E}[f]\right\|_{q}.$
-Using the inequality $\left(a+b\right)^{q}\leq 2^{q}\left(a^{q}+b^{q}\right)$ for $a,b\geq 0$, it follows that
-$\mathbb{E}\left[\left\lvert\mathbbm{1}_{S_{1}}-\mathbb{E}[\mathbbm{1}_{S_{1}}]\right\rvert^{q}\right]\leq 2^{q}\left(2^{q}\left\|f-\mathbbm{1}_{S_{1}}\right\|_{q}^{q}+\left\|f-\mathbb{E}[f]\right\|_{q}^{q}\right).$ (5)
+Using the inequality $(a+b)^{q}\leq 2^{q-1}(a^{q}+b^{q})$ for $a,b\geq 0$ and $q\geq 1$, it follows that
+$\mathbb{E}[\left\lvert\mathbbm{1}_{S_{1}}-\mathbb{E}[\mathbbm{1}_{S_{1}}]\right\rvert^{q}]\leq 2^{q-1}\left(2^{q}\left\|f-\mathbbm{1}_{S_{1}}\right\|_{q}^{q}+\left\|f-\mathbb{E}[f]\right\|_{q}^{q}\right).$ (7)
 Since $f-\mathbbm{1}_{S_{1}}$ vanishes on $S_{1}\sqcup S_{2}$ and takes values in $[0,1]$ on $S_{3}$, we have
-$\left\|f-\mathbbm{1}_{S_{1}}\right\|_{q}^{q}=\mathbb{E}\left[\left\lvert f-\mathbbm{1}_{S_{1}}\right\rvert^{q}\right]\leq\pi(S_{3}).$ (6)
+$\left\|f-\mathbbm{1}_{S_{1}}\right\|_{q}^{q}=\mathbb{E}[\left\lvert f-\mathbbm{1}_{S_{1}}\right\rvert^{q}]\leq\pi(S_{3}).$ (8)
 Inserting (6) and (8) into (7), we get
-$\mathbb{E}\left[\left\lvert\mathbbm{1}_{S_{1}}-\mathbb{E}[\mathbbm{1}_{S_{1}}]\right\rvert^{q}\right]\leq 2^{q}\left(2^{q}+\frac{1}{\mathtt{p}_{q}\cdot\mathcal{D}(S_{1},S_{2})^{q}}\right)\pi(S_{3}).$
+$\mathbb{E}[\left\lvert\mathbbm{1}_{S_{1}}-\mathbb{E}[\mathbbm{1}_{S_{1}}]\right\rvert^{q}]\leq 2^{q-1}\left(2^{q}+\frac{1}{\mathtt{p}_{q}\cdot\mathcal{D}(S_{1},S_{2})^{q}}\right)\pi(S_{3}).$
 We now derive a lower bound for $\mathbb{E}[\left\lvert\mathbbm{1}_{S_{1}}-\mathbb{E}[\mathbbm{1}_{S_{1}}]\right\rvert^{q}]$:
-$\mathbb{E}\left[\left\lvert\mathbbm{1}_{S_{1}}-\mathbb{E}[\mathbbm{1}_{S_{1}}]\right\rvert^{q}\right]$ $=\left(1-\pi(S_{1})\right)^{q}\pi(S_{1})+\pi(S_{1})^{q}\left(1-\pi(S_{1})\right)$
-$=\left(1-\pi(S_{1})\right)\pi(S_{1})\left[\left(1-\pi(S_{1})\right)^{q-1}+\pi(S_{1})^{q-1}\right]$
-$\geq\frac{1}{2^{q-1}}\left(1-\pi(S_{1})\right)\pi(S_{1})$
-$\geq\frac{1}{2^{q}}\cdot\min\left\{\pi(S_{1}),1-\pi(S_{1})\right\}$
-$\geq\frac{1}{2^{q}}\cdot\min\left\{\pi(S_{1}),\pi(S_{2})\right\},$
+$\mathbb{E}[\left\lvert\mathbbm{1}_{S_{1}}-\mathbb{E}[\mathbbm{1}_{S_{1}}]\right\rvert^{q}]$ $=(1-\pi(S_{1}))^{q}\cdot\pi(S_{1})+\pi(S_{1})^{q}(1-\pi(S_{1}))$
+$=(1-\pi(S_{1}))\cdot\pi(S_{1})\left[\left(1-\pi(S_{1})\right)^{q-1}+\pi(S_{1})^{q-1}\right]$
+$\geq\frac{1}{2^{q-1}}(1-\pi(S_{1}))\cdot\pi(S_{1})$
+$\geq\frac{1}{2^{q}}\cdot\min\{\pi(S_{1}),1-\pi(S_{1})\}$
+$\geq\frac{1}{2^{q}}\cdot\min\{\pi(S_{1}),\pi(S_{2})\},$
 where we have used the inequality $1/2\cdot\min\{x,1-x\}\leq x(1-x)$ for $x\in[0,1]$. Combining these bounds, we obtain
-$\min\left\{\pi(S_{1}),\pi(S_{2})\right\}\leq 2^{2q}\left(2^{q}+\frac{1}{\mathtt{p}_{q}\cdot\mathcal{D}(S_{1},S_{2})^{q}}\right)\pi(S_{3}),$
+$\min\{\pi(S_{1}),\pi(S_{2})\}\leq 2^{2q-1}\left(2^{q}+\frac{1}{\mathtt{p}_{q}\cdot\mathcal{D}(S_{1},S_{2})^{q}}\right)\pi(S_{3}),$
 or equivalently,
-$\pi(S_{3})\geq\frac{\mathcal{D}(S_{1},S_{2})^{q}}{2^{2q}\left(\mathtt{p}_{q}^{-1}+2^{q}\cdot\mathcal{D}(S_{1},S_{2})^{q}\right)}\cdot\min\left\{\pi(S_{1}),\pi(S_{2})\right\}.$
+$\pi(S_{3})\geq\frac{\mathcal{D}(S_{1},S_{2})^{q}}{2^{2q-1}\left(\mathtt{p}_{q}^{-1}+2^{q}\cdot\mathcal{D}(S_{1},S_{2})^{q}\right)}\cdot\min\{\pi(S_{1}),\pi(S_{2})\}.$
 The $L^{q}$-Poincaré inequality thus implies a three-set isoperimetric inequality
@@
 where
-$\Upsilon(t)=\frac{\mathtt{p}_{q}t^{q}}{2^{2q}\left(1+2^{q}\mathtt{p}_{q}t^{q}\right)}.$
+$\Upsilon(t)=\frac{\mathtt{p}_{q}t^{q}}{2^{2q-1}(1+2^{q}\mathtt{p}_{q}t^{q})}.$
 To derive a three-set isoperimetric inequality corresponding to the $L^{q}$-log-Sobolev inequality, it is useful to analyze the properties of the function $x\mapsto x\cdot\log(1/x)$.
 The proof of the following lemma is provided in Appendix A.
+Throughout, we adopt the convention that $0\cdot\log(0)=0$.
 Lemma 4.2.
@@
 Let $q\geq 1$. If the distribution $\pi$ satisfies $\mathrm{LSI}_{q}$ with constant $\mathtt{s}_{q}$, then $\pi$ satisfies a three-set isoperimetric inequality with functions
-$\Upsilon(t)=\frac{\mathtt{s}_{q}t^{q}}{1+\left(1+e^{-1}\right)\mathtt{s}_{q}t^{q}},\qquad\Psi(t)=\frac{t}{2}\cdot\log\left(\frac{2}{t}\right).$
+$\Upsilon(t)=\frac{\mathtt{s}_{q}t^{q}}{1+(1+e^{-1})\mathtt{s}_{q}t^{q}},\qquad\Psi(t)=\frac{t}{2}\cdot\log\left(\frac{2}{t}\right).$
 Proof.
@@
 $f(x)=\max\left\{0,1-\frac{\mathcal{D}(x,S_{1})}{\mathcal{D}(S_{1},S_{2})}\right\}.$
-As seen before, the map $x\mapsto\mathcal{D}(x,S_{1})$ is Lipschitz continuous with constant 1.
-It follows that the function $f$ is equal to 1 on $S_{1}$, vanishes on $S_{2}$, and is Lipschitz continuous on a subset of $S_{3}$, with the modulus of its gradient equal to $\mathcal{D}(S_{1},S_{2})^{-1}$; elsewhere on $S_{3}$, $f$ vanishes.
-In particular, $f(x)\in[0,1]$ for $x\in S_{3}$.
+As established above, the function $f$ is Lipschitz continuous with $\left\lvert\nabla f(x)\right\rvert\leq\mathcal{D}(S_{1},S_{2})^{-1}$ for almost every $x\in\mathbb{R}^{d}$.
+Moreover, $f$ takes the constant value 1 on $S_{1}$, vanishes on $S_{2}$, and satisfies $f(x)\in[0,1]$ for all $x\in S_{3}$.
 Applying (4) to $f$, we obtain
-$\mathbb{E}\left[\left\lvert f\right\rvert^{q}\cdot\log\left\lvert f\right\rvert^{q}\right]-\mathbb{E}\left[\left\lvert f\right\rvert^{q}\right]\cdot\log\mathbb{E}\left[\left\lvert f\right\rvert^{q}\right]\leq\frac{1}{\mathtt{s}_{q}}\cdot\mathbb{E}\left[\left\lvert\nabla f(x)\right\rvert^{q}\right]\leq\frac{\pi(S_{3})}{\mathtt{s}_{q}\cdot\mathcal{D}(S_{1},S_{2})^{q}}.$ (7)
-For $x\in[0,1]$, consider the function $\rho(x)=x\cdot\log(1/x)$. Using the convention that $00\cdot\log(0)=1\cdot\log(1)=$ 0, we can rewrite
-$\mathbb{E}\left[\left\lvert f\right\rvert^{q}\cdot\log\left\lvert f\right\rvert^{q}\right]={\int_{S_{3}}}\left\lvert f(x)\right\rvert^{q}\cdot\log\left(\left\lvert f(x)\right\rvert^{q}\right)\pi(\mathrm{d}x)=-{\int_{S_{3}}}\rho\left(\left\lvert f(x)\right\rvert^{q}\right)\pi(\mathrm{d}x)$
+$\mathbb{E}[\left\lvert f\right\rvert^{q}\cdot\log\left\lvert f\right\rvert^{q}]-\mathbb{E}[\left\lvert f\right\rvert^{q}]\cdot\log\mathbb{E}[\left\lvert f\right\rvert^{q}]\leq\frac{1}{\mathtt{s}_{q}}\cdot\mathbb{E}[\left\lvert\nabla f(x)\right\rvert^{q}]\leq\frac{\pi(S_{3})}{\mathtt{s}_{q}\cdot\mathcal{D}(S_{1},S_{2})^{q}}.$ (9)
+For $x\in[0,1]$, consider the function $\rho(x)=x\cdot\log(1/x)$. We can rewrite
+$\mathbb{E}[\left\lvert f\right\rvert^{q}\cdot\log\left\lvert f\right\rvert^{q}]={\int_{S_{3}}}\left\lvert f(x)\right\rvert^{q}\cdot\log(\left\lvert f(x)\right\rvert^{q})\pi(\mathrm{d}x)=-{\int_{S_{3}}}\rho(\left\lvert f(x)\right\rvert^{q})\pi(\mathrm{d}x)$
 and
-$\mathbb{E}\left[\left\lvert f\right\rvert^{q}\right]\cdot\log\mathbb{E}\left[\left\lvert f\right\rvert^{q}\right]$ $=\left(\pi(S_{1})+\int_{S_{3}}\left\lvert f(x)\right\rvert^{q}\pi(\mathrm{d}x)\right)\cdot\log\left(\pi(S_{1})+\int_{S_{3}}\left\lvert f(x)\right\rvert^{q}\pi(\mathrm{d}x)\right)$
+$\mathbb{E}[\left\lvert f\right\rvert^{q}]\cdot\log\mathbb{E}[\left\lvert f\right\rvert^{q}]$ $=\left(\pi(S_{1})+\int_{S_{3}}\left\lvert f(x)\right\rvert^{q}\pi(\mathrm{d}x)\right)\cdot\log\left(\pi(S_{1})+\int_{S_{3}}\left\lvert f(x)\right\rvert^{q}\pi(\mathrm{d}x)\right)$
 $=-\rho\left(\pi(S_{1})+\int_{S_{3}}\left\lvert f(x)\right\rvert^{q}\pi(\mathrm{d}x)\right).$
 By the first part of Lemma 4.2,
-$\mathbb{E}\left[\left\lvert f\right\rvert^{q}\cdot\log\left\lvert f\right\rvert^{q}\right]\geq-e^{-1}\cdot\pi(S_{3}).$ (8)
+$\mathbb{E}[\left\lvert f\right\rvert^{q}\cdot\log\left\lvert f\right\rvert^{q}]\geq-e^{-1}\cdot\pi(S_{3}).$ (10)
 Noting that $\mathbb{E}[\left\lvert f\right\rvert^{q}]\leq 1$, the second part of Lemma 4.2 gives
@@
 and thus
-$-\mathbb{E}\left[\left\lvert f\right\rvert^{q}\right]\cdot\log\mathbb{E}\left[\left\lvert f\right\rvert^{q}\right]\geq\pi(S_{1})\cdot\log\left(\frac{1}{\pi(S_{1})}\right)-\pi(S_{3}).$ (9)
+$\mathbb{E}[\left\lvert f\right\rvert^{q}]\cdot\log\mathbb{E}[\left\lvert f\right\rvert^{q}]\geq\pi(S_{1})\cdot\log\left(\frac{1}{\pi(S_{1})}\right)-\pi(S_{3}).$ (11)
 Inserting (10) and (11) into (9), we obtain
@@
 Combing these results, we have
-$\pi(S_{3})$ $\geq\frac{\mathcal{D}(S_{1},S_{2})^{q}}{\mathtt{s}_{q}^{-1}+(1+e^{-1})\cdot\mathcal{D}(S_{1},S_{2})^{q}}\cdot\max\left\{\pi(S_{1})\cdot\log\left(\frac{1}{\pi(S_{1})}\right),~{}\pi(S_{2})\cdot\log\left(\frac{1}{\pi(S_{2})}\right)\right\}$
-$\geq\frac{\mathcal{D}(S_{1},S_{2})^{q}}{\mathtt{s}_{q}^{-1}+(1+e^{-1})\cdot\mathcal{D}(S_{1},S_{2})^{q}}\cdot\min\left\{\pi(S_{1}),\pi(S_{2})\right\}\cdot\log\left(\frac{1}{\min\left\{\pi(S_{1}),\pi(S_{2})\right\}}\right)$
-$\geq\frac{\mathcal{D}(S_{1},S_{2})^{q}}{\mathtt{s}_{q}^{-1}+(1+e^{-1})\cdot\mathcal{D}(S_{1},S_{2})^{q}}\cdot\frac{\min\left\{\pi(S_{1}),\pi(S_{2})\right\}}{2}\cdot\log\left(\frac{2}{\min\left\{\pi(S_{1}),\pi(S_{2})\right\}}\right)$
+$\pi(S_{3})$ $\geq\frac{\mathcal{D}(S_{1},S_{2})^{q}}{\mathtt{s}_{q}^{-1}+(1+e^{-1})\cdot\mathcal{D}(S_{1},S_{2})^{q}}\cdot\max\left\{\pi(S_{1})\cdot\log\left(\frac{1}{\pi(S_{1})}\right),\pi(S_{2})\cdot\log\left(\frac{1}{\pi(S_{2})}\right)\right\}$
+$\geq\frac{\mathcal{D}(S_{1},S_{2})^{q}}{\mathtt{s}_{q}^{-1}+(1+e^{-1})\cdot\mathcal{D}(S_{1},S_{2})^{q}}\cdot\min\{\pi(S_{1}),\pi(S_{2})\}\cdot\log\left(\frac{1}{\min\{\pi(S_{1}),\pi(S_{2})\}}\right)$
+$\geq\frac{\mathcal{D}(S_{1},S_{2})^{q}}{\mathtt{s}_{q}^{-1}+(1+e^{-1})\cdot\mathcal{D}(S_{1},S_{2})^{q}}\cdot\frac{\min\{\pi(S_{1}),\pi(S_{2})\}}{2}\cdot\log\left(\frac{2}{\min\{\pi(S_{1}),\pi(S_{2})\}}\right)$
 The advantage of this last expression is that the function $x\mapsto x/2\cdot\log(2/x)$ is increasing everywhere on $(0,1/2)$, contrary to the function $x\mapsto x\cdot\log(1/x)$.
@@
 $\Upsilon(t)=\frac{\mathtt{s}_{q}t^{q}}{1+(1+e^{-1})\mathtt{s}_{q}t^{q}},\qquad\Psi(t)=\frac{t}{2}\cdot\log\left(\frac{2}{t}\right).$
-Notably, the $L^{q}$-Poincaré and $L^{q}$-log-Sobolev inequalities both lead to three-set isoperimetric inequalities with $\Upsilon(t)\sim\mathtt{p}_{q}t^{q}$ as $t\to 0$.
-Here the sign $\sim$ means that the two quantities are asymptotic, meaning their ratio converges to a constant as $t\to 0$.
+Notably, the $L^{q}$-Poincaré and $L^{q}$-log-Sobolev inequalities respectively lead to three-set isoperimetric inequalities with $\Upsilon(t)\sim\mathtt{p}_{q}t^{q}/2^{2q}$ and $\Upsilon(t)\sim\mathtt{s}_{q}t^{q}$ as $t\to 0$.
+Here, the sign $\sim$ means that the two quantities are equivalent as $t\to 0$, in the sense that their ratio converges to a finite constant.
 This regime is particularly relevant in high-dimensional settings, where $t$ typically scales inversely with the dimension, as per the discussion following Corollary 3.12.
-In contrast, when $\pi$ is log-concave, we have $\Upsilon(t)\sim t$ as $t\to 0$ (Andrieu et al., 2024, Subsection 3.2).
+In contrast, when $\pi$ is log-concave, $\Upsilon(t)\sim t$ as $t\to 0$ (Andrieu et al., 2024, Subsection 3.2).
+In many settings, one can only verify a close coupling condition on a high probability region $K\subsetneq\mathbb{R}^{d}$, which leads to lower bounds on the $s$-conductance (or restricted conductance) rather than the usual conductance; see, for example, Wu et al.
+(2022); Chen and Gatmiry (2023); Chewi et al.
+(2021); Dwivedi et al.
+(2019); Mangoubi and Vishnoi (2019); Mou et al.
+(2022a); Chen et al.
+(2020).
+In certain cases, obtaining such bounds requires working with a three-set isoperimetric inequality for the distribution $\pi$ truncated to $K$.
+For log-concave distributions, it is known that truncation preserves three-set isoperimetric inequalities; see Chen et al.
+(2020, Lemma 15) and Dwivedi et al.
+(2019, Proof of Lemma 6), both relying on Cousins and Vempala (2014, Theorem 4.4).
+Beyond the log-concave setting, however, it is unclear whether three-set isoperimetric inequalities are preserved under truncation.
+Extending the three-set isoperimetric inequalities associated with global Poincaré or log-Sobolev inequalities to truncated distributions appears delicate and is left for future work.
 4.2 Close coupling conditions for the Gibbs kernels
+In what follows, let $d\geq 2$.
 Here, we establish close coupling conditions for the systematic-scan and random-scan Gibbs kernels, assuming that the conditional distributions of the target distribution $\pi$ are TV continuous (as per Definition 2.5).
@@
 Hence, $P_{\mathrm{SS}}$ satisfies a $(\delta,\varepsilon)$-close coupling condition with
-$\delta=\left(Md\right)^{-1/\beta},\qquad\varepsilon=e^{-1}.$
+$\delta=\frac{1}{(Md)^{1/\beta}},\qquad\varepsilon=\frac{1}{4}.$
 Proof.
@@
   • Finally, set $U=U^{\left(d\right)}$, $V=V^{\left(d\right)}$.
-By construction, we have $U\sim P_{\mathrm{SS}}(x,\cdot)$ and $V\sim P_{\mathrm{SS}}(y,\cdot)$.
-Moreover, observe that for all $k\in[d]$, we have $U^{\left(k\right)}=\left(U_{1:k},x_{k+1:d}\right)$ and $V^{\left(k\right)}=\left(V_{1:k},y_{k+1:d}\right)$.
+By construction, $U\sim P_{\mathrm{SS}}(x,\cdot)$ and $V\sim P_{\mathrm{SS}}(y,\cdot)$.
+Moreover, observe that for all $k\in[d]$, $U^{\left(k\right)}=\left(U_{1:k},x_{k+1:d}\right)$ and $V^{\left(k\right)}=\left(V_{1:k},y_{k+1:d}\right)$.
 Using this, (12) can be rewritten as
@@
 We now turn our attention towards obtaining a close coupling for the random-scan Gibbs sampler.
-However, one issue immediately becomes apparent: when $x,y\in\mathbb{R}^{d}$ do not share a coordinate (i.e., when $x-y$ has a non-zero component in each coordinate direction), the supports of $P_{\mathrm{RS}}(x,\cdot)$ and $P_{\mathrm{RS}}(y,\cdot)$ are disjoint.
+However, one issue immediately becomes apparent: when $x,y\in\mathbb{R}^{d}$ do not share a coordinate (i.e., when xyx-y has a non-zero component in each coordinate direction), the supports of $P_{\mathrm{RS}}(x,\cdot)$ and $P_{\mathrm{RS}}(y,\cdot)$ are disjoint.
 In this case,
 $\left\|P_{\mathrm{RS}}(x,\cdot)-P_{\mathrm{RS}}(y,\cdot)\right\|_{\textsc{TV}}=1.$
-[…]
+To circumvent this issue, we adopt a similar approach to Narayanan and Srivastava (2022) by establishing a close-coupling condition for the iterated kernel $P_{\mathrm{RS}}^{N}$.
+Notably, when $N\geq d$, $P_{\mathrm{RS}}^{N}(x,\cdot)$ is globally supported for all $x\in\mathbb{R}^{d}$.
+It turns out that $N\sim d\cdot\log d$ as $d\to\infty$ is indeed the correct scaling in order to get a non-degenerate close coupling condition for the iterated random-scan Gibbs kernel, as formalized in the next lemma.
+Let $\lceil\cdot\rceil$ denote the ceiling function.
+The proof can be found in Appendix A.
 Lemma 4.5.
@@
 Denote by $\Xi$ the set of unique values attained by $I_{1},\dots,I_{N}$.
-If $N=\lceil 4d\cdot\log d\rceil$, then
+If $N=\lceil 2d\cdot\log d\rceil$, then
 $\mathbb{P}(\Xi=[d])\geq\frac{1}{2}.$
-Lemma 4.5 is closely related to the classic coupon collector problem in probability (Erdős and Rényi, 1961). Using the same notation, observe that when $N=d$, we have
+Lemma 4.5 is closely related to the classic coupon collector problem in probability (Erdős and Rényi, 1961). Using the same notation, observe that when $N=d$,
 $\mathbb{P}(\Xi=[d])=\frac{d!}{d^{d}}\sim\frac{\sqrt{d}}{e^{d}}$
-as $d\to\infty$, by Stirling’s formula. The probability of observing all $d$ unique values in exactly $d$ draws thus decreases exponentially with the dimension $d$.
+as $d\to\infty$. This follows from Stirling’s formula. The probability of observing all $d$ unique values in exactly $d$ draws thus decreases exponentially with the dimension $d$.
 Proposition 4.6.
-Suppose that the conditional distributions of $\pi$ are $(M,\beta)$-TV continuous. Let $N_{d}=\lceil 4d\cdot\log d\rceil$. Then,
-$\left\|P_{\mathrm{RS}}^{N_{d}}(x,\cdot)-P_{\mathrm{RS}}^{N_{d}}(y,\cdot)\right\|_{\textsc{TV}}\leq 1-\frac{1}{2}(1-M\left\lvert x-y\right\rvert^{\beta})^{N_{d}}\qquad\forall x,y\in\mathbb{R}^{d}.$
+Suppose that the conditional distributions of $\pi$ are $(M,\beta)$-TV continuous. Let $N_{d}=\lceil 2d\cdot\log d\rceil$. Then,
+$\left\|P_{\mathrm{RS}}^{N_{d}}(x,\cdot)-P_{\mathrm{RS}}^{N_{d}}(y,\cdot)\right\|_{\textsc{TV}}\leq 1-\frac{1}{2}\left(1-M\left\lvert x-y\right\rvert^{\beta}\right)^{N_{d}}\qquad\forall x,y\in\mathbb{R}^{d}.$
 Hence, $P_{\mathrm{RS}}^{N_{d}}$ satisfies a $(\delta,\varepsilon)$-close coupling condition with
-$\delta=\left(MN_{d}\right)^{-1/\beta},\qquad\varepsilon=\left(2e\right)^{-1}.$
+$\delta=\frac{1}{(MN_{d})^{1/\beta}},\qquad\varepsilon=\frac{1}{8}.$
 In particular, $\delta\geq(3Md\cdot\log d)^{-1/\beta}$.
@@
 When $x$ and $y$ differ in at least two coordinates, the event $\left\{U=V\right\}$ has a non-zero probability of occurring only if $\Xi=[d]$, i.e., when all coordinates have been updated at least once by $P_{\mathrm{RS}}^{N_{d}}$.
-More generally, for all $x,y\in\mathbb{R}^{d}$, we have
+More generally, for all $x,y\in\mathbb{R}^{d}$,
 $\mathbb{P}(U=V)\geq\mathbb{P}(\{U=V\}\cap\{\Xi=\left[d\right]\})=\mathbb{P}(\Xi=[d])\cdot\mathbb{P}(U=V\mid\Xi=\left[d\right]).$
-By Lemma 4.5, $\mathbb{P}(\Xi=\left[d\right])\geq 1/2$ when $N_{d}=\left\lceil 4d\cdot\log d\right\rceil$. Hence, for this value of $N_{d}$, we get
+By Lemma 4.5, $\mathbb{P}(\Xi=\left[d\right])\geq 1/2$ when $N_{d}=\lceil 2d\cdot\log d\rceil$. Hence, for this value of $N_{d}$, we get
 $\left\|P_{\mathrm{RS}}^{N_{d}}(x,\cdot)-P_{\mathrm{RS}}^{N_{d}}(y,\cdot)\right\|_{\textsc{TV}}$ $\leq\mathbb{P}(U\neq V)$
@@
 $\leq\mathbb{E}_{U^{\left(k-1\right)},V^{\left(k-1\right)},I_{k}}\left[M\left\lvert U^{\left(k-1\right)}-V^{\left(k-1\right)}\right\rvert^{\beta}~\middle|~\bigcap_{l=1}^{k-1}\left\{W_{l}=W_{l}^{\prime}\right\}\cap\left\{\Xi=\left[d\right]\right\}\right]$
-$\leq M\left\lvert x-y\right\rvert^{\beta},$ (15)
-where the subscript in the expectation indicates that it is taken with respect to the joint distribution of $U^{(k-1)},V^{(k-1)},I_{k}$. Combining (13), (14), (15), we finally obtain
+$\leq M\left\lvert x-y\right\rvert^{\beta},$
+where the subscript in the expectation indicates that it is taken with respect to the joint distribution of $U^{(k-1)},V^{(k-1)},I_{k}$. Recalling (15) and (16), we finally obtain
 $\left\|P_{\mathrm{RS}}^{N_{d}}(x,\cdot)-P_{\mathrm{RS}}^{N_{d}}(y,\cdot)\right\|_{\textsc{TV}}\leq 1-\frac{1}{2}\left(1-M\left\lvert x-y\right\rvert^{\beta}\right)^{N_{d}}.$
@@
 By combining the three-set isoperimetric inequalities from Subsection 4.1 with the close coupling conditions from Subsection 4.2, we can bound the conductance of the systematic-scan and random-scan Gibbs kernels.
-These results apply to target distributions that satisfy a $L^{q}$-Poincaré or $L^{q}$-log-Sobolev inequality, and whose conditional distributions are TV continuous.
+These results apply to target distributions that satisfy an $L^{q}$-Poincaré or $L^{q}$-log-Sobolev inequality, and whose conditional distributions are TV continuous.
 Throughout this subsection, we use $\lesssim$ to denote inequalities that hold up to a universal constant factor.
@@
 Suppose that the conditional distributions of $\pi$ are $(M,\beta)$-TV continuous.
-Then, the conductance of the systematic-scan Gibbs kernel is bounded by
-$\Phi(P_{\mathrm{SS}})\gtrsim\min\left\{1,\frac{\mathtt{p}_{q}}{\left(Md\right)^{q/\beta}}\right\}.$
+Then, the conductance of the lazy systematic-scan Gibbs kernel is bounded by
+$\Phi(\widetilde{P}_{\mathrm{SS}})\gtrsim\min\left\{1,\frac{\mathtt{p}_{q}}{2^{2q}(Md)^{q/\beta}}\right\}.$
 If the chain is initialized at a $\Omega$-warm start, then the $\zeta$-mixing time is bounded by
-$\tau(\zeta,P_{\mathrm{SS}})\lesssim\max\left\{1,\frac{\left(Md\right)^{2q/\beta}}{\mathtt{p}_{q}^{2}}\right\}\cdot\log\left(\frac{\sqrt{\Omega}}{\zeta}\right).$
+$\tau(\zeta,\widetilde{P}_{\mathrm{SS}})\lesssim\max\left\{1,\frac{2^{4q}(Md)^{2q/\beta}}{\mathtt{p}_{q}^{2}}\right\}\cdot\log\left(\frac{\sqrt{\Omega}}{\zeta}\right).$
 Proof.
-The bound on the conductance follows from combining Corollary 3.11, Proposition 4.1, and Proposition 4.4. The bound on the mixing time is obtained via Corollary 3.5. ∎
+Combining Corollary 3.12, Proposition 4.1, and Proposition 4.4 yields a lower bound on the conductance of $P_{\mathrm{SS}}$.
+By (5), this directly implies the stated bound on the conductance of $\widetilde{P}_{\mathrm{SS}}$.
+Since $\widetilde{P}_{\mathrm{SS}}$ is lazy, the bound on the mixing time follows from Corollary 3.7.
+∎
 Theorem 4.8.
@@
 Suppose that the conditional distributions of $\pi$ are $(M,\beta)$-TV continuous.
-Let $N_{d}=\left\lceil 4d\cdot\log d\right\rceil$.
+Let $N_{d}=\lceil 2d\cdot\log d\rceil$.
 Then, the conductance of the iterated random-scan Gibbs kernel is bounded by
-$\Phi\left(P^{N_{d}}_{\mathrm{RS}}\right)\gtrsim\min\left\{1,\frac{\mathtt{p}_{q}}{\left(MN_{d}\right)^{q/\beta}}\right\}.$
+$\Phi\left(P^{N_{d}}_{\mathrm{RS}}\right)\gtrsim\min\left\{1,\frac{\mathtt{p}_{q}}{2^{2q}(MN_{d})^{q/\beta}}\right\}.$
 The spectral gap satisfies
-$\lambda_{2}\left(P^{N_{d}}_{\mathrm{RS}}\right)\gtrsim\min\left\{1,\frac{\mathtt{p}_{q}^{2}}{\left(MN_{d}\right)^{2q/\beta}}\right\}.$
+$\lambda_{2}\left(P^{N_{d}}_{\mathrm{RS}}\right)\gtrsim\min\left\{1,\frac{\mathtt{p}_{q}^{2}}{2^{4q}(MN_{d})^{2q/\beta}}\right\}.$
 If the chain is initialized at a $\Omega$-warm start, then the $\zeta$-mixing time is bounded by
-$\tau\left(\zeta,P^{N_{d}}_{\mathrm{RS}}\right)\lesssim\max\left\{1,\frac{\left(MN_{d}\right)^{2q/\beta}}{\mathtt{p}_{q}^{2}}\right\}\cdot\log\left(\frac{\sqrt{\Omega}}{\zeta}\right).$
+$\tau\left(\zeta,P^{N_{d}}_{\mathrm{RS}}\right)\lesssim\max\left\{1,\frac{2^{4q}(MN_{d})^{2q/\beta}}{\mathtt{p}_{q}^{2}}\right\}\cdot\log\left(\frac{\sqrt{\Omega}}{\zeta}\right).$
 Proof.
-The conductance bound follows from Corollary 3.11, Proposition 4.1, and Proposition 4.6.
-The spectral gap is then bounded using Theorem 3.7, while the mixing time bound is obtained via Corollary 3.5.
+Combining Corollary 3.12, Proposition 4.1, and Proposition 4.6 yields the lower bound on the conductance.
+The spectral gap is bounded using Theorem 3.8.
+Since $P_{\mathrm{RS}}$ is reversible and positive semi-definite, the same holds for $P^{N_{d}}_{\mathrm{RS}}$.
+The bound on the mixing time then follows from Corollary 3.7.
 ∎
 We now focus on bounding the conductance and spectral gap of the random-scan Gibbs kernel $P_{\mathrm{RS}}$.
-As noted by Narayanan and Srivastava (2022, Lemma 5.3), for any reversible Markov kernel $P$ and for all $N\geq 1$, we have
+Recall that $P_{\mathrm{RS}}$ is reversible and positive semi-definite.
+As noted by Narayanan and Srivastava (2022, Lemma 5.3), for any reversible Markov kernel $P$ and for all $N\geq 1$,
 $\Phi(P)\geq\frac{1}{N}\cdot\Phi\left(P^{N}\right).$
-Let $\sigma(P)$ denote the second largest eigenvalue of $P$, so that $\lambda_{2}(P)=1-\sigma(P)$. Notably, $\sigma\left(P^{N}\right)=\sigma(P)^{N}$, which implies that
+Let $\sigma(P)$ denote the second largest eigenvalue of $P$, so that $\lambda_{2}(P)=1-\sigma(P)$. If $P$ is positive semi-definite, $\sigma\left(P^{N}\right)=\sigma(P)^{N}$, which implies that
 $\lambda_{2}(P)$ $=1-\sigma(P)$
@@
 Under the conditions of Theorem 4.8, the conductance of the random-scan Gibbs kernel is bounded by
-$\Phi(P_{\mathrm{RS}})\gtrsim\frac{1}{d\cdot\log d}\cdot\min\left\{1,\frac{\mathtt{p}_{q}}{M^{q/\beta}(d\cdot\log d)^{q/\beta}}\right\}.$
+$\Phi(P_{\mathrm{RS}})\gtrsim\frac{1}{d\cdot\log d}\cdot\min\left\{1,\frac{\mathtt{p}_{q}}{2^{2q}(3Md\cdot\log d)^{q/\beta}}\right\}.$
 The spectral gap satisfies
-$\lambda_{2}(P_{\mathrm{RS}})\gtrsim\frac{1}{d\cdot\log d}\cdot\min\left\{1,\frac{\mathtt{p}_{q}^{2}}{M^{2q/\beta}(d\cdot\log d)^{2q/\beta}}\right\}.$
+$\lambda_{2}(P_{\mathrm{RS}})\gtrsim\frac{1}{d\cdot\log d}\cdot\min\left\{1,\frac{\mathtt{p}_{q}^{2}}{2^{4q}(3Md\cdot\log d)^{2q/\beta}}\right\}.$
 The $\zeta$-mixing time is bounded by
-$\tau(\zeta,P_{\mathrm{RS}})\lesssim d\cdot\log d\cdot\max\left\{1,\frac{M^{2q/\beta}(d\cdot\log d)^{2q/\beta}}{\mathtt{p}_{q}^{2}}\right\}\cdot\log\left(\frac{\sqrt{\Omega}}{\zeta}\right).$
+$\tau(\zeta,P_{\mathrm{RS}})\lesssim d\cdot\log d\cdot\max\left\{1,\frac{2^{4q}(3Md\cdot\log d)^{2q/\beta}}{\mathtt{p}_{q}^{2}}\right\}\cdot\log\left(\frac{\sqrt{\Omega}}{\zeta}\right).$
 These results demonstrate that the systematic-scan and random-scan Gibbs samplers exhibit polynomial mixing time dependence on the dimension $d$ for a large class of target distributions.
@@
 This additional $\log d$ factor highlights a modest penalty compared to systematic scan for well-behaved target distributions.
-This aligns with prior heuristic observations about the efficiency of these sampling schemes (Diaconis, 2013) and arises naturally from spectral bounds on mixing times (Gaitonde and Mossel, 2024).
+This aligns with prior heuristic observations about the efficiency of these sampling schemes (Diaconis, 2013) and arises naturally from spectral bounds on mixing times (Gaitonde and Mossel, 2026).
 By Proposition 4.3, all results in this subsection extend to distributions satisfying $\mathrm{LSI}_{q}$ instead of $\mathrm{PI}_{q}$, with the conductance bounds remaining equivalent up to constant factors.
@@
 However, improvements may be possible when using a different norm to define the mixing time.
-For instance, in $\chi^{2}$-divergence, Chen et al.
-(2020, Theorem 1) demonstrated that the logarithmic factor in $\Psi$ from Proposition 4.3 can be exploited to improve the mixing time’s dependence on the warm start parameter (see Corollary 3.5), reducing it from $\log$ to $\log\log$.
+One possible approach is to adapt spectral and conductance profile methods under an $L^{q}$-log-Sobolev inequality, following Andrieu et al.
+(2024); Chen et al.
+(2020).
+In $\chi^{2}$-distance, these methods exploit the logarithmic factor in the function $\Psi$ from Proposition 4.3 to improve the dependence of the mixing time on the warm start parameter, reducing it from $\log$ to $\log\log$ (see Corollary 3.7).
+At a high level, this would require sharpening Corollary 3.12 along the lines of Andrieu et al.
+(2024, Corollary 16) to obtain tighter control of the conductance profile, and then plugging this refinement into the mixing time bounds of Andrieu et al.
+(2024, Theorem 8).
+Implementing this approach would require balancing several interacting quantities, such as the order of the Poincaré or log-Sobolev inequality, the form of the function $\Psi$ in the associated three-set isoperimetric inequality, and the exponent in the regularity assumption for the conditional distributions (Definition 2.5), resulting in less transparent expressions.
+Moreover, spectral profile methods rely on reversibility and do not improve upon standard spectral gap bounds for distributions that satisfy only a Poincaré inequality (Andrieu et al., 2024, Subsection 3.3).
+For these reasons, we do not pursue this direction further here.
 5 Extensions and applications
@@
 where the function $U:\mathbb{R}^{d}\to\mathbb{R}$ is known as the potential.
-Many previous works assume that the potential function is either Lipschitz continuous (Lehec, 2023) or smooth (see, e.g., Ascolani et al., 2024a).
-We show that TV continuity of the conditional distributions of $\pi$ (Definition 2.5) is a strictly weaker condition than these regularity assumptions on $U$.
+Many previous works assume that the potential function is either Lipschitz continuous (Lehec, 2023) or smooth (see, for example, Ascolani et al., 2026a).
 Definition 5.1.
@@
 Equivalently, the distribution $\pi$ is said to be $L$-log-smooth.
-To establish our result, we first demonstrate that log-Lipschitz continuity and log-smoothness extend to the marginal distributions of $\pi$. The proof of the following lemma is located in Appendix A.
+We show that TV continuity of the conditional distributions of $\pi$ (Definition 2.5) can be obtained under the above regularity assumptions on $U$.
+To establish this, we first demonstrate that log-Lipschitz continuity and log-smoothness are preserved under marginalization, provided that the conditional distributions of $\pi$ satisfy a uniform $L^{2}$-Poincaré inequality.
+The proof of the following lemma is located in Appendix A.
 Lemma 5.3.
 The following properties hold:
-  • If the distribution $\pi$ is $L$-log-Lipschitz continuous, then all marginal distributions of $\pi$ are also $L$-log-Lipschitz continuous.
-  • If the distribution $\pi$ is $L$-log-smooth, then all marginal distributions of $\pi$ are also $L$-log-smooth.
-Lemma 5.3 complements Saumard and Wellner (2014, Theorems 3.3 and 3.8), who proved that (strong) log-concavity is also inherited by the marginal distributions of $\pi$.
+  • Suppose that the distribution $\pi$ is $L$-log-Lipschitz continuous. Then, all marginal distributions of $\pi$ are also $L$-log-Lipschitz continuous.
+  • Suppose that the distribution $\pi$ is $L$-log-smooth and for all $k\in[d]$ and $x\in\mathbb{R}^{d}$, the conditional distribution $\pi(\cdot\mid x_{-k})$ satisfies $\mathrm{PI}_{2}$ with uniform constant $\mathtt{c}_{2}$.
+Then, all (d1)(d-1)-dimensional marginal distributions are $L^{\prime}$-log-smooth, where
+    $L^{\prime}=L+\frac{L^{2}}{\mathtt{c}_{2}}.$
+Lemma 5.3 complements the classical result that (strong) log-concavity is preserved under marginalization, as shown in Prékopa (1973); Brascamp and Lieb (1976); see also Saumard and Wellner (2014) for a survey.
+The assumption that the conditional distributions of $\pi$ satisfy a uniform $L^{2}$-Poincaré inequality has been widely used in the study of two-scale criteria for functional inequalities (Otto and Reznikoff, 2007; Lelièvre, 2009; Grunewald et al., 2009), in covariance estimates (Menz, 2014b), and in the analysis of spin systems (see, for example, Menz, 2014a).
+As a simple sanity check, the above assumption holds when the target distribution is strongly log-concave.
+More precisely, if $\pi$ is strongly log-concave with constant $m>0$, then each one-dimensional conditional distribution is strongly log-concave with the same constant, and therefore satisfies a uniform $L^{2}$-Poincaré inequality (see the discussion in Subsection 2.2).
+Indeed, the potential function of the conditional distribution $\pi(\cdot\mid x_{-k})$ is given by
+$x_{k}\mapsto U(x_{k},x_{-k}),$
+where $x_{-k}$ is held fixed.
+Strong log-concavity of $\pi$ with constant $m>0$ is equivalent to the Hessian bound $\nabla^{2}U(x)\succeq mI_{d}$, where $I_{d}$ denotes the $d$-dimensional identity matrix and $\succeq$ the Loewner order.
+In particular, this implies that
+$\frac{\partial^{2}U}{\partial x^{2}_{k}}(x_{k},x_{-k})\geq m,$
+which is exactly the condition that $\pi(\cdot\mid x_{-k})$ is strongly log-concave with constant $m$.
 Lemma 5.4.
@@
 For simplicity, we redefine the variables as $x_{-k}=s$, $y_{-k}=t$.
-The joint distribution can be rewritten as $\pi(\theta,s)\propto e^{-U(\theta,s)}.$ From this, the conditional distribution is expressed as
-$\pi(\theta|s)=\frac{\pi(\theta,s)}{\pi(s)}=\frac{e^{-U(\theta,s)}}{\int e^{-U(\psi,s)}\mathrm{d}\psi}=\frac{e^{-U(\theta,s)}}{e^{-V(s)}},$
+The joint distribution can be rewritten as $\pi(\theta,s)\propto\exp({-U(\theta,s)})$.
+From this, the conditional distribution is expressed as
+$\pi(\theta\mid s)=\frac{\pi(\theta,s)}{\pi(s)}=\frac{e^{-U(\theta,s)}}{\int e^{-U(\psi,s)}\mathrm{d}\psi}=\frac{e^{-U(\theta,s)}}{e^{-V(s)}},$
 where
@@
 $\left\|\pi(\cdot\mid s)-\pi(\cdot\mid t)\right\|_{\textsc{TV}}$ $\leq\sqrt{\frac{1}{2}\cdot\mathrm{KL}(\pi(\cdot\mid s),\pi(\cdot\mid t))}$
-$\leq\sqrt{L\left\lvert s-t\right\rvert}.$
+$\leq\sqrt{L\left\lvert s-t\right\rvert}$$\qed$
 Lemma 5.5.
-Suppose that the distribution $\pi$ is $L$-log-smooth. Then, the conditional distributions of $\pi$ are $(\sqrt{L},1)$-TV continuous.
+Suppose that the distribution $\pi$ is $L$-log-smooth and for all $k\in[d]$ and $x\in\mathbb{R}^{d}$, the conditional distribution $\pi(\cdot\mid x_{-k})$ satisfies $\mathrm{PI}_{2}$ with uniform constant $\mathtt{c}_{2}$.
+Then, the conditional distributions of $\pi$ are $(M^{\prime},1)$-TV continuous, where
+$M^{\prime}=\sqrt{\frac{L}{2}+\frac{L^{2}}{4\mathtt{c}_{2}}}.$
 Proof.
 We proceed similarly to Lemma 5.4. Using the same notation, recall that
-$\mathrm{KL}(\pi(\cdot\mid s),\pi(\cdot\mid t))=\int(U(\theta,t)-U(\theta,s))\pi(\mathrm{d}\theta\mid s)+V(s)-V(t).$
-Observe that $\nabla V(s)=\int\nabla_{s}U(\theta,s)\pi(\mathrm{d}\theta\mid s).$ It follows that
-$\mathrm{KL}(\pi(\cdot\mid s),\pi(\cdot\mid t))$ $=\int(U(\theta,t)-U(\theta,s))\pi(\mathrm{d}\theta\mid s)+V(s)-V(t)$
-$\leq\int\left(\nabla_{s}U(\theta,s)^{T}(t-s)+\frac{L}{2}\left\lvert s-t\right\rvert^{2}\right)\pi(\mathrm{d}\theta\mid s)$
-$\quad+\nabla V(t)^{T}(s-t)+\frac{L}{2}\left\lvert s-t\right\rvert^{2}$
-$\leq\int(\nabla_{s}U(\theta,s)-\nabla_{t}U(\theta,t))^{T}(t-s)\pi(\mathrm{d}\theta\mid s)+L\left\lvert s-t\right\rvert^{2}$
-$\leq 2L\left\lvert s-t\right\rvert^{2}.$
+$\mathrm{KL}(\pi(\cdot\mid s),\pi(\cdot\mid t))=\int(U(\theta,t)-U(\theta,s))\pi(\mathrm{d}\theta\mid s)+V(s)-V(t).$ (17)
+We now bound both differences on the right-hand side of (17). By $L$-smoothness of the function $U$,
+$U(\theta,t)\leq U(\theta,s)+\nabla_{s}U(\theta,s)^{T}(t-s)+\frac{L}{2}\left\lvert s-t\right\rvert^{2},$ (18)
+where $\nabla_{s}U(\theta,s)$ represents the gradient of $U(\theta,s)$ with respect to the coordinates of $s$.
+Moreover, by Lemma 5.3, the function $V$ is $L^{\prime}$-smooth with $L^{\prime}=L+L^{2}/\mathtt{c}_{2}$.
+Therefore,
+$V(t)\geq V(s)+\nabla V(s)^{T}(t-s)-\frac{L^{\prime}}{2}\left\lvert s-t\right\rvert^{2}.$ (19)
+Observe that $\nabla V(s)=\int\nabla_{s}U(\theta,s)\pi(\mathrm{d}\theta\mid s)$. Substituting (18) and (19) into (17), we obtain
+$\mathrm{KL}(\pi(\cdot\mid s),\pi(\cdot\mid t))$ $\leq\int\nabla_{s}U(\theta,s)^{T}(t-s)\pi(\mathrm{d}\theta\mid s)-\nabla V(s)^{T}(t-s)+\frac{L+L^{\prime}}{2}\left\lvert s-t\right\rvert^{2}$
+$=\left(L+\frac{L^{2}}{2\mathtt{c}_{2}}\right)\left\lvert s-t\right\rvert^{2}.$
 We conclude by applying Pinsker’s inequality:
 $\left\|\pi(\cdot\mid s)-\pi(\cdot\mid t)\right\|_{\textsc{TV}}$ $\leq\sqrt{\frac{1}{2}\cdot\mathrm{KL}(\pi(\cdot\mid s),\pi(\cdot\mid t))}$
-$\leq\sqrt{L}\left\lvert s-t\right\rvert.$
+$\leq\sqrt{\frac{L}{2}+\frac{L^{2}}{4\mathtt{c}_{2}}}\left\lvert s-t\right\rvert$.
 The class of log-smooth and (strongly) log-concave target distributions is widely studied in the analysis of MCMC mixing times; see Ascolani et al.
@@
 Corollary 5.6.
-Consider a $L$-log-smooth target distribution $\pi$ that satisfies $\mathrm{PI}_{1}$. Then, the conductance of the systematic-scan Gibbs kernel is bounded by
-$\Phi(P_{\mathrm{SS}})\gtrsim\min\left\{1,\frac{\mathtt{p}_{1}}{\sqrt{L}d}\right\}.$
+Consider a $L$-log-smooth target distribution $\pi$ that satisfies $\mathrm{PI}_{1}$ with constant $\mathtt{p}_{1}$.
+Suppose that for all $k\in[d]$ and $x\in\mathbb{R}^{d}$, the conditional distribution $\pi(\cdot\mid x_{-k})$ satisfies $\mathrm{PI}_{2}$ with uniform constant $\mathtt{c}_{2}$.
+Then, the conductance of the lazy systematic-scan Gibbs kernel is bounded by
+$\Phi(\widetilde{P}_{\mathrm{SS}})\gtrsim\min\left\{1,\frac{\sqrt{\mathtt{c}_{2}}\mathtt{p}_{1}}{\sqrt{\mathtt{c}_{2}L+L^{2}}d}\right\}.$
 If the chain is initialized at a $\Omega$-warm start, then the $\zeta$-mixing time is bounded by
-$\tau(\zeta,P_{\mathrm{SS}})\lesssim\max\left\{1,\frac{Ld^{2}}{\mathtt{p}_{1}^{2}}\right\}\cdot\log\left(\frac{\sqrt{\Omega}}{\zeta}\right).$
-However, the mixing time bound in Corollary 4.7 appears to be loose when $\pi$ is a Gaussian distribution, as suggested by comparisons with Roberts and Sahu (1997).
+$\tau(\zeta,\widetilde{P}_{\mathrm{SS}})\lesssim\max\left\{1,\frac{(\mathtt{c}_{2}L+L^{2})d^{2}}{\mathtt{c}_{2}\mathtt{p}_{1}^{2}}\right\}\cdot\log\left(\frac{\sqrt{\Omega}}{\zeta}\right).$
+However, the bounds in Corollary 4.7 appear to be loose when $\pi$ is a Gaussian distribution, as suggested by comparisons with Roberts and Sahu (1997).
 Corollary 5.7.
-Consider a $L$-log-smooth target distribution $\pi$ that satisfies $\mathrm{PI}_{1}$. Then, the conductance of the random-scan Gibbs kernel is bounded by
-$\Phi(P_{\mathrm{RS}})\gtrsim\frac{1}{d\cdot\log d}\cdot\min\left\{1,\frac{\mathtt{p}_{1}}{\sqrt{L}d\cdot\log d}\right\}.$
+Consider a $L$-log-smooth target distribution $\pi$ that satisfies $\mathrm{PI}_{1}$ with constant $\mathtt{p}_{1}$.
+Suppose that for all $k\in[d]$ and $x\in\mathbb{R}^{d}$, the conditional distribution $\pi(\cdot\mid x_{-k})$ satisfies $\mathrm{PI}_{2}$ with uniform constant $\mathtt{c}_{2}$.
+Then, the conductance of the random-scan Gibbs kernel is bounded by
+$\Phi(P_{\mathrm{RS}})\gtrsim\frac{1}{d\cdot\log d}\cdot\min\left\{1,\frac{\sqrt{\mathtt{c}_{2}}\mathtt{p}_{1}}{\sqrt{\mathtt{c}_{2}L+L^{2}}d\cdot\log d}\right\}.$
 The spectral gap satisfies
-$\lambda_{2}(P_{\mathrm{RS}})\gtrsim\frac{1}{d\cdot\log d}\cdot\min\left\{1,\frac{\mathtt{p}_{1}^{2}}{L(d\cdot\log d)^{2}}\right\}.$
+$\lambda_{2}(P_{\mathrm{RS}})\gtrsim\frac{1}{d\cdot\log d}\cdot\min\left\{1,\frac{\mathtt{c}_{2}\mathtt{p}_{1}^{2}}{(\mathtt{c}_{2}L+L^{2})(d\cdot\log d)^{2}}\right\}.$
 If the chain is initialized at a $\Omega$-warm start, then the $\zeta$-mixing time is bounded by
-$\tau(\zeta,P_{\mathrm{RS}})\lesssim d\cdot\log d\cdot\max\left\{1,\frac{L(d\cdot\log d)^{2}}{\mathtt{p}_{1}^{2}}\right\}\cdot\log\left(\frac{\sqrt{\Omega}}{\zeta}\right).$
-Unfortunately, the spectral gap bound in Corollary 4.9 is suboptimal when $\pi$ is strongly log-concave and $L$-log-smooth. In such cases, Ascolani et al. (2024a, Corollary 3.6) established that
+$\tau(\zeta,P_{\mathrm{RS}})\lesssim d\cdot\log d\cdot\max\left\{1,\frac{(\mathtt{c}_{2}L+L^{2})(d\cdot\log d)^{2}}{\mathtt{c}_{2}\mathtt{p}_{1}^{2}}\right\}\cdot\log\left(\frac{\sqrt{\Omega}}{\zeta}\right).$
+Unfortunately, the spectral gap bound in Corollary 4.9 is suboptimal when $\pi$ is strongly log-concave and $L$-log-smooth. In such cases, Ascolani et al. (2026a, Corollary 3.7) established that
 $\lambda_{2}(P_{\mathrm{RS}})\gtrsim\frac{\mathtt{p}_{1}^{2}}{Ld}.$ (20)
@@
 Furthermore, we believe that the lower bound on the spectral gap in Theorem 3.8 is not tight for the Gibbs sampler.
-This belief is supported by the case of the random walk on the $d$-dimensional hypercube, where Trevisan (2013, Chapter 5) showed that the conductance and spectral gap are of the same order, making the first bound from Theorem 3.7 loose in this setting.
+This belief is supported by the case of the random walk on the $d$-dimensional hypercube, for which the conductance and spectral gap are of the same order, making the first bound in Theorem 3.8 loose in this setting.
+The spectral gap for this chain is computed, for example, in Diaconis (1988, Chapter 3C), while the conductance is computed in Diaconis and Stroock (1991, Example 3.2); see also Levin and Peres (2017, Examples 12.16 and 13.11) and Trevisan (2013, Chapter 5) for more recent surveys.
 The Gibbs sampler shares a similar local structure to the $d$-dimensional hypercube: each update resamples a single coordinate from its exact conditional, in analogy to the single-bit flips in the hypercube walk.
+If the warm start parameter satisfies $\Omega=\mathcal{O}(1)$, Wadia (2024) proved an upper bound of $\mathcal{O}(L^{2}d^{7.5}(\log d)^{2}/\mathtt{p}_{1}^{4})$ on the mixing time of the lazy random-scan Gibbs sampler $\widetilde{P}_{\mathrm{RS}}$.
+However, a direct comparison of their results with Corollary 5.7 is not straightforward.
+This is due both to the laziness of the sampler and the fact that Wadia (2024) constrain the Gibbs sampler to a convex subset of $\mathbb{R}^{d}$, which leads them to analyze mixing via the more restrictive notion of $s$-conductance.
 5.2 Non-log-concave examples
@@
 (2026a); Wadia (2024).
-Continuous Gaussian mixture.
-For all $t\in\mathbb{R}^{d}$, let $\varphi_{t}(\mathrm{d}x)\propto\exp(-|x-t|^{2}/2)\mathrm{d}x$ denote a Gaussian distribution with mean $t$ and covariance matrix $I_{d}$.
-Define the mixing distribution
-$\nu(\mathrm{d}t)\propto e^{-\left\lvert\left\lvert t\right\rvert^{2}-4\right\rvert}\mathrm{d}t.$
-The resulting Gaussian mixture
-$\pi_{1}(\mathrm{d}x)=\int\varphi_{t}(\mathrm{d}x)\nu(\mathrm{d}t)$
-is log-smooth and satisfies $\mathrm{LSI}_{2}$ (Chen et al., 2021, Subsection 4.2). However, since $\nu$ is multimodal, $\pi_{1}$ is not log-concave.
-Discrete Laplace mixture.
-Let $m_{1},m_{2}\in\mathbb{R}^{d}$ with $m_{1}=-m_{2}=(1,1,\dots,1)$. Consider the two-component mixture of Laplace distributions
-$\pi_{2}(\mathrm{d}x)\propto\left(\frac{1}{2}\cdot e^{-\left\lvert x-m_{1}\right\rvert}+\frac{1}{2}\cdot e^{-\left\lvert x-m_{2}\right\rvert}\right)\mathrm{d}x.$
-This distribution is log-Lipschitz continuous, and satisfies $\mathrm{PI}_{2}$ as shown by Chafaï and Malrieu (2010); Schlichting (2019). However, $\pi_{2}$ is not log-concave.
+Spherical log-concavity without global log-concavity.
+Consider the distribution
+$\pi_{1}(\mathrm{d}x)\propto e^{-\left\lvert\left\lvert x\right\rvert-1\right\rvert}\mathrm{d}x.$
+This distribution is log-Lipschitz continuous but not log-concave. Writing the potential function of $\pi_{1}$ as
+[…]
+with $u(t)=\left\lvert t-1\right\rvert$, which is convex for $t>0$, $\pi_{1}$ is nevertheless spherically log-concave and symmetric in the sense of Bobkov (2003); Bonnefont et al.
+(2016).
+It follows from their results that $\pi_{1}$ satisfies $\mathrm{PI}_{2}$.
 Perturbed Laplace distribution.
 Consider the distribution
-$\pi_{3}(\mathrm{d}x)\propto e^{-\left\lvert x\right\rvert+2\cos\left\lvert x\right\rvert}\mathrm{d}x.$
-This distribution is log-Lipschitz continuous and satisfies $\mathrm{PI}_{1}$ (Bobkov and Houdré, 1997). However, the added perturbation prevents $\pi_{3}$ from being log-concave.
+$\pi_{2}(\mathrm{d}x)\propto e^{-\left\lvert x\right\rvert+2\cos\left\lvert x\right\rvert}\mathrm{d}x.$
+[…]
+Gaussian mixture.
+Let $m_{1},m_{2}\in\mathbb{R}^{d}$ with $m_{1}=-m_{2}=(1,1,\dots,1)$, and consider the two-component mixture of isotropic Gaussian distributions
+$\pi_{3}(\mathrm{d}x)\propto\left(\frac{1}{2}\cdot e^{-\left\lvert x-m_{1}\right\rvert^{2}/2}+\frac{1}{2}\cdot e^{-\left\lvert x-m_{2}\right\rvert^{2}/2}\right)\mathrm{d}x.$
+This distribution is log-smooth but not log-concave.
+Nevertheless, since each Gaussian distribution satisfies $\mathrm{LSI}_{2}$ (Gross, 1975), it follows from Schlichting (2019, Theorem 2) that $\pi_{3}$ also satisfies $\mathrm{LSI}_{2}$.
+Moreover, each conditional distribution of $\pi_{3}$ can be written as
+$\pi_{3}(\mathrm{d}x_{k}\mid x_{-k})\propto\left(w(x_{-k})\cdot e^{-\left\lvert x_{k}-1\right\rvert^{2}/2}+(1-w(x_{-k}))\cdot e^{-\left\lvert x_{k}+1\right\rvert^{2}/2}\right)\mathrm{d}x_{k},$
+where $w(x_{-k})\in[0,1]$.
+This is a two-component mixture of one-dimensional Gaussian distributions centered at $1$ and 1-1.
+As shown in Chafaï and Malrieu (2010, Corollary 4.5); Schlichting (2019, Corollary 1), the $L^{2}$-Poincaré constant of such mixtures remains uniformly bounded over all values of the mixing weight $w(x_{-k})$.
+Hence, the conditional distributions of $\pi_{3}$ satisfy a uniform $L^{2}$-Poincaré inequality.
 6 Conclusion
-We established verifiable conductance bounds for systematic-scan and random-scan Gibbs samplers over a broad class of target distributions, particularly those satisfying a Poincaré or log-Sobolev inequality and possessing sufficiently regular conditional distributions (Theorem 4.7 and Corollary 4.9).
+We established explicit conductance bounds for systematic-scan and random-scan Gibbs samplers over a broad class of target distributions, namely those satisfying a Poincaré or log-Sobolev inequality and possessing sufficiently regular conditional distributions (Theorem 4.7 and Corollary 4.9).
 These bounds directly provide estimates on the mixing time in total variation distance and remain valid for either log-Lipschitz or log-smooth target distributions.
-In doing so, we extend prior results beyond the log-concave setting (Ascolani et al., 2024a; Wadia, 2024) and present several examples.
-Our work also complements recent results showing that the spectral gaps and mixing times of systematic-scan and random-scan Gibbs samplers are closely related (Gaitonde and Mossel, 2024; Chlebicka et al., 2025).
+In doing so, we extend prior results beyond the log-concave setting (Ascolani et al., 2026a; Wadia, 2024) and present several examples.
+Our work also complements recent results showing that the spectral gaps and mixing times of systematic-scan and random-scan Gibbs samplers are closely related (Gaitonde and Mossel, 2026; Chlebicka et al., 2025).
 While our analysis focuses on coordinate-wise updates, our results naturally extend to blocked Gibbs samplers under corresponding block-smoothness conditions, allowing for joint updates of coordinate subsets.
@@
 Their results, combined with our bounds, provide a framework for studying the convergence of broader coordinate-wise sampling methods.
-Regarding dimension dependence, it is clear that our bounds on the spectral gap and mixing time are not tight, particularly for Gaussian targets (Amit, 1996; Roberts and Sahu, 1997) and, more generally, for log-smooth and strongly log-concave distributions (Ascolani et al., 2024a).
+Regarding dimension dependence, it is clear that our bounds on the spectral gap and mixing time are not tight, particularly for Gaussian targets (Amit, 1996; Roberts and Sahu, 1997) and, more generally, for log-smooth and strongly log-concave distributions (Ascolani et al., 2026a).
 We attribute this to the generality of our assumptions, which might not fully exploit the underlying structure of these distributions, and to the use of total variation distance in defining the mixing times.
@@
 Acknowledgments
-Alexander Goyal is fully funded by the Roth scholarship. George Deligiannidis is supported by the Engineering and Physical Sciences Research Council [grant number EP/Y018273/1].
+AG was fully funded by the Roth Scholarship provided by the Department of Mathematics, Imperial College London.
+GD was supported during a large part of this work by the Engineering and Physical Sciences Research Council [grant number EP/Y018273/1].
 Appendix A Appendix
@@
 Define the function $N_{q}(x)=\left\lvert x\right\rvert^{q}\cdot\log(1+\left\lvert x\right\rvert^{q})$ for all $x\in\mathbb{R}^{d}$.
-Given a function $f:\mathbb{R}^{d}\to\mathbb{R}$, we denote its Orlicz norm associated to $N_{q}$ by
+Given a function $f:\mathbb{R}^{d}\to\mathbb{R}$, denote its Orlicz norm associated to $N_{q}$ by
 $\left\|f\right\|_{N_{q}}=\inf\left\{\lambda>0:\int N_{q}\left(\frac{f(x)}{\lambda}\right)\pi(\mathrm{d}x)\leq 1\right\}.$
@@
 We now introduce two functional inequalities in the Orlicz space.
-Let $\mathtt{c}_{q}>0$ and $\tilde{\mathtt{c}}_{q}>0$ denote the largest constants such that, for all locally Lipschitz functions $f:\mathbb{R}^{d}\to\mathbb{R}$,
-$\mathtt{c}_{q}\left\|f-\mathbb{E}[f]\right\|_{N_{q}}^{q}$ $\leq\mathbb{E}\left[\left\lvert\nabla f\right\rvert^{q}\right],$ (17)
-$\tilde{\mathtt{c}}_{q}\left\|f-m[f]\right\|_{N_{q}}^{q}$ $\leq\mathbb{E}\left[\left\lvert\nabla f\right\rvert^{q}\right],$ (18)
+Let $\mathtt{k}_{q}>0$ and $\tilde{\mathtt{k}}_{q}>0$ denote the largest constants such that, for all locally Lipschitz functions $f:\mathbb{R}^{d}\to\mathbb{R}$,
+$\mathtt{k}_{q}\left\|f-\mathbb{E}[f]\right\|_{N_{q}}^{q}$ $\leq\mathbb{E}\left[\left\lvert\nabla f\right\rvert^{q}\right],$ (21)
+$\tilde{\mathtt{k}}_{q}\left\|f-m[f]\right\|_{N_{q}}^{q}$ $\leq\mathbb{E}\left[\left\lvert\nabla f\right\rvert^{q}\right],$ (22)
 where $m[f]$ denotes a median of $f$.
-Bobkov and Zegarlinski (2005, Proposition 3.1) showed that $\mathrm{LSI}_{q}$ is equivalent to the functional inequality (17), with constants satisfying $\mathtt{c}_{q}/16\leq\mathtt{s}_{q}\leq 7\mathtt{c}_{q}$.
-Moreover, by Milman (2009a, Lemma 2.1), we have $\tilde{\mathtt{c}}_{q}/2\leq\mathtt{c}_{q}\leq 3\tilde{\mathtt{c}}_{q}$.
+Bobkov and Zegarlinski (2005, Proposition 3.1) showed that $\mathrm{LSI}_{q}$ is equivalent to the functional inequality (21), with constants satisfying $\mathtt{k}_{q}/16\leq\mathtt{s}_{q}\leq 7\mathtt{k}_{q}$.
+Moreover, by Milman (2009a, Lemma 2.1), $\tilde{\mathtt{k}}_{q}/2\leq\mathtt{k}_{q}\leq 3\tilde{\mathtt{k}}_{q}$.
 The functional inequality (22) is therefore equivalent to $\mathrm{LSI}_{q}$, with constants satisfying
-$\frac{\tilde{\mathtt{c}}_{q}}{32}\leq\mathtt{s}_{q}\leq 21\tilde{\mathtt{c}}_{q}.$ (19)
+$\frac{\tilde{\mathtt{k}}_{q}}{32}\leq\mathtt{s}_{q}\leq 21\tilde{\mathtt{k}}_{q}.$ (23)
 Now, take any locally Lipschitz function $f$ with $m[f]=0$ and define $g=\mathrm{sgn}(f)\cdot\left\lvert f\right\rvert^{r/q}$, where $\mathrm{sgn}(\cdot)$ denotes the sign function.
@@
 Applying (22) to $g$ yields
-${\tilde{\mathtt{c}}_{q}}^{1/q}\left\|g\right\|_{N_{q}}\leq\left\|\left\lvert\nabla g\right\rvert\right\|_{q}.$ (20)
-Since the solution to the equation $1/u+1/r=1/q$ is $u=qr/(r-q)$, an application of Hölder’s inequality gives
+${\tilde{\mathtt{k}}_{q}}^{1/q}\left\|g\right\|_{N_{q}}\leq\left\|\left\lvert\nabla g\right\rvert\right\|_{q}.$ (24)
+Since the solution to the equation $1/u+1/r=1/q$ is u=qr/(rq)u=qr/(r-q), an application of Hölder’s inequality gives
 $\left\|\left\lvert\nabla g\right\rvert\right\|_{q}$ $\leq\frac{r}{q}\left\|\left\lvert f\right\rvert^{r/q-1}\cdot\left\lvert\nabla f\right\rvert\right\|_{q}$
@@
 $\frac{r-q}{qr}=\frac{1}{r}\left(\frac{r}{q}-1\right).$
-The last inequality follows from Bobkov and Zegarlinski (2005, Lemma 3.2). Furthermore, for all $\lambda>0$, we have
+The last inequality follows from Bobkov and Zegarlinski (2005, Lemma 3.2). Furthermore, for all $\lambda>0$,
 $N_{q}\left(\frac{g}{\lambda}\right)$ $=\frac{\left\lvert f\right\rvert^{r}}{\lambda^{q}}\cdot\log\left(1+\frac{\left\lvert f\right\rvert^{r}}{\lambda^{q}}\right)$
@@
 Inserting these identities into (24) gives
-${\tilde{\mathtt{c}}_{q}}^{1/q}\left\|f\right\|_{N_{r}}^{r/q}\leq\left(\frac{5}{4}\right)^{1/q-1/r}\frac{r}{q}\left\|f\right\|_{N_{r}}^{r/q-1}\cdot\left\|\left\lvert\nabla f\right\rvert\right\|_{r}.$
+${\tilde{\mathtt{k}}_{q}}^{1/q}\left\|f\right\|_{N_{r}}^{r/q}\leq\left(\frac{5}{4}\right)^{1/q-1/r}\frac{r}{q}\left\|f\right\|_{N_{r}}^{r/q-1}\cdot\left\|\left\lvert\nabla f\right\rvert\right\|_{r}.$
 Rearranging terms and raising both sides to the power $r$, we obtain
-$\left(\frac{4}{5}\right)^{r/q-1}\left(\frac{q}{r}\right)^{r}{\tilde{\mathtt{c}}_{q}}^{r/q}\left\|f\right\|_{N_{r}}^{r}\leq\mathbb{E}\left[\left\lvert\nabla f\right\rvert^{r}\right].$
+$\left(\frac{4}{5}\right)^{r/q-1}\left(\frac{q}{r}\right)^{r}{\tilde{\mathtt{k}}_{q}}^{r/q}\left\|f\right\|_{N_{r}}^{r}\leq\mathbb{E}\left[\left\lvert\nabla f\right\rvert^{r}\right].$
 Hence,
-$\left(\frac{4}{5}\right)^{r/q-1}\left(\frac{q}{r}\right)^{r}{\tilde{\mathtt{c}}_{q}}^{r/q}\leq\tilde{\mathtt{c}}_{r}.$
+$\left(\frac{4}{5}\right)^{r/q-1}\left(\frac{q}{r}\right)^{r}{\tilde{\mathtt{k}}_{q}}^{r/q}\leq\tilde{\mathtt{k}}_{r}.$
 Finally, applying (23) twice yields
@@
 Proof of Proposition 3.11.
-Let $S$ $\subseteq\mathbb{R}^{d}$ with $\pi(S)\leq 1/2$. Note that in this case, $\pi(S^{c})\geq\pi(S)$. Define the sets
-$S_{1}$ =$=\left\{x\in S:P(x,S^{c})\leq\frac{\varepsilon}{2}\right\},$
-$S_{2}$ =$=\left\{x\in S^{c}:P(x,S)\leq\frac{\varepsilon}{2}\right\},$
+Let $S$ $\subseteq\mathbb{R}^{d}$ with $0<\pi(S)\leq 1/2$. Note that in this case, $\pi(S^{c})\geq\pi(S)$. Define the sets
+$S_{1}$ $=\left\{x\in S:P(x,S^{c})\leq\frac{\varepsilon}{2}\right\},$
+$S_{2}$ $=\left\{x\in S^{c}:P(x,S)\leq\frac{\varepsilon}{2}\right\},$
 $S_{3}$ $=\left(S_{1}\sqcup S_{2}\right)^{c}.$
@@
 $\left\|P(x,\cdot)-P(y,\cdot)\right\|_{\textsc{TV}}\geq P(x,S)-P(y,S)=1-P(x,S^{c})-P(y,S)\geq 1-\varepsilon.$
-By the contrapositive of the $(\delta,\varepsilon)$-close coupling condition, this implies $\left\lvert x-y\right\rvert>\delta$, and consequently, $\mathcal{D}(S_{1},S_{2})\geq\delta$. This leads to
+By the contrapositive of the $(\delta,\varepsilon)$-close coupling condition, this implies that $\left\lvert x-y\right\rvert>\delta$, and consequently, $\mathcal{D}(S_{1},S_{2})\geq\delta$.
+This leads to
 $\int_{S}P(x,S^{c})\pi(\mathrm{d}x)$ $=\frac{1}{2}\left[\int_{S}P(x,S^{c})\pi(\mathrm{d}x)+\int_{S^{c}}P(x,S)\pi(\mathrm{d}x)\right]$
@@
 Proof of Lemma 4.2.
-The derivative of $\rho$ is given by $\rho^{\prime}(x)=\log(1/x)-1$, which is positive for $x<e^{-1}$ and negative for $x\geq e^{-1}$.
-This establishes the first part of the lemma.
-Furthermore, for $e^{-1}\leq x\leq 1$, observe that $\rho^{\prime}(x)\in[-1,0]$, implying that $\rho$ is Lipschitz continuous on this interval with constant 1.
-To prove the second part, we consider the following cases.
-First, if $x+h<e^{-1}$, then both $x$ and $x+h$ lie in the region where $\rho$ is increasing.
-Therefore, $\rho(x+h)\geq\rho(x)\geq\rho(x)-h$, as required.
-Next, consider the case where $x+h\geq e^{-1}$.
-If $x\geq e^{-1}$, since $\rho$ is decreasing and Lipschitz continuous on this region, we have as desired
-$\rho(x)-\rho(x+h)=\left\lvert\rho(x)-\rho(x+h)\right\rvert\leq h.$
-If $x<e^{-1}$, we split the integral representation of $\rho(x+h)-\rho(x)$ as follows:
-$\rho(x+h)$ $=\rho(x)+\int_{x}^{e^{-1}}\rho^{\prime}(s)\mathrm{d}s+\int_{e^{-1}}^{x+h}\rho^{\prime}(s)\mathrm{d}s$
-$\geq\rho(x)+\int_{e^{-1}}^{x+h}\rho^{\prime}(s)\mathrm{d}s$
-$\geq\rho(x)-\left(x+h-e^{-1}\right)$
-$\geq\rho(x)-h.$
-This establishes the second part of the lemma. ∎
+Since $\rho(0)=0$ and $\log(1/x)\geq 0$ for $x\in(0,1]$, the function $\rho$ is positive. The first and second derivatives of $\rho$ are given by
+$\rho^{\prime}(x)=\log\left(\frac{1}{x}\right)-1,\qquad\rho^{\prime\prime}(x)=-\frac{1}{x}.$
+Since $\rho^{\prime\prime}$ is negative on its domain, the function $\rho$ is concave.
+Moreover, $\rho^{\prime}(x)=0$ at $x=e^{-1}$, and hence $\rho$ attains its maximum at $x=e^{-1}$.
+This proves the first part of the lemma.
+For the second part, by concavity of $\rho$, a Taylor expansion at $x+h$ yields
+$\rho(x)\leq\rho(x+h)-h\rho^{\prime}(x+h).$
+Since $\rho^{\prime}(x+h)\geq-1$, it follows that $\rho(x)\leq\rho(x+h)+h$, and therefore
+$\rho(x)-h\leq\rho(x+h).$
 Proof of Lemma 4.5.
-Let $T$ be the random variable representing the number of i.i.d.
-draws from the distribution $\mathrm{Unif}[d]$ required to observe all $d$ unique values.
-Following Mitzenmacher and Upfal (2017, Subsection 2.4.1), we have
-$\mathbb{E}[T]=d\sum_{j=1}^{d}\frac{1}{j}.$
-Approximating the above sum with the logarithm, we have
-$d\cdot\log d\leq\mathbb{E}[T]\leq d\cdot\log d+d.$
-An application of Markov’s inequality finally yields
-$\mathbb{P}(\Xi=[d])$ $=\mathbb{P}(T\leq N)$
-$\geq\mathbb{P}(T\leq 4d\cdot\log d)$
-$\geq\mathbb{P}(T\leq 2(d\cdot\log d+d))$
-$=1-\mathbb{P}(T>2(d\cdot\log d+d))$
-$\geq 1-\frac{\mathbb{E}[T]}{2(d\cdot\log d+d)}$
-$\geq\frac{1}{2}.$
+For any $k\in[d]$, the probability that $k$ is not drawn is
+$\mathbb{P}(k\notin\Xi)=\left(1-\frac{1}{d}\right)^{N}\leq e^{-N/d}.$
+By a union bound,
+$\mathbb{P}(\Xi=[d])$ $=1-\mathbb{P}(\Xi\neq[d])$
+$=1-\mathbb{P}\left(\bigcup_{k=1}^{d}\{k\notin\Xi\}\right)$
+$\geq 1-\sum_{k=1}^{d}\mathbb{P}(k\notin\Xi)$
+$\geq 1-de^{-N/d}.$
+In particular, if $N=\lceil 2d\cdot\log d\rceil$,
+$\mathbb{P}(\Xi=[d])\geq 1-de^{-2\cdot\log d}=1-\frac{1}{d}.$
+Since we assume $d\geq 2$, it follows that $\mathbb{P}(\Xi=[d])\geq 1/2$. ∎
 Proof of Lemma 5.3.
-Consider a partition of the space such that $\pi=\pi\left(\theta,s\right)$ with $\theta\in\mathbb{R}^{d_{1}}$, $s\in\mathbb{R}^{d_{2}}$, where $d_{1}+d_{2}=d$.
-Write $\pi\left(\theta,s\right)\propto e^{-U\left(\theta,s\right)}$.
-The marginal distribution of $\pi$ with respect to $s$ can be expressed as $\pi\left(s\right)\propto e^{-V\left(s\right)}$, where
+Consider a partition of space such that $\pi=\pi(\theta,s)$ with $\theta\in\mathbb{R}^{d_{1}}$, $s\in\mathbb{R}^{d_{2}}$, where $d_{1}+d_{2}=d$.
+Write $\pi(\theta,s)\propto\exp(-U(\theta,s))$.
+If $U$ is $L$-Lipschitz continuous, then the gradient $\nabla U$ exists almost everywhere and satisfies $\left\lvert\nabla U(\theta,s)\right\rvert\leq L$.
+The marginal distribution of $s$ can be expressed as $\pi(s)\propto\exp(-V(s))$, where
 $V(s)=-\log\left(\int e^{-U(\theta,s)}\mathrm{d}\theta\right).$
-Additionally, the conditional distribution of $\pi$ is
-$\pi(\mathrm{d}\theta\mid s)=\frac{e^{-U(\theta,s)}\mathrm{d}\theta}{\int e^{-U(\eta,s)}\mathrm{d}\eta}.$
-If $U$ is Lipschitz continuous, then the gradient $\nabla U$ exists almost everywhere and satisfies $\left\lvert U(\theta,s)\right\rvert\leq L$. The gradient of $V$ can be written as
+Additionally, the conditional distribution given $s$ is
+$\pi(\mathrm{d}\theta\mid s)=\frac{e^{-U(\theta,s)}\mathrm{d}\theta}{\int e^{-U(\psi,s)}\mathrm{d}\psi}.$
+In particular, the gradient of $V$ can be written as
 $\nabla V(s)=\int\nabla_{s}U(\theta,s)\pi(\mathrm{d}\theta\mid s),$
-where $\nabla_{s}U(\theta,s)$ represents the gradient of $U(\theta,s)$ with respect to the coordinates of $s$. Therefore,
-$\left\lvert\nabla V(s)\right\rvert=\int\left\lvert\nabla_{s}U(\theta,s)\right\rvert\pi(\mathrm{d}\theta\mid s)\leq L$
-almost everywhere, which shows that $V$ is also $L$-Lipschitz continuous.
-Now suppose that $U$ is $L$-smooth, meaning that the Hessian $\nabla^{2}U$ exists almost everywhere and satisfies $\nabla^{2}U(\theta,s)\preceq LI_{d}$, where $I_{d}$ denotes the $d$-dimensional identity matrix.
-Writing $w^{2}=ww^{T}$ for a vector $w\in\mathbb{R}^{d_{2}}$, the Hessian of $V$ is
-$\nabla^{2}V(s)=\int\nabla^{2}U(\theta,s)\pi(\mathrm{d}\theta\mid s)+\left(\int\nabla_{s}U(\theta,s)\pi(\mathrm{d}\theta\mid s)\right)^{2}-\int\left(\nabla_{s}U(\theta,s)\right)^{2}\pi(\mathrm{d}\theta\mid s).$
-By the Cauchy-Schwarz inequality,
-$\left(\int\nabla_{s}U(\theta,s)\pi(\mathrm{d}\theta\mid s)\right)^{2}-\int\left(\nabla_{s}U(\theta,s)\right)^{2}\pi(\mathrm{d}\theta\mid s)\preceq 0.$
-Thus,
-$\nabla^{2}V(s)\preceq\int\nabla^{2}U(\theta,s)\pi(\mathrm{d}\theta\mid s)\preceq LI_{d_{2}},$
-implying that $V$ is also $L$-smooth. ∎
+where $\nabla_{s}U(\theta,s)$ represents the gradient of $U(\theta,s)$ with respect to the coordinates of $s$. Consequently,
+$\left\lvert\nabla V(s)\right\rvert\leq\int\left\lvert\nabla_{s}U(\theta,s)\right\rvert\pi(\mathrm{d}\theta\mid s)\leq L,$
+which shows that $V$ is also $L$-Lipschitz continuous.
+Suppose now that $\theta\in\mathbb{R}$ and $s\in\mathbb{R}^{d-1}$. If $U$ is $L$-smooth, then the Hessian matrix $\nabla^{2}U$ exists almost everywhere and its spectral norm satisfies
+$\left\|\nabla^{2}U(\theta,s)\right\|_{\mathrm{spec}}=\sup\left\{\frac{x^{T}\nabla^{2}U(\theta,s)y}{\left\lvert x\right\rvert\left\lvert y\right\rvert}:x,y\in\mathbb{R}^{d},\,x\neq 0,\,y\neq 0\right\}\leq L.$ (25)
+Let $z\in\mathbb{R}^{d-1}$, $z\neq 0$. Taking $x=(0,z)$ and $y=(1,0)$ in (25) gives
+$\left\lvert\nabla^{2}_{\theta s}U(\theta,s)\right\rvert=\sup\left\{\frac{z^{T}\nabla^{2}_{\theta s}U(\theta,s)}{\left\lvert z\right\rvert}:z\in\mathbb{R}^{d-1},\,z\neq 0\right\}\leq L.$
+Moreover, taking $x=y=(0,z)$ and x=y=(0,z)x=-y=(0,z) in (25) yields
+$LI_{d-1}\preceq\nabla_{ss}^{2}U(\theta,s)\preceq LI_{d-1}.$
+Writing $z^{2}=zz^{T}$ for the rank-one matrix associated with $z\in\mathbb{R}^{d-1}$, the Hessian matrix $\nabla^{2}V$ admits the decomposition
+$\nabla^{2}V(s)$ $=\int\nabla^{2}_{ss}U(\theta,s)\pi(\mathrm{d}\theta\mid s)+\left(\int\nabla_{s}U(\theta,s)\pi(\mathrm{d}\theta\mid s)\right)^{2}-\int\left(\nabla_{s}U(\theta,s)\right)^{2}\pi(\mathrm{d}\theta\mid s)$
+$=\int\nabla^{2}_{ss}U(\theta,s)\pi(\mathrm{d}\theta\mid s)-\int\left(\nabla_{s}U(\theta,s)-\int\nabla_{s}U(\theta,s)\pi(\mathrm{d}\theta\mid s)\right)^{2}\pi(\mathrm{d}\theta\mid s)$
+$=\int\nabla^{2}_{ss}U(\theta,s)\pi(\mathrm{d}\theta\mid s)-\mathrm{Var}_{\theta\sim\pi(\cdot\mid s)}[\nabla_{s}U(\theta,s)],$ (26)
+where $\mathrm{Var}_{\theta\sim\pi(\cdot\mid s)}[\nabla_{s}U(\theta,s)]$ denotes the covariance matrix of the vector $\nabla_{s}U(\theta,s)$.
+Since $\mathrm{Var}_{\theta\sim\pi(\cdot\mid s)}[\nabla_{s}U(\theta,s)]\succeq 0$, it follows from (26) that
+$\nabla^{2}V(s)\preceq\int\nabla^{2}_{ss}U(\theta,s)\pi(\mathrm{d}\theta\mid s)\preceq LI_{d-1}.$ (27)
+We now derive an upper bound on the covariance matrix $\mathrm{Var}_{\theta\sim\pi(\cdot\mid s)}[\nabla_{s}U(\theta,s)]$.
+If each one-dimensional conditional distribution $\pi(\cdot\mid s)$ satisfies $\mathrm{PI}_{2}$ with constant $\mathtt{c}_{2}$, then, for any $z\in\mathbb{R}^{d-1}$,
+$z^{T}\mathrm{Var}_{\theta\sim\pi(\cdot\mid s)}[\nabla_{s}U(\theta,s)]z$ $=\mathrm{Var}_{\theta\sim\pi(\cdot\mid s)}[z^{T}\nabla_{s}U(\theta,s)]$
+$\leq\frac{1}{\mathtt{c}_{2}}\cdot\mathbb{E}_{\theta\sim\pi(\cdot\mid s)}\left[\left(z^{T}\nabla_{\theta s}^{2}U(\theta,s)\right)^{2}\right]$
+$\leq\frac{\left\lvert z\right\rvert^{2}}{\mathtt{c}_{2}}\cdot\mathbb{E}_{\theta\sim\pi(\cdot\mid s)}\left[\left\lvert\nabla_{\theta s}^{2}U(\theta,s)\right\rvert^{2}\right]$
+$\leq\frac{L^{2}\left\lvert z\right\rvert^{2}}{\mathtt{c}_{2}}.$
+It follows that
+$\mathrm{Var}_{\theta\sim\pi(\cdot\mid s)}[\nabla_{s}U(\theta,s)]\preceq\frac{L^{2}}{\mathtt{c}_{2}}I_{d-1}.$
+Substituting this bound into (26), we obtain
+$\nabla^{2}V(s)\succeq\int\nabla^{2}_{ss}U(\theta,s)\pi(\mathrm{d}\theta\mid s)-\frac{L^{2}}{\mathtt{c}_{2}}I_{d-1}\succeq-\left(L+\frac{L^{2}}{\mathtt{c}_{2}}\right)I_{d-1}.$ (28)
+Combining (27) and (28), we conclude that
+$\left(L+\frac{L^{2}}{\mathtt{c}_{2}}\right)I_{d-1}\preceq\nabla^{2}V(s)\preceq LI_{d-1},$
+and therefore $V$ is $L^{\prime}$-smooth with
+$L^{\prime}=L+\frac{L^{2}}{\mathtt{c}_{2}}.$